A Complete Bibliography of Publications in the British Journal for the History of Science: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 September 2025 Version 1.22

Title word cross-reference

#BSHSGlobalHist [HLDS20].

$0.00 [Goo25]. $103.00 [Amu23]. $105.00 [Cro24b]. $107.95 [Tha24]. $108.00 [Hau24]. $112.50 [Tar24]. $120.00 [Fig21]. $125.00 [Bro20, Rob25]. $130.00 [Sch25c]. $150.00 [Dun23]. $19.00 [Win23a]. $19.95 [Dua21, Ste22]. $210.00 [Smi23]. $24.95 [Sta24a]. $24.95/£ [Hel20]. $25.00 [Sim22]. $26.00 [Esp25, Sta24b]. $27.00 [Düm25]. $27.50 [HA21, Sac21]. $27.95 [Far25]. $28.00 [Bet24, Ste25a]. $28.95 [BR23, Fis25, vD23]. $29.95 [Dua21, Goo20, Sle24, Ste22, Whi24]. $29.99 [Loc24]. $30.00 [Dun25, Fre25, Hol20, Kat20]. $31.95 [Sto24]. $32.00 [Hay23]. $32.50 [Git25]. $32.95 [Bla20]. $33.95 [Fis24]. $34.95 [Ris20, Van24]. $35.00 [BR25, BC25b, Bud24, HA20, Jon21, Kad24, LZ24, MWH25, Mor21a, Pal21, Sho22, Smi21b]. $37.00 [BC25a]. $39.95 [Bim20, Kuc25, Pen24]. $40.00 [Amb21b, Fig21, Hol21, Mar25a, McC23]. $44.95 [Tal21]. $45.00 [Her24, Lan22, Leh23, McC25, Nut24, Pal21, Win23b]. $455.00 [Hor23]. $49.00 [Coo24]. $49.95 [Fra20, Hem22, Web23]. $50.00 [Amb21a, Far21, Han20, Hes23, Iwa25, Roc24, Sie25b, Sle20, Spi23]. $52.00 [vD23]. $54.95 [GS21, Hin24]. $55.00 [Big25, Bro25a, Bru24, Buc21, Gre25, Hab24, Lor24, Ree25, SD24, Web24a]. $56.95 [Byf24, Dor25]. $57.00 [Ols23]. $60.00 [Bie25, Boo23a, Mül24, Ros25, Sec22, Tal21, Tom25]. $64.95 [Sch25b, Web20]. $65.00 [Aru25, Bru22, Cro24a, Cur24, Goo25, Gro24, Mox21]. $75.00 [Dun24]. $79.95 [Led21]. $82.50 [HA21]. $85.00 [Hal24]. $89.00 [Jon24]. $89.99 [Wal25b]. $90.00 [Mac20]. $99.00 [Chu23, Mar24b].

. [Mil21, Sle20, Web20].

10.14324 [Ran23]. 10.14324/111.9781800082328 [Ran23]. 111.9781800082328 [Ran23]. 1500 [Ols23]. 1660 [Exl23]. 1700 [Dun22]. 1715 [Buc21]. 1720 [Fig21]. 1750 [Ree25]. 1800 [Loc24]. 1820s [Ame20]. 1830s [Ame20]. 1840 [Far20]. 1850 [Has24, Ken25, Wal25a]. 1856-February [Bro20]. 1859 [Bro20]. 1880 [Bla20, Web20]. 1880-1950 [Wil21a]. 1890 [Dol21]. 18th [Los25]. 1900 [Hab24]. 1910 [Sle20]. 1923 [Fre24]. 1929 [For23]. 1948 [Sha22]. 1950 [Har25, Smi21b]. 1955 [Bru24]. 1960s [Pie20]. 1970s [Pie20, Gro24]. 19th [Los25].

20.00/$39.99 [Wil20]. 2002 [Sch25a]. 2015 [Ran23]. 28.00/$35.00 [Dun22]. 2nd [Fis25].

32.50/£ [Dee20]. 35.00/$38.00 [Tim20].

62 [Kan20]. 66 [BR25]. 69.99/£ [Tar22].

92 [Wal25b]. 978-0-00-829995-8 [Wil20]. 978-0-19-764303-7 [O’M25]. 978-0-19-881862-5 [Kor21]. 978-0-19-882352-0 [Ros20]. 978-0-19-884547-8 [Fre24]. 978-0-19-886395-3 [Ken25]. 978-0-19-886605-3 [Has24]. 978-0-19-887873-5 [Ste25b]. 978-0-19-889795-8 [Jal25]. 978-0-226-35542-9 [Dee20]. 978-0-226-50929-7 [Buc21]. 978-0-226-54608-7 [Hol20]. 978-0-226-55323-8 [Pal21]. 978-0-226-59226-8 [Mox21]. 978-0-226-60201-1 [HA20].

978-0-226-60568-5 [Hay23]. 978-0-226-61557-8 [Sta24b]. 978-0-226-62210-1 [Tal21]. 978-0-226-63714-3 [Jon24]. 978-0-226-65088-3 [Hol21]. 978-0-226-68072-9 [Fig21]. 978-0-226-68086-6 [Fig21]. 978-0-226-69304-0 [Smi21b]. 978-0-226-70615-3 [HA21]. 978-0-226-70629-0 [HA21]. 978-0-226-70680-1 [Far21]. 978-0-226-71070-9 [Dun22]. 978-0-226-73160-5 [Leh23]. 978-0-226-73188-9 [Sac21]. 978-0-226-73286-2 [Sim22]. 978-0-226-74129-1 [Bru22]. 978-0-226-75250-1 [Pal21]. 978-0-226-76138-1 [Lan22]. 978-0-226-81753-8 [Mar25a]. 978-0-226-81934-1 [Cro24b]. 978-0-226-82148-1 [Git25]. 978-0-226-82228-0 [Tar24]. 978-0-226-82256-3 [Her24]. 978-0-226-82301-0 [BR25]. 978-0-226-82312-6 [Coo24]. 978-0-226-82326-3 [Bru24]. 978-0-226-82328-7 [Hab24]. 978-0-226-82676-9 [Kad24]. 978-0-226-82863-3 [Fre25]. 978-0-226-82877-0 [Nut24]. 978-0-226-82898-5 [Dun25]. 978-0-226-82929-6 [MWH25]. 978-0-226-83004-9 [BC25b]. 978-0-226-83051-3 [Esp25]. 978-0-2260-2038-9 [Sho22]. 978-0-2280-0843-9 [Hig24]. 978-0-2280-1053-1 [Bol24]. 978-0-2280-1753-0 [Cho25]. 978-0-231-19435-8 [Bur24]. 978-0-231-21261-8 [Cro25]. 978-0-262-04514-8 [Law21a, Law23a]. 978-0-262-04553-7 [Mül24]. 978-0-262-04573-5 [Rob25]. 978-0-262-37464-4 [Goo25].

978-0-262-53905-0 [Amb21a]. 978-0-262-53914-2 [Jon21]. 978-0-262-54361-3 [Gro24]. 978-0-262-54450-4 [Sle24]. 978-0-262-54554-9 [Dun24]. 978-0-262-54559-4 [Goo25]. 978-0-262-54807-6 [McC25]. 978-0-2620-3961-1 [Ste22]. 978-0-2620-4514-8 [Law23b]. 978-0-2625-3897-8 [Ste22]. 978-0-300-22383-5 [Mor21a]. 978-0-300-22741-3 [Amb21b]. 978-0-300-23006-2 [Mid21]. 978-0-300-23046-8 [Kat20]. 978-0-300-23610-1 [Hol22]. 978-0-300-24373-4 [Mil21]. 978-0-300-24677-3 [Jam21]. 978-0-300-26356-5 [Aru25]. 978-0-303-78341-9 [Ram24]. 978-0-367-22842-2 [Jac21]. 978-0-5000-5200-6 [Bow22]. 978-0-511-98014-5 [Sia21]. 978-0-520-39857-3 [Dha24]. 978-0-593-65447-7 [Ste25a]. 978-0-6-9123-379-6 [Bel25b]. 978-0-674-28136-3 [Gro23]. 978-0-674-91919-8 [Far23]. 978-0-674-97201-8 [Har21]. 978-0-6745-0462-2 [BR23]. 978-0-6747-3758-7 [Bow22]. 978-0-691-17507-2 [Des20]. 978-0-691-18188-2 [Goo20]. 978-0-691-21342-2 [Gue24]. 978-0-7735-5312-5 [Hem22]. 978-0-8173-1985-4 [Tal21]. 978-0-8229-4511-6 [Bim20]. 978-0-8229-4533-8 [Bro20]. 978-0-8229-4551-2 [Han20]. 978-0-8229-4552-9 [Sle20]. 978-0-8229-4576-5 [Win23b]. 978-0-8229-4579-6 [Hes23]. 978-0-8229-4609-0 [LZ24]. 978-0-8229-4630-4 [Web24a]. 978-0-8229-4639-7 [Sie25b].

978-0-8229-4649-6 [Spi23]. 978-0-8229-4659-5 [McC23]. 978-0-8229-4681-6 [Sec22]. 978-0-8229-4706-6 [Bud24]. 978-0-8229-4707-3 [Roc24]. 978-0-8229-4734-9 [Lor24]. 978-0-8229-4735-6 [Hor23]. 978-0-8229-4737-0 [Boo23a]. 978-0-8229-4740-0 [Iwa25]. 978-0-8229-4748-6 [SD24]. 978-0-8229-4751-6 [Gre25]. 978-0-8229-4780-6 [Cur24]. 978-0-8229-4790-5 [Cro24a]. 978-0-8229-4793-6 [Bro25a]. 978-0-8229-6563-3 [Ris20]. 978-0-87071-253-1 [Kuc25]. 978-0-9955462-9-5 [Ree21]. 978-1-009-12311-2 [Arn23]. 978-1-009-31486-2 [Sch25c]. 978-1-032-13916-6 [Ant24]. 978-1-032-16223-2 [Jor23]. 978-1-032-29157-4 [Law25]. 978-1-032-33718-0 [Amb24]. 978-1-032-36479-7 [Jag25]. 978-1-032-43590-9 [Lan24]. 978-1-032-53264-6 [Smi24]. 978-1-107-08872-6 [Bow23b]. 978-1-108-34823-2 [Tar22]. 978-1-108-41828-7 [Far20]. 978-1-108-41869-0 [Smi23]. 978-1-108-47819-9 [Tur21]. 978-1-108-48415-2 [Hau24]. 978-1-108-49125-9 [Loc24]. 978-1-108-49949-1 [Sha22]. 978-1-108-68845-1 [LD21]. 978-1-108-72884-3 [Coz20]. 978-1-108-74709-7 [Tur21]. 978-1-108-78126-8 [Dol21]. 978-1-108-84483-3 [Amu23]. 978-1-108-87915-6 [Ros22]. 978-1-108-96977-2 [Wal25b]. 978-1-1084-8713-9 [Di 23]. 978-1-1087-3219-2 [Di 23]. 978-1-1088-4457-4 [Kri23].

978-1-137-33953-9 [Har25]. 978-1-1385-7015-3 [Dur23]. 978-1-316-88243-6 [Hun23]. 978-1-350-19582-0 [Bro24]. 978-1-3503-2623-1 [Vir25]. 978-1-4214-2531-3 [Led21]. 978-1-4214-2552-8 [Fra20]. 978-1-4214-2673-0 [Dua21]. 978-1-4214-2889-5 [Web20]. 978-1-4214-2951-9 [vD23]. 978-1-4214-3852-8 [Ols23]. 978-1-4214-3908-2 [Dua21]. 978-1-4214-4002-6 [Düm25]. 978-1-4214-4069-9 [BC25a]. 978-1-4214-4368-3 [Ree25]. 978-1-4214-4405-5 [Fis25]. 978-1-4214-4437-6 [Pen24]. 978-1-4214-4527-4 [vD23]. 978-1-4214-4592-2 [Ros25]. 978-1-4214-4631-8 [Bie25]. 978-1-4214-4655-4 [Tom25]. 978-1-4214-4710-0 [Hin24]. 978-1-4214-4763-6 [Whi24]. 978-1-4214-4849-7 [Sch25b]. 978-1-4426-3017-8 [Mac20]. 978-1-469-67184-0 [Mar24b]. 978-1-4696-4844-6 [Bla20]. 978-1-4696-5609-0 [Chu23]. 978-1-4696-7382-0 [Van24]. 978-1-4742-9492-8 [DGJQ24]. 978-1-4780-0642-8 [Tha24]. 978-1-4780-2537-5 [Far25]. 978-1-4875-4365-5 [Hal24]. 978-1-4875-4503-1 [Cam25]. 978-1-5017-4015-2 [Web23]. 978-1-5017-6977-1 [Byf24]. 978-1-5017-7154-5 [Sto24]. 978-1-5017-7377-8 [Dor25]. 978-1-5128-2456-8 [Big25]. 978-1-5179-1546-9 [Bet24]. 978-1-5261-3138-6 [Kan20]. 978-1-5261-5303-6 [Spa25]. 978-1-5261-6135-2 [Hes25b]. 978-1-5261-7480-2 [Bol25].

978-1-64825-069-9 [Kli24]. 978-1-68458-157-3 [Sta24a]. 978-1-78023-997-2 [Mey20]. 978-1-78735-393-0 [Bee21]. 978-1-78735-394-7 [Bee21]. 978-1-78735-395-4 [Bee21]. 978-1-7891-4455-0 [Hol23a]. 978-1-78914-008-8 [Hel20]. 978-1-78914-214-3 [Tur21]. 978-1-78914-299-0 [Des21]. 978-1-78914-314-0 [Eic21]. 978-1-78914-342-3 [Smi20]. 978-1-78914-440-6 [Win23a]. 978-1-78914-505-2 [Web24b]. 978-1-78914-639-4 [And24]. 978-1-78962-081-8 [Guf21]. 978-1-78962-082-5 [Guf21]. 978-1-8000-8234-2 [Ran23]. 978-1-80008-415-5 [Koh23]. 978-1-85124-553-6 [Mar20]. 978-1-935408-36-9 [Hay21b]. 978-1-9788-0326-8 [Fis24]. 978-1-9788-2021-0 [Dun23]. 978-2-0211-7599-8 [Bow23a]. 978-2-503-58806-3 [Nia24]. 978-2-84137-408-3 [Mat25]. 978-3-030-78972-5 [Exl23]. 978-3-030-85531-4 [Bad24]. 978-3-031-27509-8 [Bud25]. 978-3-031-29796-0 [Mar25b]. 978-3-031-34955-3 [Los25]. 978-3-525-30206-4 [Hof24]. 978-3-525-30209-5 [Sch25a]. 978-3-662-63932-0 [Ram24]. 978-3-666-30209-1 [Sch25a]. 978-3-7965-4873-4 [Cou25a]. 978-3-8353-3680-3 [Sto23]. 978-81-948316-0-0 [Bah24]. 978-90-04-41246-0 [Wal25a]. 978-90-04-41917-9 [Off25]. 978-94-6301-434-2 [Hol23b]. 978-94-6372-106-6 [Bel25a]. 978-9500845-399-0 [Man24]. 978-981-322-868-9 [McM20].

978-981-327-903-2 [Tim20]. 978-987-628-598-8 [Man24]. 987-1-1088-3708-8 [GS22]. 987-1-4214-3874-0 [GS21].

Aaron [Sch25b]. Abbate [Pen24]. Abstractions [Pen24]. Academic [DGJQ24, Vir25]. Academy [Ito24, Ken25, Ste25b, Nia24]. access [Bee21, Goo25, Koh23, Ran23, Sch25a]. Accuracy [Bel25a]. accurate [Sto22]. Achille [Sta22]. Activism [Dor25]. Adapt [Bol24]. Adaptation [Law21b]. Addiction [Sac21, Kan24]. Address [Boo23b]. Adjusting [Low21]. Adler [Har21]. Adrian [Git25]. Advice [Cro24a]. advising [Ito24]. aerodynamic [Spi20]. aeroplane [Iqb24]. Aesthetic [Fig21]. Aesthetics [Amb24]. Affect [Bet24]. afford [Hua21]. Africa [GS22, Loc24]. African [Hef24, dB20]. After [vD23]. afterlives [Sie25a]. against [Med23]. Age [Cur24, Des20, Des21, Hay21a, Hin24, Sec22, Ker20, Gep25, Roc24, Sim22]. ageing [Hau24]. Agriculture [Van24]. Ahmed [Tek22]. Aileen [Ran23]. Aires [Man24]. airliner [Spi20]. Airy [Ish20]. Alabama [Tal21]. Alan [DGJQ24, Hig24, Ram24]. Alaska [Sch22a]. Albert [Bro25b]. Alberti [And24]. Alchemy [Dun22]. alcohol [ASW21]. Alex [Pal21, Sho22]. Alexander [Fig21, Sia21]. Alexis [Exl23]. Alfred [Bel25b, Tal21]. Alfredo [Hol23b]. Alida [Ols23]. Alison [Exl23]. Allan [BC25a, Ste22]. allowances [Ayl21]. Almond [Hes25b]. Almond-Brown [Hes25b]. along [Bro24]. alternatives [BS20]. Amanda [Tur21, Dha24]. Ambassadors [Dor25]. Ambition [Wyn21]. ambitions [Tur23]. Ameghino [Man24]. Amelia [Han20]. America [Git25, Goo20, Her24, Mar25a, Sac21, Sec22, Whi24]. American [Aru25, Bim20, Bud24, Dun23, Fra20, Jon21, Kli24, Led21, Sch25b, Bei21]. amidst [LW24]. Amir [Sha22]. Ammonia [Bad24]. Amsterdam [Bel25a]. analogy [Sch22b]. ancêtres [Bow23a]. ancestors [Sec25]. Ancient [Nut24, Sia21, Spi23, Whi24, KW21]. Anctil [Bol24, Hem22]. Andreas [Boo23a]. Andrei [Hay21b]. Andrew [Kor21, Dun23, Jor23, Mil21]. Andrews [Smi20]. Animal [Bol25, Dee20, Hol20, Cow24, Hol23c, Bol24]. Animals [Bol24, Fis25, Tar24, Tom25]. Anita [Fis25]. Anna [Nia24]. Annapurna [Goo25]. Anning [Ree21]. annotée [Mat25]. Anthony [Fre24]. anthrax [Sta22]. Anthropocene [Hay21a, LZ23, Sch22a]. anthropologie [Cou25a]. anthropologist [AA24]. Anthropology [Roc24]. Anthropozoic [LZ23]. Anti [Hau24, Ros23]. Anti-ageing [Hau24]. Anti-voluntarism [Ros23]. Anticipating [Car21]. Antiquity [GS21]. Antonio [LZ23]. Antony [Har21]. Anxiety [Dun23]. Anxious [Han20]. Aotearoa [Bar25]. Appetite [Smi21b]. appliances [KW21]. appreciation [Smi21a]. Approach [Jal25, Wid23]. Approaches [Vir25]. AR$3 [Man24]. AR$9 [Man24]. Arabic [Bar24, KK22]. arbres [Bow23a]. Archaeology [Ker20]. Archer [Guf21]. Archibald [Win23a]. architectural [Fre24]. Architecture [Hor23]. archive [AA24]. archives [MS21]. Arctic [McC23]. Argentina [Amu23, dA20]. Argentinos [Man24]. Arguing [GO24]. Argument [Law25]. Arias [Mox21]. Aristotle [Fis25, Iwa25, Jor23]. Armed [Led22]. Arnaldo [Rev25]. Aronova [Lan22]. Art [Hay21b, McC25, Smi24, BRSW23, EW23, Mau22]. Arthur [SB25]. Artificial [Sle24]. Artisanal [Amb21b]. Artisanship [Dha24]. Arts [Amb21b, Jon23, Mar24a]. Ash [Sch25a]. Asia [BH24, Car21]. Asian [Ano21h, Wil21a]. Assembling [Bow22]. Assessment [HA20, Val25]. Astrologers [Pfe21]. astrology [Gar24, Pfe21]. Astronomer [Mül24, Car24]. astronomical [Bar24, Gil20]. Astronomy [Sle20, Tur23]. Astrostereon [Bie20]. Athenæum [Jam21]. Atlantic [Far20, Loc24, Ols23, Sto24]. atmosphere [Fra21]. atmospheres [Ano21h, Wil21a]. atmospheric [Ame20]. atom [Ada23]. atomic [Ker20]. attempt [Ame20]. auctions [Mac22]. Audra [Dua21]. Australia [Far25]. Australian [Win23b, RFV25]. Author [Ree25]. authority [Car24, Dis24, Kir20]. Authorship [Pal21]. automatic [Ish20]. Avant [Smi24]. Avant-Garde [Smi24]. aviation [Iqb24, Spi20].

B [Buc21, Mil21, Ris20, Sie25b]. B. [Gar24]. Babu [Bah24]. Back [Ano20a, Ano20c, Ano20e, Ano21b, Ano21d, Ano21f, Ano22a, Ano22c, Ano22e, Ano22g, Ano23a, Ano23c, Ano23e, Ano24a, Ano25a, Ano25c]. background [Led22]. Bacterial [Web24a]. balance [Hol23c]. Baltimore [BC25a, Bie25, Dua21, Düm25, Fis25, Fra20, GS21, Hin24, Led21, Ols23, Pen24, Ree25, Ros25, Sch25b, Tom25, Web20, Whi24, vD23]. Banks [Mor21a]. Baranski [Lor24]. Barnett [vD23]. Barrett [Kri23]. Barton [Bro20]. Basel [Cou25a]. Battle [Web24b]. Battles [Amb21a]. Baumgartner [Wal25a]. Bay [Mor21a]. Beatriz [Man24]. Beauty [Amb24, Her24]. Beech [Tim20]. Beeley [Ken25]. Before [Buc21]. Behavior [Hol20]. Beijing [Tim20]. Beinart [GS22]. Being [Bee21]. Belknap [Far23]. Bell [Mar25b]. Bellhouse [Cam25]. Belyakov [McM20]. Ben [Chu23, Far20]. Beneath [Tha24]. Bengal [Raj24]. Benito [Mox21]. Benjamin [Kli24, Bad24, Gor22, Ste25a, Wil20]. Bennet [Jac21]. Benoît [Mat25]. Benson [HA21]. Benton [Bow22]. Berkeley [Dha24]. Berlin [Ram24, Kai22, Mac22]. Bernard [Jac21]. Bernardo [Law25]. Bert [Cou25a]. Bertucci [Amb21b, Hin24]. between [Hal24, Hol23b, Kem23]. Beyond [Goo25, Ken25]. Bhattacharya [Cho25]. Bible [Bie20]. Biblem [BC25a]. Biblical [Mox21]. Biddell [Ish20]. Big [Sac21]. Bihari [BR25]. Biological [O’M25]. biologists [Lon20]. Biology [Far25, Gro23, Hop22, Aru25, Cro24b, Hor23]. biomedicine [McG21]. birds [Nix24]. Birkhäuser [Ram24]. Birmingham [Wyn21]. birth [Hop22, Kom24, LW24, Wes21, Coo24]. bitter [Pen24]. BJH [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, Ano24b, Ano24a, Ano25b, Ano25a, Ano25d, Ano25c]. Black [Gro23, Rev25]. Blair [Sto24]. Blankaart [Kle20]. Blind [Dor25]. Blobel [Cro24b]. Block [Boo23a]. Blood [Jor23]. Bloomsbury [Bro24, DGJQ24, Vir25]. Bodies [Whi24, ASW21]. Bodleian [Mar20]. Body [Nut24, Coo24]. Boeck [Ram24]. Bolivia [Chu23]. Bollmer [Bet24]. Bolshevik [Lan22]. Bombings [Kad24]. Bonea [Han20]. Book [Amb21a, Amb21b, Amb24, Amu23, And24, Ant24, Arn23, Aru25, Bad24, Bah24, BC25a, Bee21, Bel25a, Bel25b, BR23, BR25, Bet24, Bie25, Big25, Bim20, Bla20, Bol24, Bol25, Boo23a, Bow22, Bow23a, Bow23b, BC25b, Bro20, Bro24, Bro25a, Bru22, Bru24, Buc21, Bud24, Bud25, Bur24, Byf24, Cam25, Cho25, Chu23, Coo24, Cou25a, Coz20, Cro24a, Cro24b, Cro25, Cur24, Dee20, Des20, Des21, Dha24, Di 23, Dol21, Dor25, Dua21, Düm25, Dun24, Dun25, Dun22, Dun23, Dur23, Eic21, Esp25, Exl23, Far20, Far21, Far23, Far25, Fig21, Fis24, Fis25, Fra20, Fre25, GS21, GS22, Git25, Goo25, Goo20, Gre25, Gro23, Gro24, Gue24, Guf21, Hab24, Hal24, Han20, Har21, Har25, Has24, Hau24, Hay21b, Hay23, Hel20, Hem22, Her24, Hes23, Hes25b]. Book [Hig24, Hin24, Hof24, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, HA21, Hor23, Hun23, Iwa25, Jac21, Jag25, Jal25, Jam21, Jon21, Jon24, Jor23, Kad24, Kan20, Kat20, Ken25, Kli24, Koh23, Kor21, Kri23, Kuc25, Lan22, Lan24, Law21a, Law23a, Law25, Law23b, Led21, Leh23, LD21, Loc24, Lor24, Los25, LZ24, Mac20, Man24, Mar25a, Mar20, Mar24b, Mar25b, Mat25, McC23, McC25, McM20, Mey20, Mid21, Mil21, MWH25, Mor21a, Mox21, Mül24, Nia24, Nut24, O’M25, Off25, Ols23, Pal21, Ram24, Ran23, Ree21, Ree25, Ris20, Rob25, Roc24, Ros22, Ros25, Ros20, Sac21, Sch25a, Sch25b, Sch25c, Sec22, Sha22, Sho22, Sia21, Sie25b, Sim22, Sle20, Sle24, Smi20, Smi21b, Smi23, Smi24, Spa25, Spi23, Sta24a, Sta24b, Ste25a, Ste22, Ste25b]. Book [Sto23, Sto24, SD24, Tal21, Tar22, Tar24, Tha24, Tim20, Tom25, Tur21, Van24, Vir25, Wal25a, Wal25b, Web20, Web23, Web24a, Web24b, Whi24, Wil20, Win23a, Win23b, vD23]. Books [And24, Des21, Eic21, Hay21b, Hel20, Hol23a, Mey20, Smi20, Tur21, Web24b, Win23a, EW23]. Borch [Hol23a]. Borch-Jacobsen [Hol23a]. Border [Amu23]. Boris [Wor22]. Boscani [Wal25a]. Botanic [Cou25b]. botanical [Kai22]. Botany [Mar20, Mor21a]. Boundaries [Cro24b]. Bowl [Wid23]. Box [Gro23]. Boy [Wil20]. Boyle [Leh23]. Brahe [Car24]. Brain [Her24, Law21b]. brains [Boo23b]. Brandeis [Sta24a]. Bravo [Hel20]. Brazil [Amu23]. Bread [Mid21]. Breeding [Tar24]. Brepols [Nia24]. Bresalier [Har25]. Breyfogle [Ris20]. brig [Bek20]. Brill [Off25, Wal25a]. Britain [Far21, Ayl21, Cow24, Gre25, Han20, Hau24, Hol22, Hun23, Jon23, Lon20, McA22, Sta22]. British [Cro24a, Har25, Kan20, Ros22, Sec22, Dav21, DUS23, Hua21, Spi20, Ste24, Wyn21]. broadcasting [Pie20]. Broberg [Gue24].

Broughton [Mac20]. Brown

[Coz20, Hes25b, BR25]. Browne [Bro20]. Bruce [Led21]. Brunswick [Dun23, Fis24]. Brysse [HA20]. BSHS [BS20]. Bud [Bee21]. Buenos [Man24]. building [Bur22, Shi24]. Bulletin [Ada23]. Buning [Goo25]. Burkhardt [Bow23b]. Burnet [Ros23]. Business [Sta24b, SB25]. Byzantine [Bar24].

C

[Aru25, Bel25a, Gro24, Hol20, Los25, Ols23, SB25, Ste22, Ano21h, Bec20, Dol21, Ish20, Pfe21, Wal25b, Wil21a, Wil21b, Wor22, tH23]. C. [Led21]. CA$47.95 [Cho25]. CA$49.95 [Hig24]. CA$65.00 [Cam25]. Cain [Goo20]. calculation [Sto22]. California [Dha24]. Cambridge [Amb21a, Amu23, Arn23, BR23, Bow22, Bow23b, Coz20, Di 23, Dol21, Dun24, Far20, Far23, GS22, Goo25, Gro23, Gro24, Hau24, Hun23, Kri23, LD21, Loc24, McC25, Mül24, Rob25, Ros22, Sch25c, Sha22, Sia21, Sle24, Smi23, Tar22, Tur21, Wal25b, Har21, Jon21, Pal21, Ste22]. Camilla [Ran23]. campaign [Ayl21]. can [Hua21, Val25]. Cancer [MWH25, Med23]. Caneva [Rob25]. Cañizares [Ant24]. Capacity [Hes25b]. Capek [Sle24]. Capitalism [Mil21]. capitalist [Gil20]. Caribbean [Kan20]. Caribou [Sch22a]. Carola [Hof24]. Carolina [Bla20, Chu23, Mar24b, Van24]. Caroline [Guf21]. Cartography [Hay23]. Cartoon [DUS23]. Cary [McC25]. case [Gar24, Kir20]. cases [BH24]. Cassidy [Ste22]. Castro [SD24]. Catherine [Dun24]. Catholic [Car24]. Catholicism [dA20]. causes [Ame20]. cave [Mat22]. Cejková [Sle24]. celebrification [GL25]. Cell [Cro24b]. Centaur [Bie25]. centre [Wyn21]. centring [BEH^{+}24a]. Century [Bro25a, Cro24a, Gre25, Han20, Hor23, Jon21, Los25, Pal21, Spa25, Cow24, Dis24, Gar24, Jon23, KK22, Kir20, Mac22, McA22, Pfe21, Raj24, Spi20, Sta22, Sto21, Val25, Wyn21, dÁ24, Bee21, Cur24, Far21, Hay21b, LZ24, McC23, Ros22]. Cesare [Omo24]. Cesarean [Fra20]. Cevdet [Tek22]. Chain [Jal25]. Chair [Mül24]. Chakrabarti [GS21]. Chakrabarty [Sim22, Sta24a]. challenge [Gil20]. Challenging [Bro25a]. Cham [Bad24, Bud25, Los25, Mar25b, Ram24]. chamber [Ber21]. Change [Chu23, Hay21a, YJ21]. Chapel [Bla20, Chu23, Mar24b, Van24]. chapter [Tek22]. character [Hes25a]. Charles [Bow23b, Wil20, Win23a, Tal21]. Charlotte [Tur21]. Charting [Fre24]. Chauvinist [Jon24, Jon25]. Chemistry [DGJQ24, Käl22, Som23, DGJQ24, Dun24, Hol23b]. Chennai [Tim20]. Chernobyl [McM20]. CHF [Cou25a]. Chicago [BR25, BC25b, Bru22, Bru24, Buc21, Coo24, Cro24b, Dee20, Dun25, Dun22, Esp25, Far21, Fig21, Fre25, Git25, Hab24, Hay23, Her24, Hol20, Hol21, HA20, HA21, Jon24, Kad24, Lan22, Leh23, Mar25a, MWH25, Mox21, Nut24, Sac21, Sho22, Sim22, Smi21b, Sta24b, Tal21, Tar24]. China [Mil21, Kri23, LW24, Lia24]. Chinese [Bar23, Fra21]. Chris [Cur24, Mar20]. Christopher [Ken25, Mar24b]. ciné [Lon20]. ciné-biologists [Lon20]. cinema [tH23]. circa [Ols23]. Circle [Ros20, Gar24]. Circuit [Sec22, Wid23]. Circulation [Jor23, Sto21]. Civic [Kli24]. Claire [Exl23, Hol22]. Clarke [Kan20, SB25]. clash [Zai23]. Classification [O’M25]. Classroom [Dun23]. Claudine [Mat25, Bow23a]. clearing [Kai22]. Clements [Bim20]. Cliché [Jon24, Jon25]. Climate [Hay21a, Sim22, Sta24a, ASW21, RBB^{+}20]. Climério [Mar25b]. clinical [Ste24]. Clock [Ste25b]. Clocks [Hig24]. cloth [Bol24, Bru24, Coo24, Cro24b, Dun23, Git25, Hab24, Her24, Hig24, Jon24, Leh23, Nut24, Sta24b, SD24, Tar24, Tha24, Win23b]. cloud [Ber21]. Club [Bei21, Jam21]. co [Lon20]. co-production [Lon20]. Coats [BR25]. cobwebs [Hen24]. Coenraad [Off25]. Coercion [Far21]. Cohen [Bow23a, Mat25]. Cold [Dor25, Dua21, Goo20, Kad24, Kor21, Kri23, Lan22, Aso24, Kan24, Sie25a]. Colder [Hol21]. Cole [Cro25]. Colin [Nut24]. Collard [Tal21]. Collecting [Dol21, Nia24, McA22]. collection [Mac22]. Collective [Nia24]. Collins [Wil20]. Colombia [Van24]. Colonial [Dha24, Jag25, Kai22, Web23, Wil21b]. Colouring [EW23]. Colours [Los25]. Columbia [Bur24, Cro25]. Commerce [Far21]. Commercial [Kem23]. committee [Käl22]. Commodification [Hes23]. Common [Hes23]. Communicating [Big23]. Communication [Bel25a, Dun23]. communism [Hyu24]. Communist [Lia24]. Communities [Gre25]. community [For23, Som23]. Companion [Tal21]. Comparative [Bur22]. competition [Hyu24]. Complexity [Hol20]. Compound [Spi23]. compulsions [Ste24]. Computational [Boo23a]. computers [Bel23]. computing [Pen24, Pen24]. Concluding [BEH^{+}24a, BEH^{+}24b]. Condé [Bud25]. conference [For23, Wid23]. conferences [BRSW23, BS20, Rei23, RBB^{+}20, Som23, Zai23, Kot23]. conferencing [Big23, HLDS20, Wid23]. confessional [GO24]. Conflict [Van24]. Congress [Mou23]. Connecting [Wal25a]. connections [AA24]. consensus [Wes21]. Consequence [Fra20]. Conservation [Rob25, Hyu24, Sch22a, dB20]. Constructing [Ish20, Fre24]. Construction [Hig24]. consultants [Kan24]. Contemporary [Dur23, Vir25]. Contested [Bus22, Kuc25]. context [BD20, HS21, LZ23]. Contexts [Rob25]. Contraception [Wal25b]. contrasting [Tur23]. contribution [Omo24]. control [Aso24, Kan24]. conventional [BS20]. Conversation [BEH^{+}24a, BEH^{+}24b]. converted [Car24]. Cooley [Tar24]. Corbett [Bro20]. Cornell [Byf24, Sto24, Web23]. corps [Cou25a]. Correlations [Jal25]. Correspondence [Bow23b, Bro20]. Corrigendum [BEH^{+}24a, Sec21b, Sid25b, Jon25, Law23a]. Corvallis [Kuc25]. cosmic [Rev25]. cosmology [Bie20, Omo24]. Costa [Bel25b, Tal21]. Council [Ito24]. Courant [Shi24]. Cover [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, Ano24b, Ano24a, Ano25b, Ano25a, Ano25d, Ano25c]. crafting [Kot23]. Craig [Web20]. Creation [Rob25]. Creatures [Hem22, Goo20]. credibility [Mat22]. Credible [Roc24]. Cremonini [Omo24]. Crisis [Jon24, Jon25, RBB^{+}20]. CRISPR [Fis25]. Cristina [Dor25]. critical [SS23, Dee20, Win23a]. Crosland [Smi21a]. Crossing [Cro24b]. crossings [Sch22a]. Crowther [Whi24]. CS$49.95 [Bol24]. Csiszar [Pal21]. CT [Amb21b, Aru25, Kat20, Mid21, Mil21, Mor21a]. Cuban [Rev25]. cudgelling [Boo23b]. Cueto [Coz20]. Cult [Hau24]. Cultivating [LD20]. Cultural [Bee21, DGJQ24, Law21a, Law23a, Law23b, Mül24, Sch25b, Gar24]. Culture [Dol21, Guf21, Jac21, Lan24, Ros20, Spa25, Kir20, Omo24, Hel20]. Cunningham [Jor23]. curated [For23]. Curating [Nic22]. Curies [Med23]. Curiosity [Spa25]. Curious [And24]. Cutting [Jal25]. Cyrus [Gro24, Zai23].

D [Bro20, Di 23, Kat20, Ken25]. D. [Bah24]. D.E [Mar24b]. Dagmar [Goo25]. Dale [HA20]. Daniel [BR23]. Daniels [Mar25a]. dans [Bow23a]. Dark [Sta24b]. Darwin [Bow23b, Kuc25, Bow23b, Hes25a, Khr24, Sec21a, Sec21b, Win23a]. Data [Sho22, Bol23, LD20]. David [Ste22, Tal21, Win23a, Cam25, Hal24, Sac21]. De-centring [BEH^{+}24a]. Death [Web23, Hop22]. debates [Val25]. Debating [Vir25]. debt [Sto22]. Decentring [BH24, BEH^{+}24b]. decision [Käl22]. decolonization [SS23]. deconstructive [MS21]. Defining [Hes25b]. Della [Bru22]. demystification [Ame20]. Denis [Sto21]. derived [LW24]. Descartes [Sch22b]. Design [Des21]. Designs [Big25]. d’Etaples [Ros20]. d’ethnographie [Mau22]. deutschen [Sch25a]. Developing [Fre25]. Development [Bur24, Cur24, Hol23b, Kan20, Bec23, Hef24, Sie25b]. developmental [Ma25a, Ma25b]. Devices [And24]. dialogues [SS23]. Diarmid [Sec22]. Dick [Guf21, Pen24]. Dickson [Han20]. did [Car24]. Digby [Hen24]. digester [Sto21]. digestion [Sch22b]. digital [McG21]. DiMeo [Leh23]. diminution [Mat25]. Dinosaur [Bow22]. Dinosaurs [Bow22]. Dipesh [Sim22, Sta24a]. Diplomacy [BEH^{+}24a, Ada23, Aso24, Bar23, BEH^{+}24b, BH24, DUS23, GB23, Hyu24, Iqb24, Lia24, Med23, TA23, Kri23]. diplomatic [Tur23]. Diplomatie [Hof24]. disaster [Hol23c]. Discerning [HA20]. discipline [Joh21]. Disciplines [Jac21]. Discontents [Smi21b]. Discovered [Jor23]. discovery [HS21, Bad24]. discussion [Bie20]. discussions [Hua21]. disease [Sta22]. dismantling [Bur22]. disorder [Joh21]. Disorders [Hol20]. Disparate [Cho25]. Displays [Fis24]. disputed [Bus22]. Disquiet [Mox21]. distribution [Ish20]. Diversion [Lia24]. divide [GO24]. DNA [HS21]. Doctor [Hol22]. Does [Esp25]. Dog [Bol25]. Doherty [Bel25a]. Doi [Ran23]. doing [Wil21b]. Domains [O’M25]. doppelgängers [Bus22]. Dora [Bru22].

Dorner [Far21]. Dovecote [Ree21]. Down [Cow24, Kom24]. Dr [Raj24]. dread [Sta22]. Dreams [Far20]. Drugs [Sac21]. DSM [BC25a]. Dublin [Kir20]. Dubow [GS22]. Duel [Web24b]. Duke [Far25, Tha24]. d’un [Mat25]. Duncan [Des20, Fre24]. Duplicate [Kai22]. Duplicates [Mac22, Bus22, HMB22, Mar22, Nic22]. duplication [Mau22]. Durham [Far25, Tha24]. during [Kan24, KK22]. Dutch [tH23]. Dylan [Law21a, Law23a, Law23b]. Dynamics [Boo23a, Dav21].

eagle [SS23]. Earle [LD21]. Early [Bah24, Bee21, Bel25a, Bie25, Coo24, Her24, Kli24, Los25, Nia24, Ros25, Smi24, Tar22, vD23, Bar24, Bie20, Bil20, Käl22, Lia24, Mac22, Mat22, Med23, Sta22]. Earth [Bru22, Mat22, Ros23, Tim20]. Eastern [MB22]. Eat [Smi21b]. Eating [dB20]. Eaton [Zai23]. ebook [Bud25]. Eburon [Hol23b]. ecological [Hol23c]. Ecology [Ris20]. Economic [Bur24]. Economy [Mid21]. ed [Dee20, Ris20, Sle24, Vir25]. Edge [Di 23]. Edhasa [Man24]. Édition [Mat25]. Editions [Bow23a]. editors [Bow23b]. edn [Fis25]. Edney [Hay23]. eds [Ant24, Bee21, Boo23a, Bow23b, Bro20, Bud25, DGJQ24, Exl23, Goo25, Guf21, Jac21, Ken25, Koh23, Kuc25, Los25, Nia24, Pen24, Sia21, Sie25b, Smi23, SD24, Tal21, Wal25a]. education [Kan24]. Edward [Bie20]. Edwardian [ASW21]. efficiency [Spi20]. Efran [Roc24]. Egle [Byf24]. Eighteenth [Far21, Spa25, Jon23, Sto21]. Eighteenth-Century [Spa25, Jon23, Sto21]. Einstein [Ste22, Ste22]. Einstein-Marić [Ste22]. Einsteinian [Gor22]. electricity [Wes21]. electronic [McG21]. Elena [Lan22]. elite [Dav21]. Elizabeth [Coz20, Big25, Smi21b]. embedding [Led22]. Embodiments [Pen24]. Emergence [Off25]. emergency [RBB^{+}20]. emerging [Bek20]. Emma [Dun25, Far25, Mid21]. Emotion [Bet24]. Empire [Arn23, Kan20, Sto24, SW25, Bil20, KK22, Yal22]. Empires [O’M25]. Encounters [Bro24, Loc24]. End [Kan20, Lan22, Web24b, Bow24]. Energy [Rob25]. engineer [Dis24]. engineer-like [Dis24]. Engineering [Gro24, Dis24, Web20]. Engineers [SD24]. England [Bel25a]. English [Dun22]. Engraving [Bel25a]. Enlightenment [Amb21b, Guf21]. Enough [BR23, dÁ24]. Ensayo [Man24]. Ensnared [Hal24]. enters [Ker20]. Entomologists [Gre25]. entomology [Wor22]. Entretiens [Mat25]. Environment [Bim20, Düm25, vD23, HA24]. Environmental [Chu23, HA20, HA24, Sie25b, Sto24]. Environmentalisms [HA21]. Environments [HA21, Ris20]. Epilogue [Sid25a, Sid25b]. episode [Dav21]. Epistemology [Dur23]. era [Bow24, Fis24]. Erika [Koh23, Goo20]. Erratum [Ma25b]. Esdaile [Raj24]. Esguerra [Ant24]. Espionage [Hin24]. Essay [Pen24, SW25]. Essays [Cro25]. establishment [Ito24]. Esterson [Ste22]. estimate [Sto22]. établie [Mat25]. ethnographic [Bus22]. ethological [Nix24]. ethology [Kem23]. Etienne [HA21]. eugenics [Ayl21]. Eulàlia [Off25]. EUR0.00 [Sch25a]. EUR117.00 [Bel25a]. EUR119.00 [Wal25a]. EUR129.00 [Off25]. EUR23.00 [Bow23a]. EUR28.00 [Sto23]. EUR33.00 [Mat25]. EUR36.00 [Hol23b]. EUR70.00 [Sch25a]. EUR80.00 [Hof24]. EUR85.00 [Nia24]. Eurasian [Ris20]. Europe [Los25, Tar22, vD23]. European [Sto21]. Evangelists [LZ24]. Everest [Bim20]. Evolution [Hor23, Tal21, Tur21, End25, Des20, Bow23a]. Evolutionary [Hor23]. Ewing [Kor21]. examining [MB22]. examples [Yal22]. exchange [RFV25]. Exhibiting [Fis24].

Exhibitions [Win23b]. Expanses [Mey20]. Expedition [Bim20]. Experience [Jon21, HLDS20, Lan24]. experiences [MB22]. experiment [Bel22, EW23]. Experimental [Dun22]. experimentation [LD20]. Experimenting [Fis25]. Experiments [Hol20, Lan22, Hee23]. expertise [KW21, Kan24]. Experts [HA20, Kan20]. Exploration [Hol21]. Explorations [McC23]. Exploring [And24, Wal25a]. Exposition [Tom25]. Exposure [Kuc25]. expressions [Kem23]. Extreme [Bim20, Hol21]. eye [Ber21, Nix24].

Fabian [Bie25]. Fabricated [Hin24]. facial [Kem23]. Faculty [Mar24a]. failure [Wyn21]. Falling [Tim20]. Falls [Amu23]. family [Ayl21]. famine [LD20]. Fanny [Bru24]. Fantasy [Har21]. Farrah [Koh23]. Father [Cam25]. Fear [Har21]. feasts [Pfe21]. February [Bro20]. Fee [Coz20]. Feeding [LD21]. Feld [Hof24]. Felix [Sto23]. Feminist [Jon24, Jon25]. fermentation [Sch22b]. Fernández [Big25]. fertility [Bro25b]. fictive [SW25]. fight [Med23, Tho24]. Figures [Coo24]. film [Lon20, tH23]. filogenia [Man24]. Finnegan [Sec22]. Fire [Dun22]. First [Mar25b, Cou25b, Som23, YJ21]. Fisher [Ayl21]. Fiss [Dun23]. Flannery [Tal21]. Flawed [Cam25]. Fleetwood [Arn23]. Flood [Lia24, vD23]. Florentino [Man24]. flowers [EW23]. Flu [Har25]. Follow [Jor23]. Fontaine [Di 23]. food [LD20]. forecast [Car21]. forerunner [Khr24]. Forest [Hay21b]. forged [Low21]. Forging [Sle20]. Form [Hor23]. Forms [Des20]. Forum [Ano20g]. Fossil [Bow22, Ree21]. fostering [SS23]. Foundation [Bec23, Amb21a]. Foundations [Mar25b]. Founders [Aru25]. Framing [Ano21h, Wil21a]. francais [Mat25]. France [Amb21b, Gar24]. Francois [Cou25a].

Frank [Bee21, Boo23b]. Fraser [Bol25]. Frederick [Bow23b]. Frederico [Amu23]. Freedom [Dua21]. Freitas [Amu23]. French [Amb24, Buc21, Jal25]. Freud [Hol23a]. Friedrich [Los25, HS21]. fringe [Mor21b]. Fritz [Bad24]. Frøland [Web24b]. Front [Ano20b, Ano20d, Ano20f, Ano21a, Ano21c, Ano21e, Ano21g, Ano22b, Ano22d, Ano22f, Ano22h, Ano23b, Ano23d, Ano23f, Ano24b, Ano25b, Ano25d]. Frontier [Chu23, Fra21, Ros22]. Frontiers [Boo23a]. fuel [Spi20]. fuel-efficiency [Spi20]. function [Boo23b]. Functional [Kot23]. Funding [Amb21a, Bro25b]. Fungorum [Bol23]. Future [Byf24, McC25, Bek20, Big23]. Fyfe [Ran23].

G [Exl23, Sch25a]. Galenic [Spi23].

Galmarini [Dor25]. game [dB20]. Gansen [Big25]. Garde [Smi24]. Garden [Hol22]. gardener [Hes25a]. Gardens [Cou25b]. Garza [SD24]. Gas [Sch25c]. Gassendi [Gar24]. Gassó [Off25]. gastrulation [Hop22]. gathering [BRSW23]. GBP [Web24b]. Gemma [Hes25b]. gen [DGJQ24]. Genealogies [Bru22]. genealogy [McG21]. Genes [McG21]. genetic [Bur22]. Genetics [Hol20, Sha22, Bro25a]. Genius [Cam25]. genomics [Low21]. geocentric [Ma25a, Ma25b]. Geographical [Bru22]. Geographien [Sto23]. Geological [Mou23]. Geology [GS21]. Geometry [Iwa25, Wil20, MS21]. Geopolitics [Ant24]. George [Ish20]. German [Bus22]. Germany [Sha22, Sch25c, Sie25a]. Germline [Esp25]. Gesellschaft [Hof24, Sch25a]. get [Car24]. Gil [Bur24]. Gilded [Sec22]. Gin [Mar20]. Gisela [Ram24, Ram24]. Glahn [McC25]. Glaishers [Win24]. Glasgow [Cou25b]. Global [Bur24, Fre25, Web20, vD23, Bre25, Gep25, Mar22, SB25]. Globalization [Lor24, MB22, Tek22]. globalizations [Bur22]. Globalizing [Yal22]. gnosis [Bro25b]. go [McG21]. goats [Cow24]. goddess [Som23]. Going [Tim20]. Gonaver [Bla20]. Goncalves [Law25]. Gonzalo [Big25]. Good [Hes23, BR23]. Gordon [Kot23, Kri23]. Göttingen [Hof24, Sch25a, Sto23]. Graeco [Bar24]. Graft [Bro25a]. Granada [Kem23]. Grand [Hin24]. Grant [Bet24, Boo23a]. Graphics [Cam25]. gravid [Bel22]. Greece [Whi24]. Greek [Nut24]. Green [Hol23b, Van24, Spi20]. Greenhalgh [Bee21]. Greenwich [Bel23, Ish20, Joh21]. Grenoble [Mat25]. Gresham [Bie20]. Gribenski [Bru24]. Griffin [Mid21]. Group [Aru25]. Gruen [Dee20]. guardians [YJ21]. Guerrini [Fis25]. Guinea [Bus22]. Gunnar [Gue24]. Gunpowder [Wil20]. Günter [Cro24b].

H [Fra20, Hay23, Win23b]. Haber [Bad24]. Haida [Nic22, SS23]. Hamblin [Kuc25]. hammer [Mac22]. Handbook [Exl23, Smi23, Bud25]. Hannah [Koh23]. Hannan [Spa25]. Hans [Hes23]. hardback [Amb24, And24, Ant24, Arn23, Bee21, Bel25a, BR23, Bow22, Bow23b, Bro20, Bro24, Bru22, Bud24, Chu23, Di 23, DGJQ24, Dol21, Dua21, Dun22, Eic21, Far20, Far21, Far23, Fig21, Fra20, Fre24, GS21, GS22, Guf21, Har21, Has24, Hay21b, Hem22, Hof24, Hol23a, Hol21, Hol22, HA21, Jac21, Jag25, Jam21, Jor23, Kan20, Kat20, Ken25, Koh23, Kor21, Kri23, Lan22, Lan24, Law25, Led21, LD21, Mac20, Mar20, McC23, Mey20, Mid21, Mox21, Nia24, O’M25, Off25, Ols23, Pal21, Ree21, Ros22, Ros20, Sac21, Sch25c, Sec22, Sha22, Sho22, Sia21, Smi20, Smi23, Smi24, Spi23, Ste22, Ste25b, Sto23, Tal21, Tar22, Tur21, Vir25, Wal25a, Web23, Wil20, vD23]. hardback/paperback [Tar22]. hardcover [Amb21b, Amu23, Aru25, Bad24, Bah24, BC25a, Bel25b, Bie25, Big25, Bol25, Boo23a, Bro25a, Byf24, Cam25, Cro24a, Cur24, Des20, Des21, Dor25, Exl23, Goo20, Gre25, Gro23, Gue24, Hal24, Han20, Har25, Hau24, Hes23, Hes25b, Hin24, Hor23, Hun23, Iwa25, Jal25, Kli24, Lor24, Los25, LZ24, Mar24b, Mil21, Mor21a, Ram24, Ran23, Ree25, Rob25, Roc24, Ros25, Sch25a, Sch25b, Sie25b, Sle20, Spa25, Ste25a, Tom25, Wal25b, Web20, Web24a, Whi24]. Harrison [Koh23]. Hartogsohn [Jon21]. Harvard [BR23, Bow22, Gro23, Har21, Pal21]. Harvey [Jor23, McA22]. Haunting [Far25]. Haven [Amb21b, Aru25, Hol22, Jam21, Kat20, Mid21, Mil21, Mor21a]. head [Sec21a, Sec21b]. Healing [Loc24]. Health [Coz20, Far21, Fis24, Kli24, Mar24b, Dav21, Fre25, Sta24b]. heavens [Mar24a]. Heggie [Hol21]. Heidi [Fre25]. Helen [Mey20]. Helmholtz [Rob25]. Henry [EW23, McA22]. heritage [Bar24]. hermanos [Man24]. Hertz [Bru24]. Herzberg [Sac21]. Hickman [Hol22]. Hidden [Sac21, Wil21b]. High [O’M25, Wid23]. High-Level [O’M25, Wid23]. Higher [Hol21]. Hill [Bla20, Chu23, Mar24b, Van24]. Hilton [Shi24]. Himalaya [Arn23]. Hiroshima [Kad24]. Hiscock [Mar20]. Historian [Mor21a]. historical [Hee23]. Histories [Bro24, Kuc25, BH24, BRSW23, Pen24]. Historiography [Bud25, MS21]. historique [Cou25a]. History [BC25a, Bar25, Bet24, Boo23a, DGJQ24, Dor25, Düm25, Dun23, Fra20, Gro23, Hem22, Hol21, Hol23b, HA21, Kat20, Koh23, Lan22, Led21, Mey20, Mid21, Mil21, O’M25, Ran23, Sac21, Sho22, Sia21, Sim22, Smi20, Smi23, Tha24, Tim20, Tur21, Vir25, Wal25b, Bel23, Ber21, Bil20, Boo23b, Bre25, Ker20, Kle20, Kom24, Lon20, Mac22, McA22, Mor21b, MB22, SS23, SW25, Bow22, Coz20, Eic21, Hay23, Mül24, Ris20, Sch25b]. Hitler [Hal24]. Hoffenberg [Win23b]. hole [GB23]. Hollings [Ken25]. Holmes [Bro25a]. Home [Spa25]. Hong [Tim20, Hua21, Wil21b]. Hooker [Cou25b].

Hopkins [BC25a, Bie25, Dua21, Düm25, Fis25, Fra20, GS21, Hin24, Led21, Ols23, Pen24, Ree25, Ros25, Sch25b, Tom25, Web20, Whi24, vD23]. Horizontal [DGJQ24]. hormone [Bec20]. Horticulture [Hol22]. Horwitz [BC25a]. hospitalized [Raj24]. house [Kai22]. household [EW23]. Howard [Ree25, Bro25b]. HPV [MWH25]. Hudson [Bow22]. humaine [Bow23a]. Human [Eic21, Goo20, Her24, Kli24, Lan24, Tur21, Bel22, Des20, Tur21]. Humanities [Hab24, Hay21a]. Humanity [Tur21]. Humans [Fis25]. Humboldt [Ant24]. Humphreys [Iwa25]. Humphry [Des21]. Humus [Bro25b]. Hunt [Goo20]. Hunter [Bel22]. Hunters [Bow22]. Hutton [Wil20]. Hybrid [Bro25a, Fre24]. hydraulic [Val25].

IAEA [Ada23]. Ian [Des20]. iatromechanism [Sch22b]. Ibn [Tek22]. Icy [McC23]. idea [HA24, Düm25]. Ideal [Hay23]. Ido [Jon21]. Iguazu [Amu23]. Illinois [Dor25]. Illness [Kli24]. illusions [Ame20]. im [Hof24, Sch25a, Sto23]. image [TA23, McC25]. images [GB23]. Imagination [Amb24, GS22, Iwa25]. Imagining [vD23]. Impact [Bee21]. Imperial [Ris20, Ano21h, DUS23, Wil21a, Wor22]. importance [Hes25a]. impossibility [Car21]. Impossible [Sto22]. inactivated [LW24]. Incomparable [Leh23]. independence [Aso24]. India [BR25, Ros22, Bah24, Cho25, Dha24, Jag25, Kan24, Ma25a, Ma25b, Mil21, YJ21]. Indian [Iqb24, Raj24]. indien [Mat25]. Indigeneity [Far25]. Industrial [Hin24]. Industry [Bud24, Ste25b]. Inference [BC25b]. Influenza [Har25]. informality [Kot23]. Information [Git25]. infrastructure [Val25]. Inheritance [McG21]. Innovation [Guf21, LW24, RBB^{+}20]. Innovations [Mey22]. inoculation [Mey22].

Inscriptions [GS21]. insects [Kle20]. Insider [Led21]. Institute [Low21]. Institution [Bla20, Nic22]. institutions [Kai22]. instrument [Ada23]. Instruments [Mar25b, Kom24]. Intellectual [Goo25, Smi23, Bar24]. interaction [Kot23]. interested [Sto22]. Interlopers [Ros25]. International [Dor25, Sie25b, Tom25, Ada23, BRSW23, Low21, Mou23, Som23, dB20, Bec23, Mou23]. internationale [Hof24]. internationalism [Aso24, Rei23]. internationalisms [Zai23]. internationalization [Bol23]. Interwar [Sch25c, Ayl21, Lon20]. Intimate [Mid21]. Introduction [Dur23, HMB22, TA23, Hol23c]. Inventing [Dun22]. Invention [Ant24, Big25, Iwa25]. investigation [Bel22]. Investigators [Roc24]. invisible [Mat22]. Ireland [Wal25b, Kir20]. Irina [Man24, Man24]. Irons [Web24b]. Isaac [Hen24]. ISBN [Amb21a, Amb21b, Amb24, Amu23, And24, Ant24, Arn23, Aru25, Bad24, Bah24, BC25a, Bee21, Bel25a, Bel25b, BR23, BR25, Bet24, Bie25, Big25, Bim20, Bla20, Bol24, Bol25, Boo23a, Bow22, Bow23a, Bow23b, BC25b, Bro20, Bro24, Bro25a, Bru22, Bru24, Buc21, Bud24, Bud25, Bur24, Byf24, Cam25, Cho25, Chu23, Coo24, Cou25a, Coz20, Cro24a, Cro24b, Cro25, Cur24, Dee20, Des20, Des21, Dha24, Di 23, DGJQ24, Dol21, Dor25, Dua21, Düm25, Dun24, Dun25, Dun22, Dun23, Dur23, Eic21, Esp25, Exl23, Far20, Far21, Far23, Far25, Fig21, Fis24, Fis25, Fra20, Fre25, Fre24, GS21, GS22, Git25, Goo25, Goo20, Gre25, Gro23, Gro24, Gue24, Guf21, Hab24, Hal24, Han20, Har21, Har25, Has24, Hau24, Hay21b, Hay23, Hel20, Hem22, Her24]. ISBN [Hes23, Hes25b, Hig24, Hin24, Hof24, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, HA21, Hor23, Hun23, Iwa25, Jac21, Jag25, Jal25, Jam21, Jon21, Jon24, Jor23, Kad24, Kan20, Kat20, Ken25, Kli24, Koh23, Kor21, Kri23, Kuc25, Lan22, Lan24, Law21a, Law23a, Law25, Law23b, Led21, Leh23, LD21, Loc24, Lor24, Los25, LZ24, Mac20, Man24, Mar25a, Mar20, Mar24b, Mar25b, Mat25, McC23, McC25, McM20, Mey20, Mid21, Mil21, MWH25, Mor21a, Mox21, Mül24, Nia24, Nut24, O’M25, Off25, Ols23, Pal21, Pen24, Ram24, Ran23, Ree21, Ree25, Ris20, Rob25, Roc24, Ros22, Ros25, Ros20, Sac21, Sch25a, Sch25b, Sch25c, Sec22, Sha22, Sho22, Sia21, Sie25b, Sim22, Sle20, Sle24, Smi20, Smi21b, Smi23, Smi24, Spa25, Spi23, Sta24a, Sta24b]. ISBN [Ste25a, Ste22, Ste25b, Sto23, Sto24, SD24, Tal21, Tar22, Tar24, Tha24, Tim20, Tom25, Tur21, Van24, Vir25, Wal25a, Wal25b, Web20, Web23, Web24a, Web24b, Whi24, Wil20, Win23a, Win23b, vD23]. Islam [Dha24]. Islamic [Tek22]. issue [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, Ano24b, Ano24a, Ano25b, Ano25a, Ano25d, Ano25c, HMB22, Mar22]. Italian [Bec20]. Italy [Bar24, Val25]. Ithaca [Byf24, Dor25, Sto24, Web23]. itself [Low21]. Ivanova [Amb24]. ix [Coz20, HA20, Ram24].

J [Bow22, Mac20, Ram24, Ros20, Smi23]. J. [Buc21, Gar24, SD24, Win23a]. J.-B. [Gar24]. J.A [Sto24]. J.M.M [And24]. J.T [DGJQ24]. Jackson [Cou25b, Cro24a, Dun24]. Jacob [Off25, Kuc25]. Jacobsen [Hol23a]. Jacqueline [Far23, Fra20]. Jähn [Sie25a]. Jahrhundert [Sto23]. Jamaica [Hol23c]. James [Bee21, Bow23b, Dun25, Raj24, SD24, Tal21, Bel25b, Guf21, Hau24, Sto24]. Jamieson [HA20]. Jan [Kle20]. Janet [Bro20, Pen24]. Janssen [Fre24]. Japan [Ito24]. Jean [Cou25a]. Jefferson [Di 23]. Jeffrey [LZ24]. Jennifer [Han20, Dun22, Fis24]. Jenny [Sie25b]. Jérôme [Mat25]. Jersey [Tim20]. Jessica [HA20]. Jesuits [Car24]. Jingjiang [Lia24]. Jitka [Sle24]. Joan [Kor21]. JoAnne [Web20]. job [Wil21b]. Jodhpur [Iqb24]. Johan [Hol23b]. John [Bro20, Mar25a, Web24b, Hes25a, Pie20]. Johns [BC25a, Bie25, Dua21, Düm25, Fis25, Fra20, GS21, Git25, Hin24, Led21, Ols23, Pen24, Ree25, Ros25, Sch25b, Tom25, Web20, Whi24, vD23]. Johnson [Bad24]. Johnston [Hig24]. Jonas [LW24]. Jonathan [Cro25]. Jones [Exl23, Koh23, Sia21]. Jorge [Ant24]. Joseph [Bar23]. Joshua [Sle20]. Journal [Pal21]. Journals [Ran23]. Julie [Ran23]. Just [Jam21, Wil21b]. Justin [SD24, Iwa25].

Kaalund [McC23]. Kalle [Loc24].

Kamlesh [Jag25]. Kampourakis [Dur23]. Kananoja [Loc24]. Kantorovich [BD20]. Karel [Sle24]. Karl [Cro24b]. Kate [Esp25]. Kathleen [Whi24]. Katrine [McC23]. Katzir [Ste25b]. Keith [Kor21]. Keller [Nia24, Ros25]. Kelly [Wal25b]. Ken [Bro20]. Kenelm [Hen24]. Kenneth [Rob25]. Kerri [Smi20]. Kevin [Dur23]. Kew [Mor21a]. Keynyn [HA20]. Khaldun [Tek22]. Kinasthesia [Lan24]. Kingdoms [O’M25]. Kingston [Hem22]. Kleinwächter [Los25]. Knapp [Tom25]. Knittel [Wal25a]. Knowledge [Dur23, Goo25, Loc24, Pal21, Jon23, Led22, Lon20, Mat22, RFV25, Ant24, Mar25a, Ros25]. Kong [Tim20, Hua21, Wil21b]. Korea [Hyu24]. Korenjak [Has24]. Koslow [Fis24]. Kostas [Dur23]. Kowal [Far25]. Kraemer [Bie25]. Krige [Mar25a].

L [Bud25, Rob25]. Lab [Bet24, Led21]. Labatut [Ste25a]. laboratories [Bur22]. Laboratory [Dua21, Har21, Wyn21, Led21]. Labour [Dha24]. Lachlan [Arn23]. Lady [Leh23]. Land [Hin24]. Landscape [Des21, Mac22, Chu23]. language [Gor22]. Lanka [SB25]. Lanzillo [Dha24]. Late [Roc24, Bil20, Yal22]. Latin [Has24]. Launius [Kat20]. Laura [MWH25, Wal25b]. Law [Kor21]. Lawrence [Koh23, Led21]. Lawrence-Mackey [Koh23]. lay [Bol23]. leadership [YJ21]. leading [Low21]. League [For23]. Leapfrogging [Ma25a, Ma25b]. Learned [Ken25]. Lebanon [Bur22]. Lebens [Sto23]. Lecture [Sec22, Bar24]. Lee [Gor22]. Leeuwenhoek [tH23]. Lefèvre [Ros20]. Legacy [Sie25b, Med23]. Leiden [Off25, Wal25a]. Leigh [Sie25b]. Leoni [Wal25a]. Leonid [BD20]. Leonie [Spa25]. Levant [Mey22]. Level [O’M25, Wid23]. Levine [McC25]. Liba [Sia21]. Libby [Düm25]. Library [Mar20]. Lies [Hig24]. Life [Bel25b, Cro24b, Gue24, Leh23, Ree21, Wil20, Bro25b, Hop22, Far23, Sle24]. Light [Hem22]. Lighteners [Tha24]. Lightman [Jac21]. like [Dis24, Hen24]. Limits [Bet24]. Lina [Sta24b]. Linda [Kuc25]. linguistic [Gor22]. Linnaeus [Gue24]. Linthorst [Hol23b]. Liquidator [McM20]. Lisa [Fis24]. literary [KK22]. Literature [Has24]. Liu [Mil21]. live [Cow24]. Livermore [Led21]. Liverpool [Guf21]. Lives [Win23a, Sie25a, Hol23a]. Livestock [Tom25]. Living [Hem22]. Loath [Ree25]. local [Ano21h, Wil21a]. locusts [Wor22]. London [Amb21b, Amb24, And24, Ant24, Bee21, Bie25, Bow22, Bro24, DGJQ24, Dun22, Dur23, Eic21, Exl23, Fig21, Har21, Har25, Hel20, Hol23a, Hol20, Hol21, HA20, Jag25, Jam21, Jor23, Koh23, Lan24, Mac20, Mey20, Mil21, Mox21, Ran23, Smi20, Smi24, Tal21, Tim20, Tur21, Vir25, Web24b, Wil20, Win23a, Des21, Kem23, Law21a, Law23a, Law23b, Pfe21]. Long [Cur24, Far21, Gro24, Hay21b]. lore [KW21]. Lorei [Dee20]. Lorek [Van24]. Lorraine [Goo20]. Lothar [Ram24]. Lowengard [Los25]. Lüders [McC23]. Lukas [Bow22, Vir25]. Luminous [Hem22]. lunar [Bel23]. Lutheran [Car24]. Lüttge [Sto23]. lviii [Bro20]. Lydia [vD23]. Lynn [Tha24].

M [Coz20, Dun22, Gro24, Law21b, Mey20, Mox21, Sec21a, Sec21b, Tha24, Vir25, Whi24]. MA [Amb21a, BR23, Bow22, Dun24, Far23, Goo25, Gro23, Gro24, Har21, Jon21, McC25, Mül24, Pal21, Rob25, Sle24, Sta24a, Ste22]. MacCord [Esp25]. Machine [Bol24]. Machines [And24, Tom25]. Mackenzie [Tar24]. Mackey [Koh23]. Macmillan [Exl23, Har25]. magic [Raj24, Tar22]. Mahoney [Kor21]. Maillet [Mat25]. maintenance [Kir20]. Making [Bla20, Bow22, Cho25, Dun24, Fre25, KW21, Käl22, SB25, Bad24, Gre25, Kuc25, Ros20, Tom25, Van24]. malacology [Mar22]. Malcolm [Guf21]. Malinowski [Mar22]. Malthus [Sec21a, Sec21b]. Mamidipudi [Goo25]. Mammals [Cur24]. Mamo [MWH25]. Man [Gue24, McG21]. management [Dis24]. Managing [Joh21]. Manchester [Bol25, Hes25b, Kan20, Spa25]. MANIAC [Ste25a]. Manias [Cur24]. Mankind [Web24b]. Manning [Tur21]. Mantle [Bru22]. Many [Sta24a, Sid25a, Sid25b]. Mapping [Mat22, Ols23]. maps [Mat22]. Marci [Lor24]. Marco [Hor23]. Marcos [Coz20]. Margarete [Smi24]. Margery [EW23]. marginalization [Gar24]. Maria [Bud24, Dor25, Mox21, Off25]. Mariam [Bol25]. Marić [Ste22]. Marie [Kuc25, Nia24]. Mario [Mar25a]. Marius [Goo25]. Mark [Amb21a, Ant24, O’M25, Tar22]. Market [Sac21]. Marketplace [Bud24]. Marlon [Bud25]. marriage [Hop22]. Marrs [Sch25b]. Mars [Sle20]. Marsh [Far20]. Martin [Exl23, Koh23, Has24, Tim20]. Marvels [Hin24]. Mary [Ree21]. Mask [Sch25c, SS23]. Mass [Sle20]. Masters [Mar24b]. material [Hee23, Kai22].

Materializing [Mar25b]. Mathematical [Ros20, Shi24]. Mathematician [Wil20]. Mathematics [Dun23, Ken25, MS21, Bah24]. Matlin [Cro24b]. Matter [Sta24b, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, Ano24b, Ano24a, Ano25b, Ano25a, Ano25d, Ano25c]. Matthew [HA20, Bro25a, Gre25, Hay23]. Maurice [Smi21a]. Mauricio [BC25b]. Mauro [Bud25]. Max [Hof24, Sch25a, Gor22]. Max-Planck-Gesellschaft [Hof24, Sch25a]. Maynard [Pie20]. McCain [Dur23]. McDougal [Ran23]. McDougal-Waters [Ran23]. McGill [Bol24, Cho25, Hem22, Hig24]. McGill-Queen [Bol24, Cho25, Hem22, Hig24]. McMillan [Bro20]. MD [Dun25]. me [Val25]. means [Val25]. measurement [Dis24]. Measures [Bah24]. Measuring [Bet24, Hes25b]. Mechanical [Amb21b, Jon23]. Mechanics [Buc21, Fre24, Jal25, Mar25b]. mechanisms [Led22]. Media [Git25, Sle20, HLDS20]. mediatic [Med23]. mediating [Big23]. Medical [Dav21, Har25, Loc24, Mar24b, Aso24, Kan24]. Medicine [Bro24, Han20, Hol22, KK22, Mar24a, Smi21b, ASW21, Mey22, Raj24, Nut24]. Medicines [Cho25, Far21]. Mediocrity [BR23]. Mediterranean [Spi23]. Meghan [Bel25a]. Meike [Wal25a]. Melissa [Han20]. memory [tH23]. Mendelian [McG21]. Mendelism [Sha22]. mensuration [Dis24]. Merbold [Sie25a]. Merchants [Far21]. merm [Mat25]. Merpeople [Eic21]. mesmeric [Raj24]. metaphor [Wes21, Bru22]. Metcalf [Ols23]. meteorologists [Wil21b]. method [Bei21].

Mexico [SD24]. Meyer [Ram24]. MI5 [Kor21]. Michael [Bow22, Tal21, Bro20, Har25, Hel20, HA20, Jam21]. Michel [Fre24, Bol24, Gro23, Hem22]. Michelle [Leh23]. Microbes [Web24b]. microscope [Bei21]. Microscopical [Bei21]. mid [Raj24, Spi20]. mid-nineteenth-century [Raj24]. mid-twentieth [Spi20]. Middle [MB22]. Midlife [Jon24, Jon25]. Miescher [HS21]. Mighty [And24]. Migration [Chu23]. Mikkel [Hol23a]. Milam [Goo20]. Milena [HA20, Amb24]. Mileva [Ste22]. Millon [Mat25]. Milo [BR23]. Mind [Git25, Law21b, Ber21]. Minds [Kad24]. Minneapolis [Bet24]. Minnesota [Bet24]. Minster [Ree21]. Miracle [Off25]. miracles [Ros23]. mirages [Ame20]. missionnaire [Mat25]. Mitchell [Sch25a]. Mitton [Far23]. Mobile [McC25]. Mobility [Chu23]. Model [Hol20]. Modeling [BC25b]. Modern [Bah24, Bee21, Bel25a, Bie25, Bla20, Cho25, Coo24, Dun24, Git25, Hau24, Hun23, Nia24, Roc24, Tar22, Wal25b, vD23, Bar24, Bie20, KW21, Mat22, MS21, Har25, Ram24]. Moderne [Ram24]. Modernity [Han20, Sie25b]. Mody [Gro24]. Mohan [Jag25]. Molecular [Aru25, Cro24b, Gro23, Dun24]. monsoon [Car21]. Montano [Mox21]. Montreal [Bol24, Cho25, Hem22, Hig24, GB23]. Moon [Kat20]. Morag [Bee21]. Morange [Gro23]. Morefield [Fre25]. Moretta [Kor21]. Morin [Gar24]. Mørk [Ran23]. Morley [Fig21]. Morris [DGJQ24]. Motamedi [Bol25]. Mount [Bim20]. Moxham [Ran23]. Mr [Mor21a]. Mukharji [BR25]. Multifarious [Mor21a]. Mulvin [Law21a, Law23a, Law23b]. Muqaddimah [Tek22]. Murphy [Web20]. Musée [Mau22]. museumization [SS23]. museums [Boo23b, RFV25, And24]. Musgrave [Mor21a]. Music [Bru24]. my [Boo23b]. mycology [Bol23]. mysteries [Bol23].

N [Web20]. nacional [Man24]. Nagasaki [Kad24]. Nall [Sle20]. Nandini [Cho25]. Nanna [McC23]. Naomi [HA20]. Naramore [Kli24]. Narratives [McC23]. NASA [GB23]. Nasim [Mül24]. National [Amb21a, Amu23, Ito24, Led21, Mar25a, Aso24, Hyu24]. Nationalizing [Amu23]. Natural [Kle20, Mor21a, Mox21, Sto24, Tal21, Kom24, Lon20, Mac22, McA22, Ros23, Big25]. natural-history [Mac22, McA22]. Naturalization [GS21]. Nature [Amu23, Bad24, Bel25b, BR23, Bud25, Cur24, Fig21, GS21, Goo20, Gue24, Hel20, Her24, Los25, Mar25b, Ris20, Tar24, Wal25a, Boo23b, Hol23c, Big25, Tal21]. NC [Chu23, Far25, Tha24]. Neal [Tom25]. necessary [Led22]. necessity [Car21]. Necropolis [Web23]. Needham [Bar23]. Neeraja [Web24a]. Negotiating [Hyu24, Mou23]. negotiations [Bus22, GB23]. Nelson [Hol20]. Neptune [Har21]. Nernst [Bad24]. Neto [Mar25b]. nets [Hen24]. network [Wid23]. networks [KW21, Kai22, Low21]. never [Spi20]. News [Sle20]. Newton [Hen24]. Newtonian [Buc21]. Niche [Bec23]. Nicholas [Ris20]. Nicole [Hol20, Ree25]. nihilists [Khr24]. Nineteenth [Cro24a, Cur24, Gre25, Han20, Hay21b, McC23, Pal21, Ros22, Cow24, Dis24, KK22, Mac22, McA22, Raj24, Val25]. Nineteenth-Century [Cro24a, Gre25, Han20, Cow24, Dis24, Mac22, McA22, Val25]. NJ [Bel25b, Des20, Dun23, Fis24, Goo20, Gue24]. Noah [Ran23]. Noakes [Hun23]. Nobbs [Chu23]. Nobbs-Thiessen [Chu23]. Nobel [Käl22]. noise [Tur23]. nomenclature [dÁ24]. non [Omo24]. non-theological [Omo24]. normalizing [Bec20]. Norman [Bro20]. norms [Käl22, Mou23]. North [Bla20, Chu23, Mar24b, McC23, Van24, Bek20, Hel20]. Northern [Dor25, Mac20]. Nos [Bow23a]. Novel [Des20]. November [Bro20]. Nuclear [Sho22, YJ21, Kad24]. nuclein [HS21]. Numbers [Bah24]. NY [Byf24, Dor25, Sto24, Web23].

object [Bel22]. Observation

[Bie25, Fra21, Nix24]. Observatory [Bel23, Joh21, Gil20]. observing [Kle20]. Occult [Hun23]. occupied [Ito24]. OCD [Ste24]. Oceans [Mey20]. Old [Amb21b]. Omar [Mül24]. one [Ber21, Sta24a]. online [BS20, HLDS20]. ontology [Bel22]. Oosterhoff [Ros20]. open [Bee21, Goo25, Koh23, Ran23, Sch25a, KW21]. open-access [Goo25, Koh23, Ran23, Sch25a]. operationalizing [Nic22]. Oppenheimer [HA20]. Optimizing [Bec20]. Orchestrating [Byf24]. order [Joh21]. Ordering [Los25]. Oregon [Kuc25]. O’reilly [HA20]. Oreskes [HA20]. Organisms [Nut24, Hem22]. Organization [Coz20]. Organized [Gue24]. Origin [Jac21]. original [HS21]. Origins [Buc21, Cro24b, Jon24, Jon25]. Other [Amb21a, Sta24b]. Ottoman [Bil20, KK22, MB22, Sto22, Yal22]. Our [Win23b]. Oviedo [Big25]. Own [Win23b, Ber21]. Ownership [Goo25]. Oxford [Bah24, Fre24, Has24, Jal25, Ken25, Kor21, Mar20, O’M25, Ros20, Ste25b]. ozone [GB23].

P [Wor22]. PA [Bro20, Bud24]. Pacific [Bek20]. Padua [Mar24a, Omo24]. Paduan [Bar24]. Paleontology [Cur24]. Palgrave [Exl23, Har25]. pandemic [RBB^{+}20]. Paola [Amb21b, Hin24]. paper [BR25, BC25b, Dun25, Esp25, Fre25, Mar25a, MWH25, Sta24a]. paperback [Amb21a, Bee21, Bet24, Bim20, Bla20, Bow22, Bow23a, Buc21, Bur24, Cho25, Cou25a, Coz20, Cro25, Dee20, Dha24, Di 23, Dua21, Düm25, Dun24, Dur23, Far25, Fig21, Fis24, Fis25, Goo25, Gro24, Guf21, Hay23, Hel20, Hol20, Hol23b, HA20, HA21, Jon21, Kad24, Koh23, Kuc25, Law21a, Law23a, Law23b, Loc24, Mat25, McC25, McM20, Mül24, Pal21, Pen24, Ris20, Sim22, Sle24, Smi21b, Ste22, Sto24, Tar22, Tim20, Tur21, Van24, Web24b, Win23a, vD23]. paperwork [Kle20]. Papin [Sto21]. Parallax [Sta24a]. Parallels [Web24a]. Paris [Bow23a, Dol21]. Parks [Amu23, Hyu24]. Parliament [Led22]. Parré [Guf21]. partial [AA24]. Past [Dol21, Big23]. Pathways [Ram24]. Patients [Hol23a]. Patrick [Tur21]. Patronage [Gar24]. Paul [Bee21, Düm25]. Paula [Spi23]. Paulo [Mar25b]. PDF [Goo25, Koh23, Sch25a, Ran23]. Pearson [Web23]. Pebbles [Cro25]. peces [Man24]. Peculiar [Bla20]. Penguin [Ste25a]. Pennsylvania [Big25]. pense [Cou25a]. Penser [Bow23a]. People [LD21, Wal25a]. perceptions [Mey22]. Perfection [Tar24]. performance [Ame20]. Performing [Aso24]. period [Bie20]. Periodensystem [Ram24]. Periodicals [Gre25]. Periodicity [Ram24]. persona [SB25]. personas [Gep25]. Peter [DGJQ24, Mac20, Sch25c, Win23b]. Petty [Kom24]. Phage [Aru25]. Pharma [Sac21]. Pharmacy [Spi23, Kan24]. Phenomenological [Jal25]. Philadelphia [Big25]. Philip [Bim20, Ken25, McC25]. Philippe [Di 23]. philosophe [Mat25]. Philosopher [Dun25]. Philosophy [Boo23a, Lan24, Mox21, Tek22]. photographs [Bar23, Mau22]. photography [Nix24]. Physical [Gro24, Ame20]. physician [Kle20, Dun25]. Physics [Ste25b, Ber21, Käl22, Hun23]. physiological [Hes25a]. Physiology [Hol21]. Pickren [Smi23]. pictured [Win24]. Picturing [Bar23, End25]. Pierre [Smi21a, Gar24]. pig [Low21]. Pious [Dha24]. Pipeline [Sch22a]. Pit [Wil20]. Pittsburgh [Bim20, Boo23a, Bro20, Bro25a, Bud24, Cro24a, Cur24, Gre25, Han20, Hes23, Hor23, Iwa25, Lor24, LZ24, McC23, Ris20, Roc24, Sec22, Sie25b, Sle20, Spi23, SD24, Web24a, Win23b]. pivotal [Dav21]. place [Bie20]. plague [Hol23c]. Planck [Hof24, Sch25a]. Planet [Sta24a]. Planetary [Sim22, Bel23]. plant [Kai22]. Plaskett [Mac20]. Playfair [Cam25]. plots [Tur23]. Podgorny [Man24]. Polar [KW21]. Pole [Hel20]. polemics [Gar24]. Policing [Whi24]. Policy [Di 23, Wid23, HA20]. polio [LW24]. political [Wes21]. Politics [Bol25, Bru24, Hol23b, Lan22, LD21, MWH25, Pal21, Sha22, Web23, Gar24, Kan24, Ste24, dA20, dÁ24]. Politik [Hof24]. Pooley [Di 23]. Pop [Hay21b]. population [Bec20]. Portugal [Car24, DUS23]. Portuondo [Mox21]. Post [Dol21, Whi24, LW24, Tur23]. post-punk [Tur23]. Post-revolutionary [Dol21]. Post-Roe [Whi24]. post-socialist [LW24]. Postal [Bei21]. Postwar [Di 23, Mar25a]. potable [Val25]. Potato [LD21]. Power [Ros22, TA23, EW23]. Pp [Amb21a, Amb21b, Amb24, Amu23, And24, Ant24, Arn23, Aru25, Bad24, Bah24, BC25a, Bee21, Bel25a, Bel25b, BR23, BR25, Bet24, Bie25, Big25, Bim20, Bol24, Bol25, Boo23a, Bow22, Bow23a, Bow23b, BC25b, Bro20, Bro24, Bro25a, Bru22, Bru24, Buc21, Bud24, Bud25, Bur24, Byf24, Cam25, Cho25, Chu23, Coo24, Cou25a, Coz20, Cro24a, Cro24b, Cro25, Cur24, Dee20, Des20, Des21, Dha24, Di 23, DGJQ24, Dol21, Dor25, Dua21, Düm25, Dun24, Dun25, Dun22, Dun23, Dur23, Eic21, Esp25, Exl23, Far20, Far21, Far23, Far25, Fig21, Fis24, Fis25, Fra20, Fre25, Fre24, GS21, Git25, Goo25, Goo20, Gre25, Gro23, Gro24, Gue24, Guf21, Hab24, Hal24, Han20, Har21, Har25, Has24, Hau24, Hay21b, Hay23, Hel20, Hem22, Her24, Hes23, Hes25b]. Pp [Hig24, Hin24, Hof24, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, HA21, Hor23, Hun23, Iwa25, Jac21, Jag25, Jal25, Jam21, Jon21, Jon24, Jor23, Kad24, Kan20, Kat20, Ken25, Kli24, Koh23, Kor21, Kri23, Kuc25, Lan22, Lan24, Law21a, Law23a, Law25, Law23b, Led21, Leh23, LD21, Loc24, Lor24, Los25, LZ24, Mac20, Man24, Mar25a, Mar20, Mar24b, Mar25b, Mat25, McC23, McC25, McM20, Mey20, Mid21, Mil21, MWH25, Mor21a, Mox21, Mül24, Nia24, Nut24, O’M25, Off25, Ols23, Pal21, Pen24, Ram24, Ran23, Ree21, Ree25, Ris20, Rob25, Roc24, Ros22, Ros25, Ros20, Sch25a, Sch25b, Sch25c, Sec22, Sha22, Sho22, Sia21, Sie25b, Sim22, Sle20, Sle24, Smi20, Smi21b, Smi23, Smi24, Spa25, Spi23, Sta24a, Sta24b, Ste25a, Ste22, Ste25b]. Pp [Sto23, Sto24, SD24, Tal21, Tar22, Tar24, Tha24, Tim20, Tom25, Tur21, Van24, Vir25, Wal25a, Wal25b, Web20, Web23, Web24a, Web24b, Whi24, Win23a, Win23b, vD23, GS22, Wil20]. practicalities [McA22]. Practice [Ken25, Käl22]. Practices [HA20]. Pratik [GS21]. pratiques [Cou25a]. precarity [Low21]. Predict [Byf24]. Pregnant [Coo24, Whi24]. presence [Big23]. Present [McC25, Big23, Chu23, GS22]. présentée [Mat25]. Presidential [Boo23b]. Press [Amb21a, Amb21b, Amu23, Arn23, Aru25, Bah24, BC25a, Bee21, Bel25a, Bel25b, BR23, BR25, Bet24, Bie25, Big25, Bim20, Bla20, Bol24, Bol25, Boo23a, Bow22, Bow23b, BC25b, Bro20, Bro25a, Bru22, Bru24, Buc21, Bud24, Bur24, Byf24, Cam25, Cho25, Chu23, Coo24, Coz20, Cro24a, Cro24b, Cro25, Cur24, Dee20, Des20, Dha24, Di 23, Dol21, Dor25, Dua21, Düm25, Dun24, Dun25, Dun22, Dun23, Esp25, Far20, Far21, Far23, Far25, Fig21, Fis24, Fis25, Fra20, Fre25, Fre24, GS21, GS22, Git25, Goo25, Goo20, Gre25, Gro23, Gro24, Gue24, Guf21, Hab24, Hal24, Han20, Har21, Has24, Hau24, Hay23, Hem22, Her24, Hes23, Hes25b, Hig24, Hin24, Hol20, Hol21, Hol22, HA20, HA21, Hor23, Hun23, Iwa25, Jal25, Jam21, Jon21]. Press [Jon24, Kad24, Kan20, Kat20, Ken25, Kli24, Koh23, Kor21, Kri23, Kuc25, Lan22, Law21a, Law23a, Law23b, Led21, Leh23, LD21, Loc24, Lor24, LZ24, Mac20, Mar25a, Mar24b, McC23, McC25, Mid21, Mil21, MWH25, Mor21a, Mox21, Mül24, Nut24, O’M25, Ols23, Pal21, Pen24, Ran23, Ree21, Ree25, Ris20, Rob25, Roc24, Ros22, Ros25, Ros20, Sac21, Sch25b, Sch25c, Sec22, Sha22, Sho22, Sia21, Sie25b, Sim22, Sle20, Sle24, Smi21b, Smi23, Spa25, Spi23, Sta24a, Sta24b, Ste22, Ste25b, Sto24, SD24, Tal21, Tar22, Tar24, Tha24, Tom25, Tur21, Van24, Wal25b, Web20, Web23, Web24a, Whi24, Win23b, vD23]. Prevention [MWH25]. primate [Kem23]. Princeton [Bel25b, Des20, Goo20, Gue24]. Print [Ree25, Ros20]. Prints [Coo24]. Priorities [Dav21]. prizes [Käl22]. problem [Ano21h, Hen24, Wil21a]. Production [Hem22, KK22, Lon20]. productivity [Kot23]. profit [Cou25b]. Program [Hab24]. Programming [BD20]. Progress [Dor25]. Progressive [Fis24]. Project [Bow23b, Lia24]. projecting [Som23]. Projects [Ros25]. Projit [BR25]. promise [Ma25a, Ma25b]. Property [Goo25]. Prophet [Tal21, SB25]. proteins [dB20]. Protocol [GB23]. provide [Sto22]. providence [Ros23]. Proxies [AA24, Law21a, Law23a, Law23b]. Prozess [Sch25a]. Psyche [Iwa25]. Psychedelic [Jon21]. Psychiatric [Hol20]. Psychiatry [BC25a, Bla20, Wyn21]. Psychic [Roc24]. Psychics [Hun23]. Psychological [Kad24]. Psychologist [Dun25]. Psychology [Lan24, Ste24, Smi23]. Psychotechnics [Smi24]. Public [Amb21a, Di 23, Fis24, Win23b, Boo23b, Sto22]. Publishing [McM20, Ran23, McG21]. Pugwash [Zai23]. punk [Tur23]. Purchase [Dol21]. python [Cow24].

Qian [GL25]. quality [Val25]. Quantum [Jal25, Mar25b, Fre24]. Quartz [Ste25b]. Queen [Bol24, Cho25, Hem22, Hig24]. Quest [Bol24]. qui [Cou25a]. quotation [Hop22].

R [Cam25, Cro25, Lor24, Mac20]. R. [Ayl21]. R.U.R. [Sle24]. Race [BR25, Bur24, Sha22, Tar24, Kat20]. Rachel [Her24]. Racial [Mar24b]. Radder [Hes23]. Radiation [Kuc25, LZ24]. Radical [Bel25b]. radio [Pie20, Tur23]. radiocarbon [Ker20]. Radium [Bud24, Kir20]. Ragan [O’M25]. railways [Gil20]. Rampling [Dun22]. Ramsey [Boo23a]. Ran [Kad24]. Ranelagh [Leh23]. rarefaction [Hen24]. rattles [Nic22]. raw [Bel23]. Re [MB22]. Re-examining [MB22]. Reaching [Kat20]. Reading [Bie25, Git25]. Reaktion [And24, Des21, Eic21, Hel20, Hol23a, Mey20, Smi20, Tur21, Web24b, Win23a]. Real [Ste22]. Realities [Hin24]. Rebecca [Koh23, Coo24, LD21]. Rebel [Wil20]. rebranding [dB20]. reception [Car24, Khr24, Rob25]. Reconstructing [Hes23]. reductions [Bel23]. Rees [Tur21]. reflections [Big23]. Reform [For23, LW24]. Refugee [Hal24]. Regenerate [Esp25]. Regime [Amb21b]. Regiomontanus [Bar24]. regret [Fra21]. Regulation [Mar25a]. rehabilitation [Kan24]. relativity [Gor22]. Religion [Tar22, Omo24, Yal22]. Reluctant [Ree25]. Remaking [Arn23, tH23]. Remedies [Cho25, Spi23]. remedy [Sta22]. Remnants [Bur24]. Renaissance [Tar24]. Rentetzi [Bud24]. ReOrienting [Bro24]. Representation [BC25b, Hee23]. Representing [Fig21, Tur23]. Repton [Des21]. Republic [Kli24]. republican [Bil20]. reputation [RFV25]. Research [Dav21, Kot23, Web24a, Bro25b, Low21, Hol23b, HLDS20]. residues [Hef24]. Resources [Sto24]. responsibility [Pie20]. Restricted [Sho22]. resuscitation [Pfe21]. rethinking [RBB^{+}20]. Retrospectives [Bil20]. Review [Amb21a, Amb21b, Amb24, Amu23, And24, Ant24, Arn23, Aru25, Bad24, Bah24, BC25a, Bel25a, Bel25b, BR23, BR25, Bet24, Bie25, Big25, Bim20, Bla20, Bol24, Bol25, Boo23a, Bow23a, Bow23b, BC25b, Bro20, Bro24, Bro25a, Bru22, Bru24, Buc21, Bud24, Bud25, Bur24, Byf24, Cam25, Cho25, Chu23, Coo24, Cou25a, Coz20, Cro24a, Cro24b, Cro25, Cur24, Dee20, Des20, Des21, Dha24, Dol21, Dor25, Düm25, Dun24, Dun25, Dun22, Dun23, Dur23, Eic21, Esp25, Exl23, Far20, Far21, Far23, Far25, Fis24, Fis25, Fra20, Fre25, Fre24, GS21, GS22, Git25, Goo25, Goo20, Gre25, Gro23, Gro24, Gue24, Hab24, Hal24, Han20, Har21, Har25, Has24, Hau24, Hay21b, Hay23, Hel20, Hem22, Her24, Hes23, Hes25b, Hig24, Hin24, Hof24, Hol23a, Hol20]. Review [Hol21, Hol22, Hol23b, HA20, Hor23, Hun23, Iwa25, Jac21, Jag25, Jal25, Jam21, Jon21, Jon24, Jor23, Kad24, Kan20, Kat20, Ken25, Kli24, Koh23, Kor21, Kri23, Kuc25, Lan22, Lan24, Law21a, Law23a, Law25, Law23b, Led21, Leh23, LD21, Loc24, Lor24, Los25, LZ24, Mac20, Mar25a, Mar20, Mar24b, Mar25b, Mat25, McC23, McC25, McM20, Mey20, Mid21, Mil21, MWH25, Mor21a, Mox21, Mül24, Nia24, Nut24, O’M25, Off25, Ols23, Pen24, Ran23, Ree21, Ree25, Ris20, Rob25, Roc24, Ros22, Ros25, Ros20, Sac21, Sch25b, Sch25c, Sec22, Sha22, Sho22, Sia21, Sie25b, Sim22, Sle20, Sle24, Smi20, Smi21b, Smi23, Smi24, Spa25, Spi23, Sta24a, Sta24b, Ste25a, Ste25b, Sto23, Sto24, SD24, Tar22, Tar24, Tha24, Tim20, Tom25, Van24, Vir25]. Review [Wal25a, Wal25b, Web20, Web23, Web24a, Web24b, Whi24, Wil20, Win23a, Win23b, SW25, DGJQ24]. Reviews [Bee21, Bow22, Di 23, Dua21, Fig21, Guf21, HA21, Man24, Pal21, Ram24, Sch25a, Ste22, Tal21, Tur21, vD23]. revisited [Law21b]. Revolution [Bow22, Des21, Lan22, Sch25b, Van24, Rev25, Gro23]. Revolutionary [Bel25b, Wes21, Dol21]. Revolutions [Sec21a, Sec21b]. revolving [Hua21]. Rewriting [Bow22]. Riaño [Bur24]. Richard [Hun23, Ros20]. Richards [Kuc25].

Rieppel [Bow22]. Rindzeviciute [Byf24]. Rise [Bru24, Win23b]. Risk [Fra20]. rivalries [DUS23]. river [Dis24]. road [Bec23]. Roads [Bro24]. Robert [Leh23, Bee21, Law21b, Sec21a, Sec21b, Smi23]. Robertson [Sie25b]. Robin [Düm25]. Rochester [Kli24]. rock [Sid25a, Sid25b]. Rocke [DGJQ24, Ram24]. Rocket [Gep25, Sie25a]. Roe [Whi24]. Roger [Kat20, Lan24]. Roland [Cro24a]. role [Boo23b, Low21, Spi20]. Roman [Nut24]. Ronit [Bro24]. Roof [Arn23]. Roos [Nia24]. roots [Bar24]. Rosario [Man24]. Roslin [Low21]. Røstvik [Ran23]. Routledge [Amb24, Ant24, Dur23, Jac21, Jag25, Jor23, Lan24, Law25, Smi24]. Royal [Bel25a, Fig21, Ran23, Bel23, Gil20, Kir20]. Rozwadowski [Mey20]. Rs [Bah24]. Rubin [Far23]. Rule [Kor21]. Rules [Web20]. Ruprecht [Hof24, Sch25a]. Rurik [Bek20]. Rush [Kli24]. Russel [Bel25b, Tal21]. Russia [Khr24]. Russian [Ris20]. Rutgers [Dun23, Fis24].

S [BR23, Cro24b, HA21, Mac20, Web24b]. Sabin [LW24]. Sabine [Kan20]. Saccardo [Bol23]. Sachse [Hof24]. Sadie [Koh23]. safe [Bei21]. sailor [Jon23]. Salk [LW24]. Sally [Han20]. Salomon [Bud25]. Salvaging [Sto24]. Samuel [And24]. Sankaran [Web24a]. Sarabhai [Ma25a, Ma25b]. Sarah [Los25, Wal25a, Kli24]. Satellite [GB23]. Saul [GS22]. savantes [Cou25a]. Scaffold [Fre24]. Scale [Fre25]. Schäfer [Goo25]. Schmidt [Jon24, Jon25]. scholarship [RBB^{+}20]. School [HA24]. Schools [Mar24b]. Schwabe [Cou25a]. Science [Amb21a, Amb21b, Amb24, And24, BEH^{+}24a, Bee21, BR25, Boo23a, Bru24, Bud24, Bur24, Cou25b, Di 23, Exl23, Far25, Fig21, Git25, Har21, Har25, Hay21b, Hem22, Her24, Hes23, Hig24, Hin24, Hol22, Hol23b, Hor23, Kad24, Käl22, Koh23, Kri23, Kuc25, Mar24b, McC23, Ros22, Sec22, Sia21, Smi21b, Smi24, Spa25, Sta24b, Sto24, Tar22, Tim20, Win23b, dA20, Ada23, Ano21h, Bar23, BEH^{+}24b, BH24, Bar25, Bec20, Bek20, Big23, Bil20, Boo23b, Bre25, For23, Gil20, Hee23, Hes25a, Led22, Mor21b, Mou23, MB22, SS23, SW25, Ste24, Tek22, Tho24, Wid23, Wil21a, Yal22, tH23, Arn23, Bec23, Bie25, Bim20, Boo23a, BC25b, Bud25, Byf24, Cro25, Dua21, Dur23, Ito24, Jag25, Kan20, Sia21, TA23, Vir25, Win23b]. Sciences [Hab24, Hun23, Shi24, Bol25, HA24, Ito24]. Scientific [Bad24, Bow22, Dur23, GS22, Gre25, Has24, Hol23b, HA20, McM20, Pal21, Pie20, Ran23, Ree25, Tim20, Wil20, BRSW23, Hes25a, Ito24, Käl22, Kir20, Kot23, Mey22, Wyn21, ASW21, Cro24a, Lan22]. scientist [Zai23]. Scientists [Gro24, Hal24, Sec22]. Scientizing [HA24]. Sclavo [Sta22]. Scott [Hes25a, Hig24]. Scottish [Raj24]. screw [Bre25]. Scribner [Eic21]. scrutiny [Led22]. Sea [Har21, Bek20, KW21]. seamanship [Jon23]. seasonal [Car21]. Sebastián [Bur24]. Secord [Bow23b]. Secrecy [Sho22]. Section [Fra20]. secular [Omo24]. Security [Mar25a, LD20]. Seduced [Bud24]. Seeing [Ber21]. Selection [Tal21]. selenography [dÁ24]. Semi [Web23]. Semi-colonial [Web23]. sensation [Bel22]. sense [Bel22]. Senthil [Bah24]. Sera [Roc24]. Sera-Shriar [Roc24]. Sergei [McM20]. serial [Som23]. sermons [Pfe21]. serum [Sta22]. serve [Som23]. service [Khr24]. Set [Jon21]. Setting [Jon21, Web20]. Seuil [Bow23a]. seventeenth [Gar24, Pfe21, dÁ24]. seventeenth-century [Gar24, Pfe21, dÁ24]. Sex [For23]. Sexual [For23]. Sexualizing [MWH25]. Shanghai [Tim20]. Shank [Buc21]. Shaped [Gre25, McC23, Mor21a]. Shaping [Hes25b, MS21]. sharing [HLDS20]. Sharpe [Ree21]. Shaul [Ste25b]. Sherwood [Boo23b]. Shiach [Bee21]. Shindell [HA20]. shock [Wes21]. Short [Kat20, Ker20]. Showcasing [Ada23, Sie25a]. Shriar [Roc24]. Shuttleworth [Han20]. Siam [Web23]. Siam/ [Web23]. side [GO24]. Sigmund [Sie25a]. Silk [Bro24, Far20]. Silva [Mar25b]. similarity [Nic22]. Simon [Far23, Mar20]. Simona [Wal25a]. Simpson [Ros22]. since [Exl23, Web20]. Singapore [McM20, Tim20]. Sir [Bro25b]. Sister [Leh23]. sixth [Tek22]. Skin [Tha24]. Skins [BR25]. Slavery [Mar24b]. Sleigh [Tur21]. slides [Bei21]. smallpox [Mey22]. Smith [Lan24, Sie25b, Tal21, Pie20]. Smithsonian [Nic22]. Smoother [Cro25]. snakes [Cow24]. snowflakes [Win24]. Sociability [Kir20]. Social [Di 23, Dor25, Sch25b, Wal25b, HLDS20, Kot23, Pen24, Pie20, Amb21a, Sha22]. socialist [Aso24, LW24]. Society [Bah24, Fig21, Hig24, Hol23b, Ran23, SD24, Bel25a, BR23, Di 23, Hes23, Kir20, Pen24, Pfe21]. Sociology [Cro25]. softcover [Man24, Mar25b, Ram24]. soil [Bro25b]. Solly [HA24]. Solovey [Amb21a]. Some [Boo23b]. Sonar [Ste25b]. Sörlin [Düm25]. Soul [Dua21]. Sound [Hab24]. sounds [Dis24]. Sourcebook [Koh23]. South [Bah24, GS22, Sto24, Bur24, Car21, Hef24, Hyu24]. southern [GO24]. Sovereign [Web23]. Sovereignty [Sto24]. Soviet [Ris20, Smi24]. Space [Kat20, Ros22, Gep25, Rev25, Sie25a, SB25, Gep25]. spaceflight [Ma25a, Ma25b]. Spain [Spi23]. Spanish [Mox21]. speaking [Ber21]. Species [Bol25]. specimens [Mau22]. Spectacle [Bow22]. Spectacles [Hes25b]. Spektrum [Ram24]. Spiritualism [Roc24]. spore [Bol23]. Springer [Bad24, Bud25, Los25, Mar25b, Ram24]. Spuren [Sto23]. Squares [Gro24]. Sri [SB25]. stacked [Tur23]. stakeholdership [Sch22a]. Stalin [Hal24]. Stammers [Dol21]. Standard [Web20]. standardization [Mou23]. Standing [Law21a, Law23a, Law23b]. Star [Mac20]. Stark [Hau24]. stars [GO24, Gep25, Sid25a, Sid25b, Sie25a]. State [Kuc25, Hef24, Iqb24, Spi20, Cro24a]. States [Ito24, Di 23, Sho22]. stations [Fra21]. Statistical [Cam25]. statistics [Kom24]. Steinle [Los25]. Stephanie [Pen24]. Sternberg [Smi23]. Steven [Amb24, Jal25, Kle20]. Stig [Web24b]. stimulants [ASW21]. Stoppani [LZ23]. Stories [Bol25]. storms [Hua21]. Story [Cam25, Ste22, McM20]. strains [LW24]. strains-derived [LW24]. strategies [DUS23]. Stratifying [Jon23]. Struggle [Dua21]. Stuart [Ros25]. studies [Bar24, Dee20]. Study [BC25b, Kir20, Kle20]. style [Fre24]. Suárez [BC25b]. subservient [Cou25b]. substance [Kan24]. sugar [Pen24]. Summers [Aru25]. Super [Wid23]. Surface [Tha24]. surgery [Raj24]. surrounding [Bus22]. Surroundings [HA21]. Survey [Kom24]. Susanne [Jon24, Jon25]. Sustainability [Hay21a]. Sutton [Dun25]. Sverker [Düm25]. Swammerdam [Kle20]. Sweden [Bec23]. sweet [Pen24]. Sylloge [Bol23]. Symbols [Hay21b]. system [Ish20, Sch22a, Tur21]. Systematics [Off25]. systems [Mar22].

T [Bel25b, Tal21]. Taipei [Tim20]. Tale [Web24a]. Talmey [Gor22]. Tamayo [Rev25]. Tamborini [Hor23]. Tanja [Los25]. Tarter [Led21]. Taub [Sia21]. taxonomy [Bol23]. Taylor [Boo23b]. Tea [Mil21]. teaching [Ber21, Hee23]. Technical [Rei23]. technique [Rei23]. Technocratic [SD24]. Technology [Dha24, Fra20, Fre25, Hes23, Jag25, Lia24, LZ24, McC25, Nut24, Smi24, Ste25b, SD24, Spi20, TA23]. Technopolitics [Hef24]. Teicher [Sha22]. telegraphic [Ish20]. teleology [Sch22b]. television [Kem23, Pie20]. tell [Val25]. Telliamed [Mat25]. Telling [Hig24]. Temminck [Off25]. term [Dis24]. Terms [Dee20]. territorial [Kom24]. Territories [Wal25a]. Test [Law25, Hes25a]. Tests [Mar25b]. Thailand [Kan24]. Thames [Bow22]. their [Ame20, Car24, Wil21b]. Theodore [Coz20]. theological [Omo24]. Theology [Tal21]. Theorien [Ram24]. Theories [Ram24]. Theory [Iwa25, Ros23, Fre24, Spi20]. therapies [Bec20]. Therapy [LZ24, Sta22]. Thiessen [Chu23]. Thinking [Hab24]. Thomas [Tha24, Ros22, Ros23, Sie25b]. Thompson [Sch25c]. Thorogood [Mar20]. throats [Cow24]. thrust [Cow24]. Thurner [Ant24]. Tibet [Fra21]. ticks [Hol23c]. time [Ish20, Hig24]. Times [Han20, Omo24]. Timothy [Van24]. Tkaczyk [Hab24]. Tlalim [Bro24]. Toby [Mor21a]. Tokyo [Tim20]. Tolerance [BR23]. Tom [Des21, Dol21, Ree21]. tools [GB23, Nut24]. Toronto [Cam25, Hal24, Mac20]. Tour [Hin24]. trace [Ame20]. tradition [Sec25]. Trais [Web23]. Trajectories [Web24a]. trans [Web24b, Sch22a]. Trans-Alaska [Sch22a]. Transformations [Hee23]. Transformed [Bud24]. transits [Gil20]. translated [HS21]. translation [Khr24, Tek22]. transmission [Bei21]. Transnational [Ito24, Tha24, Med23]. Transplanting [Sie25b]. Transportation [Sch25b]. Travel [McC23, ASW21]. treatments [Med23]. Tremoring [Gil20]. Trip [Jon21]. Trocadéro [Mau22]. Tropical [Chu23, ASW21]. troubles [McA22]. Truth [Hay21b]. tuberculosis [Aso24]. Tumor [Web24a]. Tuning [Bru24]. Turing [Law25]. Turkey [Bil20]. Turn [LZ24, Bre25]. Turnhout [Nia24]. Turning [Sta24b]. Tuscaloosa [Tal21]. Twentieth [Bee21, Bro25a, Hor23, Jon21, Kir20, Spi20, Sta22, Wyn21]. Twentieth-Century [Bro25a, Hor23, Kir20, Sta22, Wyn21]. Two [BS20, Web24a, Hua21]. Tycho [Car24]. Tycoons [Bow22]. Tyndall [Bro20]. typhoon [Hua21].

U.S [Mar24b]. UCL [Bee21, Koh23, Ran23]. UK [Led22]. Ulf [Sie25a]. Ultimatum [DUS23]. Uncertainty [LZ24]. Underground [Tim20]. Understand [Bol24]. Understanding [Amb24]. understandings [ASW21]. Undisciplining [Ros25]. UNESCO [Bur24]. Unfriendly [YJ21]. Union [Smi24]. United [Di 23, Sho22, Ito24]. Universal [dÁ24]. University [Amb21b, Amu23, Arn23, Aru25, Bah24, Bel25a, Bel25b, BR23, BR25, Bet24, Bie25, Big25, Bim20, Bla20, Bol24, Bol25, Boo23a, Bow22, Bow23b, BC25b, Bro20, Bro25a, Bru22, Bru24, Buc21, Bud24, Bur24, Byf24, Cam25, Cho25, Chu23, Coo24, Coz20, Cro24a, Cro24b, Cro25, Cur24, Dee20, Des20, Dha24, Di 23, Dol21, Dor25, Dua21, Düm25, Dun25, Dun22, Dun23, Esp25, Far20, Far21, Far25, Fig21, Fis24, Fis25, Fra20, Fre25, Fre24, GS21, GS22, Git25, Goo20, Gre25, Gro23, Gue24, Guf21, Hab24, Hal24, Han20, Har21, Has24, Hau24, Hay23, Hem22, Her24, Hes23, Hes25b, Hig24, Hin24, Hol20, Hol21, Hol22, HA20, HA21, Hor23, Hun23, Iwa25, Jal25, Jam21, Jon24, Kad24, Kan20, Kat20, Ken25, Kli24, Kor21, Kri23]. University [Kuc25, Lan22, Led21, Leh23, LD21, Loc24, Lor24, LZ24, Mac20, Mar25a, Mar20, Mar24b, McC23, Mid21, Mil21, MWH25, Mor21a, Mox21, Nut24, O’M25, Ols23, Pal21, Pen24, Ree25, Ris20, Roc24, Ros22, Ros25, Ros20, Sac21, Sch25b, Sch25c, Sec22, Sha22, Sho22, Sia21, Sie25b, Sim22, Sle20, Smi21b, Smi23, Spa25, Spi23, Sta24a, Sta24b, Ste25b, Sto24, SD24, Tal21, Tar22, Tar24, Tha24, Tom25, Tur21, Van24, Wal25b, Web20, Web23, Web24a, Whi24, Win23b, vD23, Shi24]. Unravelled [Far20]. Unseen [Kuc25]. Urge [Smi21b]. Uses [Bil20]. using [HLDS20]. USSR [BD20, Hal24]. uterus [Bel22]. Utrecht [Hol23b]. Uvarov [Wor22].

V [BC25a, BD20]. vaccine [LW24]. Value [RFV25, Mar22]. Vandenhoeck [Hof24, Sch25a]. Vanessa [Hol21]. Vast [Mey20]. Vaughn [Eic21]. Vera [Nia24, Far23, Ros25]. Verburgt [Vir25]. Vereinigung [Sch25a]. Verlag [Cou25a, Sto23]. Veronica [Bru22]. very [Dis24]. Vestiges [Khr24]. Victoria [Hab24]. Victorian [Mid21, Roc24, ASW21, Gil20, Hes25a, Jac21, SW25]. Victorians [Hes25b]. victory [Hyu24]. victory-over-communism [Hyu24]. vienen [Man24]. Vietnam [Aso24]. Vikram [Ma25a, Ma25b]. Viralisation [Har25]. Viruses [Web24a, Web24a]. Visible [Kuc25, Ste24]. Vision [Sle24]. Visions [SD24, Mat22]. visit [Lia24]. Visual [Bel25a, Hes25b, Mau22, Mül24, Ada23, DUS23, GB23, Med23, TA23]. visualizations [GB23]. Viterbo [Man24]. Vöhringer [Smi24]. Voice [Sec22]. vol [Bro20, Fre24, Sia21]. vols [DGJQ24]. Voltaire [Buc21]. volume [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, Ano24b, Ano24a, Ano25b, Ano25a, Ano25d, Ano25c, Bow23b]. voluntarism [Ros23]. Vos [Spi23]. Voyaging [Bek20].

W [Bim20, Mül24, Sch25b, Van24]. Waddell [Tar22]. Wade [Smi23]. Wale [Gre25]. Wales [Mor21b]. Walker [Her24]. Walking [Smi20]. Wallace [Tal21, Bel25b]. Wallis [Han20]. Wallstein [Sto23]. Wals [Sto23]. Waltham [Sta24a]. Walther [Bad24]. Wanderers [Smi20]. War [Aso24, Dua21, Goo20, Kad24, Kor21, Kri23, Mil21, Ste25b, Kan24, Lan22, Sie25a]. Warde [Düm25]. Wardhaugh [Wil20]. Warm [Dor25]. wars [Omo24]. wartime [Bar23]. was [Spi20]. Waste [Sta24b]. Watching [Nix24]. water [Val25]. Waters [Ran23]. Watt [Guf21]. Wazeck [HA20]. Wealth [Sta24b]. weasels [Hol23c]. weather [Ano21h, Car21, Fra21, Wil21a]. Webster [Nut24]. Wege [Ram24]. welfare [Cow24]. Wellerstein [Sho22]. Wendy [Bla20]. Wheat [Lor24]. Wheeler [Jam21]. White [BR25, Sac21]. Whiteley [Coo24]. Who [Gue24, Mor21a, Kan24, Val25]. Whorf [Gor22]. Wife [Ste22]. wildlife [dB20]. Will [Byf24]. William [Cam25, Jor23, Wil20, Aru25, Bel22, Cou25b, Dun25, GS22, McA22]. Williams [Smi21b]. Williamson [Des21]. Willoughby [Mar24b]. Wills [Koh23]. Wilson [Ber21]. Wimborne [Ree21]. Winner [Mid21]. winter [Bel23]. Wisdom [Nia24]. Wissenschaft [Hof24]. without [Web24b]. Witnessing [Roc24]. Wolf [Exl23, Fra20]. Wolfe [Dua21]. Womack [LZ24]. Woman [Ree21]. Women [Exl23, Smi20, Koh23]. Work [Bol24, Law21a, Law23a, Law23b, McC25, Mou23]. works [Mau22]. World [Arn23, Big25, Bru24, Coz20, Dun24, Far20, McM20, Ols23, Tim20, McA22, Wid23, For23, Mor21a]. Worlds [Sta24a, Sid25a, Sid25b]. Wragge [Fig21]. Wragge-Morley [Fig21].

x [Dua21, Far23]. xi [Bud24, Ram24, Tim20]. xii [Jor23]. xiii [Sec22, Tur21]. xiv [Far20, Hol22, Sha22]. xix [Sia21]. Xuesen [GL25]. xvi [Bad24, Fre24, Mox21, Ros22]. xvii [Bim20, Hem22]. xx [Bud25, Mac20]. xxvii [Har25]. xxviii [Koh23].

Yale [Amb21b, Aru25, Hol22, Jam21, Kat20, Mid21, Mil21, Mor21a]. Yates [Web20]. years [Boo23b]. Yoeli [Bro24]. Yoeli-Tlalim [Bro24]. York [Bur24, Cro25, Hay21b, Jac21, Law25, Shi24, Ste25a]. Young [Law21b, Sec21a, Sec21b]. Youth [Hau24].

Zachary [Far21]. Zealand [Bar25]. Zeldovich [Sta24b]. Zimmerman [Hal24]. Zon [Jac21]. Zone [Hay21b]. Zoo [Kem23]. zoological [RFV25]. zoos [Cow24]. Zuckerman [HA24]. zum [Ram24]. Zwigenberg [Kad24].

References

Arzyutov:2024:PPC

[AA24] Dmitry V. Arzyutov and David G. Anderson. Proxies and partial connections in an anthropologist’s archive. British Journal for the History of Science, 57(3): 389–405, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/proxies-
and-partial-connections-in-
an-anthropologists-archive/
655B67A5B0AC0A5FB83C6ED925FB0F1B.

Adamson:2023:SIA

[Ada23] Matthew Adamson. Showcasing the international atom: the IAEA Bulletin as a visual science diplomacy instrument, 1958–1962. British Journal for the History of Science, 56:205–223, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/showcasing-
the-international-atom-the-
iaea-bulletin-as-a-visual-
science-diplomacy-

instrument-19581962/
74ADFFE477CDA3169B3141A97A1A290E.

Ambler:2021:BRM

[Amb21a] Katherine Ambler. Book review: Mark Solovey, Social Science for What? Battles over Public Funding for the ‘Other Sciences’ at the National Science Foundation. Cambridge, MA: MIT Press, 2020. Pp. 398. ISBN: 978-0-262-53905-0. $50.00 (paperback). British Journal for the History of Science, 54(1): 113–114, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mark-
solovey-social-science-for-
what-battles-over-public-
funding-for-the-other-
sciences-at-the-national-
science-foundation-
cambridge-ma-mit-press-
2020-pp-398-isbn-9780262539050-
5000-paperback/151265806FABF2BEA3F69B0DA4EA3EA3.

Amborg:2021:BRP

[Amb21b] Jens Amborg. Book review: Paola Bertucci, Artisanal Enlightenment: Science and the Mechanical Arts in Old Regime France. New Haven, CT and London: Yale University Press, 2017. Pp. 312. ISBN 978-0-300-22741-3. $40.00 (hardcover). British Journal for the History of Science, 54 (1):130–131, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

org/core/journals/british-
journal-for-the-history-of-
science/article/paola-
bertucci-artisanal-
enlightenment-science-and-
the-mechanical-arts-in-old-
regime-france-new-haven-ct-
and-london-yale-university-
press-2017-pp-312-isbn-
9780300227413-4000-
hardcover/0A5EBBD2A86ACD736BCBB8DFC9F37698.

Ambrosio:2024:BRM

[Amb24] Chiara Ambrosio. Book review: Milena Ivanova and Steven French, The Aesthetics of Science: Beauty, Imagination and Understanding, London: Routledge, 2022. Pp. 224. ISBN 978-1-032-33718-0. £110.00 (hardback). British Journal for the History of Science, 57(1):125–127, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/milena-
ivanova-and-steven-french-
the-aesthetics-of-science-
beauty-imagination-and-
understanding-london-
routledge-2022-pp-224-isbn-
9781032337180-11000-
hardback/47B7F4A2AAC9EB4674BF4D6AA0EE928E.

Amery:2020:ATI

[Ame20] Fiona Amery. ‘An attempt to trace illusions to their physical causes’: atmospheric mirages and the performance of their demystification in the 1820s and 1830s. British Journal for the History of Science, 53(4): 443–467, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/an-attempt-
to-trace-illusions-to-
their-physical-causes-
atmospheric-mirages-and-
the-performance-of-their-
demystification-in-the-
1820s-and-1830s/
B6987E17A024924D6939138D6AA0A01A.

Amuchastegui:2023:BRF

[Amu23] Maria Amuchastegui. Book review: Frederico Freitas, Nationalizing Nature: Iguazu Falls and National Parks at the Brazil– Argentina Border Cambridge: Cambridge University Press, 2021. Pp. 333. ISBN 978-1-108-84483-3. $103.00 (hardcover). British Journal for the History of Science, 56(4): 579–582, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/frederico-
freitas-nationalizing-
nature-iguazu-falls-and-
national-parks-at-the-
brazilargentina-border-
cambridge-cambridge-
university-press-2021-pp-
333-isbn-9781108844833-10300-
hardcover/6FD4245159AA4B561FE5645F62920881.

Anderson:2024:BRS

[And24] Robert G. W. Anderson. Book review: Samuel J.M.M. Alberti, Curious Devices and Mighty Machines: Exploring Science Museums, London: Reaktion Books, 2022. Pp. 272. ISBN 978-1-78914-639-4. £20.00 (hardback). British Journal for the History of Science, 57(2):289–291, June 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/samuel-jmm-alberti-
curious-devices-and-mighty-
machines-exploring-science-
museums-london-reaktion-
books-2022-pp-272-isbn-
9781789146394-2000-
hardback/
D21FFA7326A548EC94E4860909C8ABAF.

Anonymous:2020:BVIb

[Ano20a] Anonymous. BJH volume 53 issue 1 cover and back matter. British Journal for the History of Science, 53(1):b1–b4, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-
issue-1-cover-and-back-
matter/
0058932D8AAB0E578F856ED5A199303F.

Anonymous:2020:BVIa

[Ano20b] Anonymous. BJH volume 53 issue 1 cover and front matter. British Journal for the History of Science, 53(1):f1–f2, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-
issue-1-cover-and-front-
matter/
F40087CC5173FE67ACD0BD426C9E69F2.

Anonymous:2020:BVId

[Ano20c] Anonymous. BJH volume 53 issue 2 cover and back matter. British Journal for the History of Science, 53(2):b1–b4, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-
issue-2-cover-and-back-
matter/
6C9C4EFEF420657B595064B14477FD8C.

Anonymous:2020:BVIc

[Ano20d] Anonymous. BJH volume 53 issue 2 cover and front matter. British Journal for the History of Science, 53(2):f1–f2, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
53-issue-2-cover-and-front-
matter/
34AE32A46E4D651DCBE984A5EC825532.

Anonymous:2020:BVIf

[Ano20e] Anonymous. BJH volume 53 issue 4 cover and back matter. British Journal for the History of Science, 53(4):b1–b4, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-
issue-4-cover-and-back-
matter/
407CA57FCF97156253F3606256A9CAB4.

Anonymous:2020:BVIe

[Ano20f] Anonymous. BJH volume 53 issue 4 cover and front matter. British Journal for the History of Science, 53(4):f1–f2, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-
issue-4-cover-and-front-
matter/
6923463CF99F5B2E58D220DEFFF52B44.

Anonymous:2020:F

[Ano20g] Anonymous. Forum. British Journal for the History of Science, 53(1):??, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/scientific-
broadcasting-as-a-social-

                                        responsibility-john-
                                        maynard-smith-on-radio-and-
                                        television-in-the-1960s-
                                        and-1970s/
                                        6F06499A7EDA8C0282E56DB79BA95F9A.

Anonymous:2021:BVIa

[Ano21a] Anonymous. BJH volume 54 issue 1 cover and front matter. British Journal for the History of Science, 54(1):f1–f2, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-54-
                                        issue-1-cover-and-front-
                                        matter/
                                        455321967C798D30144F23675BD3875F.

Anonymous:2021:BVIc

[Ano21b] Anonymous. BJH volume 54 issue 2 cover and back matter. British Journal for the History of Science, 54(2):b1–b4, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-54-
                                        issue-2-cover-and-back-
                                        matter/
                                        910735495A9A03ECA1D195554274BD34.

Anonymous:2021:BVIb

[Ano21c] Anonymous. BJH volume 54 issue 2 cover and front matter. British Journal for the History of Science, 54(2):f1–f2, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
54-issue-2-cover-and-front-
matter/
4B348B7B1B26AA55D60263B31D4ECC71.

Anonymous:2021:BVIe

[Ano21d] Anonymous. BJH volume 54 issue 3 cover and back matter. British Journal for the History of Science, 54(3):b1– b4, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
54-issue-3-cover-and-back-
matter/
5EF77664B0EE32FC1A801AFA6748EA7D.

Anonymous:2021:BVId

[Ano21e] Anonymous. BJH volume 54 issue 3 cover and front matter. British Journal for the History of Science, 54(3):f1–f2, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-54-
issue-3-cover-and-front-
matter/
B46886B9FCB9DE5B2A4254019517AE56.

Anonymous:2021:BVIg

[Ano21f] Anonymous. BJH volume 54 issue 4 cover and back matter.

British Journal for the History of Science, 54(4):b1–b5, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-54-
issue-4-cover-and-back-
matter/
E7FE61616D71B77A245A4AB3E55F52DC.

Anonymous:2021:BVIf

[Ano21g] Anonymous. BJH volume 54 issue 4 cover and front matter. British Journal for the History of Science, 54(4):f1–f2, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-54-
issue-4-cover-and-front-
matter/
F1A37074075A0BDCC17B8A3449185A9E.

Anonymous:2021:FAA

[Ano21h] Anonymous. Framing Asian atmospheres: imperial weather science and the problem of the local c. 1880–1950. British Journal for the History of Science, 54(3), September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ancient-
lore-with-modern-
appliances-networks-

expertise-and-the-making-
of-the-open-polar-sea-
18511853/
9FB26207EDAE5DBD31EE96D5EC829B0B.

Anonymous:2022:BVIb

[Ano22a] Anonymous. BJH volume 55 issue 1 cover and back matter. British Journal for the History of Science, 55(1):b1–b3, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-
issue-1-cover-and-back-
matter/
57A59E54EC0BF2B8AC08947ACC86C761.

Anonymous:2022:BVIa

[Ano22b] Anonymous. BJH volume 55 issue 1 cover and front matter. British Journal for the History of Science, 55(1):f1–f2, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-
issue-1-cover-and-front-
matter/
E5D8287DB2F729BB84764062C2A7602C.

Anonymous:2022:BVId

[Ano22c] Anonymous. BJH volume 55 issue 2 cover and back matter. British Journal for the History of Science, 55(2):b1–b4, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-55-
                                        issue-2-cover-and-back-
                                        matter/
                                        7C6A3E61393F8710651FBF50B01D318F.

Anonymous:2022:BVIc

[Ano22d] Anonymous. BJH volume 55 issue 2 cover and front matter. British Journal for the History of Science, 55(2):f1–f2, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bjh-volume-
                                        55-issue-2-cover-and-front-
                                        matter/
                                        39C5A56421E11409DB89DA16C25C041A.

Anonymous:2022:BVIf

[Ano22e] Anonymous. BJH volume 55 issue 3 cover and back matter. British Journal for the History of Science, 55(3):b1– b4, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bjh-volume-
                                        55-issue-3-cover-and-back-
                                        matter/
                                        86A1733562C5BAE49FA2AF8D42F5A621.

Anonymous:2022:BVIe

[Ano22f] Anonymous. BJH volume 55 issue 3 cover and front matter. British Journal for the History of Science, 55(3):f1–f2, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-
issue-3-cover-and-front-
matter/
6AF6A22729353181A0AD498CA9E0FF66.

Anonymous:2022:BVIh

[Ano22g] Anonymous. BJH volume 55 issue 4 cover and back matter. British Journal for the History of Science, 55(4):b1–b4, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-
issue-4-cover-and-back-
matter/
0E65CD2BA9FE6AB768DEBD818C0B46F3.

Anonymous:2022:BVIg

[Ano22h] Anonymous. BJH volume 55 issue 4 cover and front matter. British Journal for the History of Science, 55(4):f1–f2, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-
issue-4-cover-and-front-
matter/
01FB4E52063F5DDF376949B3FCB3F4D7.

Anonymous:2023:BVIb

[Ano23a] Anonymous. BJH volume 56 issue 1 cover and back matter. British Journal for the History of Science, 56(1):b1–b2, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-56-
                                        issue-1-cover-and-back-
                                        matter/
                                        95DDB64D05811F6FF802F15B9074E695.

Anonymous:2023:BVIa

[Ano23b] Anonymous. BJH volume 56 issue 1 cover and front matter. British Journal for the History of Science, 56(1):f1–f2, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-56-
                                        issue-1-cover-and-front-
                                        matter/
                                        A9A421918B98C972F43EF9ECD087AD89.

Anonymous:2023:BVId

[Ano23c] Anonymous. BJH volume 56 issue 2 cover and back matter. British Journal for the History of Science, 56:b1–b4, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bjh-volume-

56-issue-2-cover-and-back-
matter/
1BC8E28A19A51FDDCCF91FC78AF9D0B8.

Anonymous:2023:BVIc

[Ano23d] Anonymous. BJH volume 56 issue 2 cover and front matter. British Journal for the History of Science, 56:f1–f2, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
56-issue-2-cover-and-front-
matter/
2582A3954C0F09A1B82A2A3255621B35.

Anonymous:2023:BVIf

[Ano23e] Anonymous. BJH volume 56 issue 3 cover and back matter. British Journal for the History of Science, 56(3):b1– b4, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/bjh-volume-
56-issue-3-cover-and-back-
matter/
491FE04838AED2E13469EE5413B95E93.

Anonymous:2023:BVIe

[Ano23f] Anonymous. BJH volume 56 issue 3 cover and front matter. British Journal for the History of Science, 56(3):f1–f2, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/

                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-56-
                                        issue-3-cover-and-front-
                                        matter/
                                        C012C36C3F76BF0FC4E42C25EB119AFF.

Anonymous:2024:BVIb

[Ano24a] Anonymous. BJH volume 57 issue 1 cover and back matter. British Journal for the History of Science, 57(1):b1–b2, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-57-
                                        issue-1-cover-and-back-
                                        matter/
                                        8505ED0E149FFBE3043186F03AFE3C5A.

Anonymous:2024:BVIa

[Ano24b] Anonymous. BJH volume 57 issue 1 cover and front matter. British Journal for the History of Science, 57(1):f1–f2, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-57-
                                        issue-1-cover-and-front-
                                        matter/
                                        6FC90134915BD891DA86F9D9F6C693B7.

Anonymous:2025:BVIb

[Ano25a] Anonymous. BJH volume 58 issue 1 cover and back matter. British Journal for the History of Science, 58(1):b1–b2, March 2025.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-58-
issue-1-cover-and-back-
matter/
8F8D2898D4FC1BAE4DEA1CAE2E45554D.

Anonymous:2025:BVIa

[Ano25b] Anonymous. BJH volume 58 issue 1 cover and front matter. British Journal for the History of Science, 58(1):f1–f2, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-58-
issue-1-cover-and-front-
matter/
E56A8D59532E0CF6B2BB794DE6DC48B3.

Anonymous:2025:BVId

[Ano25c] Anonymous. BJH volume 58 issue 2 cover and back matter. British Journal for the History of Science, 58(2):b1–b2, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-58-
issue-2-cover-and-back-
matter/
5A1AE4663DDC509BE4741390E15AF8BE.

Anonymous:2025:BVIc

[Ano25d] Anonymous. BJH volume 58 issue 2 cover and front matter. British Journal for the History of Science, 58(2):f1–f2, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bjh-volume-
                                        58-issue-2-cover-and-front-
                                        matter/
                                        7629AF7F7BF204504689ED985C3596C8.

Anthony:2024:BRM

[Ant24]  Patrick Anthony.  Book re-

view: Mark Thurner and Jorge Cañizares–Esguerra (eds), The Invention of Humboldt: On the Geopolitics of Knowledge, London: Routledge, 2023. Pp. 342. ISBN 978-1-032-13916-6. £96.00 (hardback). British Journal for the History of Science, 57(1): 134–136, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/mark-
                                        thurner-and-jorge-
                                        canizaresesguerra-eds-the-
                                        invention-of-humboldt-on-
                                        the-geopolitics-of-
                                        knowledge-london-routledge-
                                        2023-pp-342-isbn-
                                        9781032139166-9600-
                                        hardback/5EC7C5A9A8907461F7A682844980BEF7.

Arnold:2023:BRL

[Arn23] Katherine Arnold. Book review:

Lachlan Fleetwood, Science on the Roof of the World: Empire and the Remaking of the Himalaya. Cambridge: Cambridge University Press, 2022. Pp. 294. ISBN 978-1-009-12311-2. £75.00 (hardback). British Journal for the History of Science, 56(4): 582–583, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/lachlan-
fleetwood-science-on-the-
roof-of-the-world-empire-
and-the-remaking-of-the-
himalaya-cambridge-
cambridge-university-press-
2022-pp-294-isbn-
9781009123112-7500-
hardback/EF437D1372EB122CC05BFD9CBE705A76.

Arun:2025:BRW

[Aru25] Ashitha B. Arun. Book review:

William C. Summers, The American Phage Group: The Founders of Molecular Biology, New Haven, CT: Yale University Press, 2023. Pp. 312. ISBN 978-0-300-26356-

5. $65.00 (hardcover). British Journal for the History of Science, 58(2):369–370, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/william-c-summers-
the-american-phage-group-
the-founders-of-molecular-
biology-new-haven-ct-yale-

                                        university-press-2023-pp-
                                        312-isbn-9780300263565-
                                        6500-hardcover/
                                        0BDBAD0137CECE0A6A7A02BB7D69D01C.

Aso:2024:PNI

[Aso24] Michitake Aso. Performing national independence through medical diplomacy: tuberculosis control and socialist internationalism in Cold War Vietnam. British Journal for the History of Science, 57 (2):205–220, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/performing-
                                        national-independence-
                                        through-medical-diplomacy-
                                        tuberculosis-control-and-
                                        socialist-internationalism-
                                        in-cold-war-vietnam/
                                        8E24CEED5CD91825097D61B7E734A11E.

Armston-Sheret:2021:ATM

[ASW21] Edward Armston-Sheret and Kim Walker. Is alcohol a tropical medicine? Scientific understandings of climate, stimulants and bodies in Victorian and Edwardian tropical travel. British Journal for the History of Science, 54 (4):465–484, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/is-alcohol-
                                        a-tropical-medicine-
                                        scientific-understandings-

of-climate-stimulants-and-
bodies-in-victorian-and-
edwardian-tropical-travel/
0B3E72F2CE3EC4E56D609F62C3546CC4.

Aylward:2021:RFE

[Ayl21] Alex Aylward. R. A. Fisher, eugenics, and the campaign for family allowances in interwar Britain. British Journal for the History of Science, 54(4): 485–505, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ra-fisher-
eugenics-and-the-campaign-
for-family-allowances-in-
interwar-britain/
BD28704590AC49EAD4B6FC8F4471C455.

Badino:2024:BRB

[Bad24] Massimiliano Badino. Book review: Benjamin Johnson, Making Ammonia: Fritz Haber, Walther Nernst, and the Nature of Scientific Discovery, Cham: Springer Nature, 2022. Pp. xvi + 278. ISBN 978-3-030-85531-4. £44.99 (hardcover). British Journal for the History of Science, 57 (2):291–293, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/benjamin-
johnson-making-ammonia-
fritz-haber-walther-nernst-
and-the-nature-of-
scientific-discovery-cham-

                                        springer-nature-2022-pp-
                                        xvi-278-isbn-9783030855314-
                                        4499-hardcover/
                                        5220346A03CD757686920E3613571229.

Bahl:2024:BRD

[Bah24] Christopher D. Bahl. Book review: D. Senthil Babu, Mathematics and Society: Numbers and Measures in Early Modern South India, Oxford: Oxford University Press, 2022. Pp. 384. ISBN 978-81-948316-0-0. Rs. 1895.00 (hardcover). British Journal for the History of Science, 57 (2):293–295, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/d-senthil-
                                        babu-mathematics-and-
                                        society-numbers-and-
                                        measures-in-early-modern-
                                        south-india-oxford-oxford-
                                        university-press-2022-pp-
                                        384-isbn-9788194831600-
                                        189500-hardcover/
                                        552C9F078C111AB035CA28BC91158E76.

Barrett:2023:PCS

[Bar23] Gordon Barrett. Picturing Chinese science: wartime photographs in Joseph Needham’s science diplomacy. British Journal for the History of Science, 56:185–203, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/picturing-

chinese-science-wartime-
photographs-in-joseph-
needhams-science-diplomacy/
27697B43FABBBBFCC089FBB4F76F427E.

Bardi:2024:RPL

[Bar24] Alberto Bardi. Regiomontanus’s Paduan lecture of 1464, the Byzantine intellectual heritage and the Graeco–Arabic roots of astronomical studies in early modern Italy. British Journal for the History of Science, 57(4): 529–544, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
regiomontanuss-paduan-
lecture-of-1464-the-
byzantine-intellectual-
heritage-and-the-
graecoarabic-roots-of-
astronomical-studies-in-
early-modern-italy/D9F627B2C5EC27F6D8CD33B96269957F.

Barton:2025:HSA

[Bar25] Ruth Barton. History of science in Aotearoa New Zealand. British Journal for the History of Science, 58(1):117–126, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/history-of-science-
in-aotearoa-new-zealand/
208FC9B1FFAC432F334862AB2992D13F.

Barlow:2025:BRA

[BC25a] Jacob Barlow and Rachel Cooper. Book review: Allan V. Horwitz, DSM: A History of Psychiatry’s Biblem Baltimore: Johns Hopkins Press, 2021. Pp. 232. ISBN 978-1-4214-4069-9. $37.00 (hardcover). British Journal for the History of Science, 58 (2):353–354, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/allan-v-
horwitz-dsm-a-history-of-
psychiatrys-bible-
baltimore-johns-hopkins-
press-2021-pp-232-isbn-
9781421440699-3700-
hardcover/
73647C84C184292779EFCB68A2961F81.

Brewer:2025:BRM

[BC25b] Matthew Brewer and Matilde Carrera. Book review: Mauricio Suárez, Inference and Representation: A Study in Modeling Science, Chicago: University of Chicago Press, 2024. Pp. 328. ISBN 978-0-226-83004-

9. $35.00 (paper). British Journal for the History of Science, 58 (1):163–164, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mauricio-
suarez-inference-and-
representation-a-study-in-

modeling-science-chicago-
university-of-chicago-
press-2024-pp-328-isbn-
9780226830049-3500-paper/
0DC7A2A175C1EAD2AFF096EB631700EC.

Boldyrev:2020:PUL

[BD20] Ivan Boldyrev and Till Düppel.

Programming the USSR: Leonid V. Kantorovich in context. British Journal for the History of Science, 53(2):255–278, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/programming-the-
ussr-leonid-v-kantorovich-
in-context/
4BF0F0D89079DD94AF595EA25A991299.

Beccalossi:2020:ONP

[Bec20] Chiara Beccalossi. Optimizing and normalizing the population through hormone therapies in Italian science, c. 1926–1950. British Journal for the History of Science, 53(1):67–88, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/optimizing-
and-normalizing-the-
population-through-hormone-
therapies-in-italian-
science-c19261950/
766EBC9D01796500ADA4EA4CE1AFFEA7.

Beckman:2023:NDI

[Bec23] Jenny Beckman. Niche development: the International Foundation for Science and the road to Sweden. British Journal for the History of Science, 56(4): 553–566, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/niche-
development-the-
international-foundation-
for-science-and-the-road-
to-sweden/
F6FA6A9AE66F24725778556423EE284A.

Beeston:2021:BRR

[Bee21] Erin Beeston. Book reviews: Robert Bud, Paul Greenhalgh, Frank James and Morag Shiach (eds.), Being Modern: The Cultural Impact of Science in the Early Twentieth Century, London: UCL Press, 2018. Pp. 438. ISBN 978-1-78735-395-4. £50.00 (hardback). ISBN 978-1-78735-394-7. £30.00 (paperback). ISBN 978-1-78735-393-0 (open access). British Journal for the History of Science, 54(2):237–238, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/robert-bud-paul-
greenhalgh-frank-james-and-
morag-shiach-eds-being-
modern-the-cultural-impact-

of-science-in-the-early-
twentieth-century-london-
ucl-press-2018-pp-438-isbn-
9781787353954-5000-hardback-
isbn-9781787353947-3000-paperback-
isbn-9781787353930-open-access/
6F31B2333BB0A68D8D4A1910FDFF77D0.

Barrett:2024:CCC

[BEH^{+}24a] Gordon Barrett, Claire Edington, Aya Homei, Kate Sullivan de Estrada, and Zuoyue Wang. Concluding conversation: De-centring science diplomacy — corrigendum. British Journal for the History of Science, 57(2):287, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/concluding-
conversation-decentring-
science-diplomacy-
corrigendum/
B20301AEF1985AA674B2A202A176D452.

See [BEH^{+}24b].

Barrett:2024:CCD

[BEH^{+}24b] Gordon Barrett, Claire Edington, Aya Homei, Kate Sullivan de Estrada, and Zuoyue Wang. Concluding conversation: decentring science diplomacy. British Journal for the History of Science, 57(2):273–285, June 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/concluding-

                                        conversation-decentring-
                                        science-diplomacy/
                                        336030B810E483389CDF3C2AAFB1BE4C.

See corrigendum [BEH^{+}24a].

Beiermann:2021:MST

[Bei21] Lea Beiermann. ‘A method for safe transmission’: the microscope slides of the American Postal Microscopical Club. British Journal for the History of Science, 54(4): 403–422, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/method-for-
                                        safe-transmission-the-
                                        microscope-slides-of-the-
                                        american-postal-
                                        microscopical-club/
                                        1269011CF843B0CC6758A38883ED6ABC.

Bekasova:2020:VTF

[Bek20]  Alexandra Bekasova.  Voyag-

ing towards the future: the brig Rurik in the North Pacific and the emerging science of the sea. British Journal for the History of Science, 53(4): 469–495, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/voyaging-
                                        towards-the-future-the-
                                        brig-rurik-in-the-north-
                                        pacific-and-the-emerging-
                                        science-of-the-sea/
                                        52837ED83D7D9069F3B5B61465F0252D.

Bellis:2022:OSE

[Bel22] Richard T. Bellis. ‘The object of sense and experiment’: the ontology of sensation in William Hunter’s investigation of the human gravid uterus. British Journal for the History of Science, 55(2):227–246, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/object-of-
sense-and-experiment-the-
ontology-of-sensation-in-
william-hunters-
investigation-of-the-human-
gravid-uterus/
9C603438D58E747B15C1BFA50C2A533D.

Belteki:2023:WRC

[Bel23]  Daniel Belteki.  The winter

of raw computers: the history of the lunar and planetary reductions of the Royal Observatory, Greenwich. British Journal for the History of Science, 56(1):65–81, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/winter-of-
raw-computers-the-history-
of-the-lunar-and-planetary-
reductions-of-the-royal-
observatory-greenwich/
A71791F9C2C0AB7125E49C7B56DFE5F1.

Bellis:2025:BRM

[Bel25a] Richard T. Bellis. Book review: Meghan C. Doherty, Engraving Accuracy in Early Modern England: Visual Communication and the Royal Society, Amsterdam: Amsterdam University Press, 2022. Pp. 244. ISBN 978-94-6372-106-6. EUR117.00 (hardback). British Journal for the History of Science, 58(1): 187–188, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/meghan-c-
doherty-engraving-accuracy-
in-early-modern-england-
visual-communication-and-
the-royal-society-
amsterdam-amsterdam-
university-press-2022-pp-
244-isbn-9789463721066-11700-
hardback/C903C59EE18DA29C828FEF82FD895F26.

Bellon:2025:BRJ

[Bel25b] Richard Bellon. Book review: James T. Costa, Radical by Nature: The Revolutionary Life of Alfred Russel Wallace, Princeton, NJ: Princeton University Press, 2023. Pp. 552. ISBN 978-0-6-9123-379-6. £35.00 (hardcover). British Journal for the History of Science, 58 (1):137–140, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/james-t-
costa-radical-by-nature-
the-revolutionary-life-of-

alfred-russel-wallace-
princeton-nj-princeton-
university-press-2023-pp-
552-isbn-9780691233796-
3500-hardcover/C16E96DDBCA384F24D4DC88510048656.

Bertozzi:2021:SOO

[Ber21] Eugenio Bertozzi. ‘Seeing with one’s own eyes’ and speaking to the mind: a history of the Wilson cloud chamber in the teaching of physics. British Journal for the History of Science, 54(2):177–193, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/seeing-
with-ones-own-eyes-and-
speaking-to-the-mind-a-
history-of-the-wilson-
cloud-chamber-in-the-
teaching-of-physics/
8B43708F0415BFE3941805F7328D6ADB.

Betzler:2024:BRG

[Bet24] Riana Betzler. Book review: Grant Bollmer, The Affect Lab: The History and Limits of Measuring Emotion, Minneapolis: University of Minnesota Press, 2023. Pp. 290. ISBN 978-1-5179-1546-9. $28.00 (paperback). British Journal for the History of Science, 57(2):295–297, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/grant-

                                        bollmer-the-affect-lab-the-
                                        history-and-limits-of-
                                        measuring-emotion-
                                        minneapolis-university-of-
                                        minnesota-press-2023-pp-
                                        290-isbn-9781517915469-
                                        2800-paperback/EAECAFE4E71BF25F8C82273CA542F165.

Barrett:2024:DHS

[BH24] Gordon Barrett and Aya Homei. Decentring histories of science diplomacy: cases from Asia. British Journal for the History of Science, 57(2):165–173, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/decentring-
                                        histories-of-science-
                                        diplomacy-cases-from-asia/
                                        4D3D0F94A97BBBECE75B236DD2171213.

Bienias:2020:PEG

[Bie20] Barbara Bienias. The place of Edward Gresham’s Astrostereon (1603) in the discussion on cosmology and the Bible in the early modern period. British Journal for the History of Science, 53(4):417–442, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/place-of-edward-
                                        greshams-astrostereon-1603-
                                        in-the-discussion-on-
                                        cosmology-and-the-bible-in-

the-early-modern-period/
4A0BBEA8D46280AC9C28FA0639A8BA2E.

Bienias:2025:BRF

[Bie25] Barbara Bienias. Book review:

Fabian Kraemer, A Centaur in London: Reading and Observation in Early Modern Science, Baltimore: Johns Hopkins University Press, 2023. Pp. 344. ISBN 978-1-4214-4631-8. $60.00 (hardcover). British Journal for the History of Science, 58 (2):375–376, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/fabian-
kraemer-a-centaur-in-
london-reading-and-
observation-in-early-
modern-science-baltimore-
johns-hopkins-university-
press-2023-pp-344-isbn-
9781421446318-6000-
hardcover/E931C09515A280F1C310BC7496909FD7.

Bigg:2023:CSM

[Big23]  Charlotte Bigg.  Communi-

cating science, mediating presence: reflections on the present, past and future of conferencing. British Journal for the History of Science, 56(4):567–577, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/communicating-
science-mediating-presence-

                                        reflections-on-the-present-
                                        past-and-future-of-
                                        conferencing/
                                        E747A8A7BE94485649E8466D279A6C58.

Bigelow:2025:BRE

[Big25] Allison Margaret Bigelow. Book review: Elizabeth Gansen, Natural Designs: Gonzalo Fernández de Oviedo and the Invention of New World Nature, Philadelphia: University of Pennsylvania Press, 2023. Pp. 248. ISBN 978-1-5128-2456-8. $55.00 (hardcover). British Journal for the History of Science, 58(2):347–348, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/elizabeth-gansen-
                                        natural-designs-gonzalo-
                                        fernandez-de-oviedo-and-
                                        the-invention-of-new-world-
                                        nature-philadelphia-
                                        university-of-pennsylvania-
                                        press-2023-pp-248-isbn-
                                        9781512824568-5500-hardcover/
                                        EDC9D73B961ED87567C4D120C3E056D0.

Bilgili:2020:RUH

[Bil20] Alper Bilgili. Retrospectives: Uses of history of science in the late Ottoman Empire and early republican Turkey. British Journal for the History of Science, 53 (1):109–117, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-

science/article/
retrospectives-uses-of-
history-of-science-in-the-
late-ottoman-empire-and-
early-republican-turkey/
FFDECA807EDF80ED84E638B221B19D54.

Bimm:2020:BRP

[Bim20] Jordan Bimm. Book review:

Philip W. Clements, Science in an Extreme Environment: The 1963 American Mount Everest Expedition. Pittsburgh: University of Pittsburgh Press, 2018. Pp. xvii + 269. ISBN 978-0-8229-4511-6. $39.95 (paperback). British Journal for the History of Science, 53(1):121–123, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/philip-w-clements-
science-in-an-extreme-
environment-the-1963-
american-mount-everest-
expedition-pittsburgh-
university-of-pittsburgh-
press-2018-pp-xvii-269-
isbn-9780822945116-3995-paperback/
756229A3BEE4019305B76C757EE1A605.

Blakley:2020:BRW

[Bla20] Christopher M. Blakley. Book review: Wendy Gonaver, The Peculiar Institution and the Making of Modern Psychiatry, 1840–1880. Chapel Hill: University of North Carolina Press, 2019. ISBN 978-1-4696-4844-6. $32.95 (paperback). British Journal for the History of Science, 53(4):596–597, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/wendy-gonaver-the-
                                        peculiar-institution-and-
                                        the-making-of-modern-
                                        psychiatry-18401880-chapel-
                                        hill-university-of-north-
                                        carolina-press-2019-isbn-
                                        9781469648446-3295-
                                        paperback/5AB051FF8ECEB727321707156A0715E6.

Bolman:2023:WML

[Bol23] Brad Bolman. What mysteries lay in spore: taxonomy, data, and the internationalization of mycology in Saccardo’s Sylloge Fungorum. British Journal for the History of Science, 56(3):369–390, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/what-mysteries-lay-
                                        in-spore-taxonomy-data-and-
                                        the-internationalization-
                                        of-mycology-in-saccardos-
                                        sylloge-fungorum/
                                        621C128689D7D3A31788AFEB40C39BF3.

Bolman:2024:BRM

[Bol24]  Brad  Bolman.

Book review: Michel Anctil, Animal as Machine: The Quest to Understand How Animals Work and Adapt, Montreal: McGill-Queen’s University Press, 2022. Pp. 334. ISBN 978-0-2280-1053-1. CS$49.95 (cloth). British Journal for the History of Science, 57(2):297–299, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michel-
anctil-animal-as-machine-
the-quest-to-understand-
how-animals-work-and-adapt-
montreal-mcgillqueens-
university-press-2022-pp-
334-isbn-9780228010531-
cs4995-cloth/4F37A527BFB8B520CA96B2662ECF0819.

Bolman:2025:BRM

[Bol25] Brad Bolman. Book review:

Mariam Motamedi Fraser, Dog Politics: Species Stories and the Animal Sciences, Manchester: Manchester University Press, 2024. Pp. 288. ISBN 978-1-5261-7480-2. £85.00 (hardcover). British Journal for the History of Science, 58 (2):339–342, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mariam-
motamedi-fraser-dog-
politics-species-stories-
and-the-animal-sciences-
manchester-manchester-
university-press-2024-pp-
288-isbn-9781526174802-
8500-hardcover/09730D3F170C315F5A60B9B1C3D42DCE.

Boon:2023:BRG

[Boo23a] Tim Boon. Book review: Grant Ramsey and Andreas De Block (eds), The Dynamics of Science: Computational Frontiers in History and Philosophy of Science. Pittsburgh: University of Pittsburgh Press, 2022. Pp. 308. ISBN 978-0-8229-4737-0. $60.00 (hardcover). British Journal for the History of Science, 56(4): 584–585, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/grant-
ramsey-and-andreas-de-
block-eds-the-dynamics-of-
science-computational-
frontiers-in-history-and-
philosophy-of-science-
pittsburgh-university-of-
pittsburgh-press-2022-pp-
308-isbn-9780822947370-6000-
hardcover/8331671B7B0B20EB4058E51DEB780209.

Boon:2023:PAS

[Boo23b] Tim Boon. Presidential address: ‘Some years of cudgelling my brains about the nature and function of science museums’: Frank Sherwood Taylor and the public role of the history of science. British Journal for the History of Science, 56(3):283– 307, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
presidential-address-some-
years-of-cudgelling-my-
brains-about-the-nature-

and-function-of-science-
museums-frank-sherwood-
taylor-and-the-public-role-
of-the-history-of-science/
0904E615DFB0335C3DD1FA3104FF6D14.

Bowler:2022:BRL

[Bow22] Peter J. Bowler. Book reviews:

Lukas Rieppel, Assembling the Dinosaur: Fossil Hunters, Tycoons, and the Making of a Spectacle, Cambridge, MA: Harvard University Press, 2019. Pp. 325. ISBN 978-0-6747-3758-7. £23.95 (hardback). — Michael J. Benton, The Dinosaurs: How a Scientific Revolution Is Rewriting History, London: Thames and Hudson, 2020. Pp. 320. ISBN 978-0-5000-5200-6. £10.99 (paperback). British Journal for the History of Science, 55(4): 529–531, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/lukas-
rieppel-assembling-the-
dinosaur-fossil-hunters-
tycoons-and-the-making-of-
a-spectacle-cambridge-ma-
harvard-university-press-
2019-pp-325-isbn-
9780674737587-2395-
hardback-michael-j-benton-the-
dinosaurs-how-a-scientific-revolution-
is-rewriting-history-london-
thames-and-hudson-2020-pp-320-
isbn-9780500052006-1099-paperback/
95B884C81AF2EA08943DCF8F5114EAED.

Bowler:2023:BRC

[Bow23a] Peter J. Bowler. Book review: Claudine Cohen, Nos ancêtres dans les arbres: Penser l’évolution humaine. Paris: Editions du Seuil, 2021. Pp. 319. ISBN 978-2-0211-7599-

8. EUR23.00 (paperback). British Journal for the History of Science, 56(1):128–129, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/claudine-cohen-nos-
                                        ancetres-dans-les-arbres-
                                        penser-levolution-humaine-
                                        paris-editions-du-seuil-
                                        2021-pp-319-isbn-
                                        9782021175998-2300-
                                        paperback/1B9E1DC71BD23A28E7D221B034F6FF6A.

Bowler:2023:BRF

[Bow23b] Peter J. Bowler. Book review: Frederick Burkhardt, James Secord and the editors of the Darwin Correspondence Project (eds.), The Correspondence of Charles Darwin, volume 22, 1874 Cambridge: Cambridge University Press, 2015. Pp. 855. ISBN 978-1-107-08872-6. £105.00 (hardback). British Journal for the History of Science, 56(3):409– 410, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/frederick-

burkhardt-james-secord-and-
the-editors-of-the-darwin-
correspondence-project-eds-
the-correspondence-of-
charles-darwin-volume-22-
1874-cambridge-cambridge-
university-press-2015-pp-855-
isbn-9781107088726-10500-hardback/
BAD36B9D7EA0A808041FCBFA6B0C26D8.

Bowler:2024:EE

[Bow24] Peter J. Bowler. The end of an era. British Journal for the History of Science, 57(1):113–117, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/end-of-an-era/
9C5CAEA1ACC3225117C1C219F15F574E.

Bernhardt-Radu:2023:BRD

[BR23] Stefan Bernhardt-Radu. Book review: Daniel S. Milo, Good Enough: The Tolerance for Mediocrity in Nature and Society. Cambridge, MA: Harvard University Press, 2019. Pp. 310. ISBN 978-0-6745-0462-2. $28.95 (hardback). British Journal for the History of Science, 56(1): 117–119, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/daniel-s-
milo-good-enough-the-
tolerance-for-mediocrity-
in-nature-and-society-
cambridge-ma-harvard-

                                        university-press-2019-pp-
                                        310-isbn-9780674504622-
                                        2895-hardback/
                                        273C7DDDED760004C87A561DF0ADB8E7.

Bernhardt-Radu:2025:BRP

[BR25] Stefan Bernhardt-Radu. Book review: Projit Bihari Mukharji, Brown Skins, White Coats: Race Science in India, 1920–66, Chicago: University of Chicago Press, 2022. Pp. 360. ISBN 978-0-226-82301-

0. $35.00 (paper). British Journal for the History of Science, 58 (1):177–179, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/projit-
                                        bihari-mukharji-brown-
                                        skins-white-coats-race-
                                        science-in-india-192066-
                                        chicago-university-of-
                                        chicago-press-2022-pp-360-
                                        isbn-9780226823010-3500-
                                        paper/DFD58FCC7872CD5D8BED548998C7DE4A.

Bremner:2025:HSG

[Bre25] Lewis Bremner. The history of science and the global turn of the screw. British Journal for the History of Science, 58 (1):127–131, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/history-of-
                                        science-and-the-global-

turn-of-the-screw/
022FE32550598D9BDB0241D28FF91619.

Brock:2020:BRM

[Bro20] William H. Brock. Book review: Michael D. Barton, Janet Browne, Ken Corbett and Norman McMillan (eds.), The Correspondence of John Tyndall, vol. 6: The Correspondence, November 1856-February 1859. Pittsburgh, PA: University of Pittsburgh Press, 2018. Pp. lviii + 537. ISBN 978-0-8229-4533-8. $125.00. (hardback). British Journal for the History of Science, 53(4): 598–599, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michael-d-
barton-janet-browne-ken-
corbett-and-norman-
mcmillan-eds-the-
correspondence-of-john-
tyndall-vol-6-the-
correspondence-november-
1856february-1859-
pittsburgh-pa-university-of-
pittsburgh-press-2018-pp-lviii-
537-isbn-9780822945338-12500-
hardback/6926EABB3ABCCF08F8278192EB75CFF4.

Brown:2024:BRR

[Bro24] Robert Brown. Book review:

Ronit Yoeli-Tlalim, ReOrienting Histories of Medicine: Encounters along the Silk Roads, London: Bloomsbury, 2022. Pp. 256. ISBN 978-1-350-19582-0. £90.00 (hardback). British Journal for the History of Science, 57(1):

146–148, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/ronit-
                                        yoelitlalim-reorienting-
                                        histories-of-medicine-
                                        encounters-along-the-silk-
                                        roads-london-bloomsbury-
                                        2022-pp-256-isbn-
                                        9781350195820-9000-
                                        hardback/
                                        69D0985E41B429294110C2738E959B86.

Brown:2025:BRM

[Bro25a] Tad Brown. Book review: Matthew Holmes, The Graft Hybrid: Challenging Twentieth-Century Genetics, Pittsburgh: University of Pittsburgh Press, 2024. Pp. 320. ISBN 978-0-8229-4793-6. $55.00 (hardcover). British Journal for the History of Science, 58 (2):371–372, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/matthew-
                                        holmes-the-graft-hybrid-
                                        challenging-
                                        twentiethcentury-genetics-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2024-pp-
                                        320-isbn-9780822947936-
                                        5500-hardcover/0634779D8108DB6992FAF4CF61BE993A.

Brown:2025:HGS

[Bro25b] Tad Brown. Humus gnosis: soil fertility, research and funding in the life of Sir Albert Howard. British Journal for the History of Science, 58(1):61–80, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/humus-gnosis-soil-
fertility-research-and-
funding-in-the-life-of-sir-
albert-howard/
94F7392834F3CF0A242FAAA56CF93A33.

Bigg:2023:AGH

[BRSW23] Charlotte Bigg, Jessica Reinisch, Geert Somsen, and Sven Widmalm. The art of gathering: histories of international scientific conferences. British Journal for the History of Science, 56(4):423–433, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/art-of-gathering-
histories-of-international-
scientific-conferences/
69855DA669EB68A587CCD85977C07878.

Bruinsma:2022:BRV

[Bru22]  Mette Bruinsma.  Book re-

view: Veronica Della Dora, The Mantle of the Earth: Genealogies of a Geographical Metaphor. Chicago: The University of Chicago Press, 2021. Pp. 416. ISBN 978-0-226-74129-1. $65.00 (hardback). British Journal for the History of Science, 55 (2):249–250, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/veronica-
                                        della-dora-the-mantle-of-
                                        the-earth-genealogies-of-a-
                                        geographical-metaphor-
                                        chicago-the-university-of-
                                        chicago-press-2021-pp-416-
                                        isbn-9780226741291-6500-
                                        hardback/C896AA3CB8B66787A260726FD3D266FC.

Bruyninckx:2024:BRF

[Bru24] Joeri Bruyninckx. Book review: Fanny Gribenski, Tuning the World: The Rise of 440 Hertz in Music, Science, and Politics, 1859–1955, Chicago: University of Chicago Press, 2023. Pp. 280. ISBN 978-0-226-82326-

3. $55.00 (cloth). British Journal for the History of Science, 57 (1):127–129, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/fanny-
                                        gribenski-tuning-the-world-
                                        the-rise-of-440-hertz-in-
                                        music-science-and-politics-
                                        18591955-chicago-
                                        university-of-chicago-
                                        press-2023-pp-280-isbn-
                                        9780226823263-5500-cloth/
                                        0A00F4BE95C4FB4CF48EEDB5FB9C9D4B.

Boon:2020:TBO

[BS20] Tim Boon and Charlotte Sleigh. Two BSHS online alternatives to conventional conferences. British Journal for the History of Science, 53 (4):553–554, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/two-bshs-
online-alternatives-to-
conventional-conferences/
8149048F6884315513E5231A47EEC67A.

Bucciarelli:2021:BRJ

[Buc21]  Lucia Bucciarelli.  Book re-

view: J. B. Shank, Before Voltaire: The French Origins of ‘Newtonian’ Mechanics, 1680– 1715. Chicago: The University of Chicago Press, 2018. Pp. 464. ISBN: 978-0-226-50929-7. $55.00 (paperback). British Journal for the History of Science, 54(3): 396–398, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jb-shank-
before-voltaire-the-french-
origins-of-newtonian-
mechanics-16801715-chicago-
the-university-of-chicago-
press-2018-pp-464-isbn-
9780226509297-5500-
paperback/6146D818DCB238492D873694768018F8.

Bud:2024:BRM

[Bud24] Robert Bud. Book review: Maria Rentetzi, Seduced by Radium: How Industry Transformed Science in the American Marketplace, Pittsburgh, PA: University of Pittsburgh Press, 2022. Pp. xi + 292. ISBN 978-0-8229-4706-6. $35.00 (hardback). British Journal for the History of Science, 57 (1):119–121, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/maria-
rentetzi-seduced-by-radium-
how-industry-transformed-
science-in-the-american-
marketplace-pittsburgh-pa-
university-of-pittsburgh-
press-2022-pp-xi-292-isbn-
9780822947066-3500-
hardback/241891E696E363ACC1129E72206662B6.

Bud:2025:BRM

[Bud25]  Robert Bud.  Book review:

Mauro L. Condé and Marlon Salomon, eds., Handbook for the Historiography of Science, Cham: Springer Nature, 2023. Pp. xx + 530. ISBN 978-3-031-27509-8. £199.99 (ebook). British Journal for the History of Science, 58(2):383–384, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mauro-l-
conde-and-marlon-salomon-
eds-handbook-for-the-
historiography-of-science-
cham-springer-nature-2023-
pp-xx-530-isbn-
9783031275098-19999-ebook/
7630978B15490BF8D17CFDA4FD89C8CE.

Burton:2022:CGB

[Bur22] Elise K. Burton. Comparative globalizations: building and dismantling genetic laboratories in Lebanon. British Journal for the History of Science, 55(4): 495–513, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
comparative-globalizations-
building-and-dismantling-
genetic-laboratories-in-
lebanon/
5E248827702CC0CC79D148EB7C7E61FF.

Burton:2024:BRS

[Bur24] Elise K. Burton. Book review:

Sebastián Gil–Riaño, The Remnants of Race Science: UNESCO and Economic Development in the Global South, New York: Columbia University Press, 2023. Pp. 392. ISBN 978-0-231-19435-8. £30.00 (paperback). British Journal for the History of Science, 57(3): 489–491, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/sebastian-
gilriano-the-remnants-of-
race-science-unesco-and-
economic-development-in-
the-global-south-new-york-
columbia-university-press-
2023-pp-392-isbn-

                                        9780231194358-3000-paperback/
                                        E1570784223F75BBDFDBA55B3BB96CF0.

Buschmann:2022:CDD

[Bus22] Rainer F. Buschmann. Contested duplicates: disputed negotiations surrounding ethnographic doppelgängers in German New Guinea, 1898– 1914. British Journal for the History of Science, 55(3):297– 318, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/contested-
                                        duplicates-disputed-
                                        negotiations-surrounding-
                                        ethnographic-doppelgangers-
                                        in-german-new-guinea-
                                        18981914/
                                        BCEAE036B3140E806AA6D03309CAA154.

Byford:2024:BRE

[Byf24] Andy Byford. Book review: Egle Rindzeviciute, The Will to Predict: Orchestrating the Future through Science, Ithaca, NY: Cornell University Press, 2023. Pp. 306. ISBN 978-1-5017-6977-1. $56.95 (hardcover). British Journal for the History of Science, 57 (1):155–157, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/egle-
                                        rindzeviciute-the-will-to-
                                        predict-orchestrating-the-
                                        future-through-science-

ithaca-ny-cornell-
university-press-2023-pp-
306-isbn-9781501769771-
5695-hardcover/
DE23B55FBE51DC177AC73F0A8950624E.

Campolo:2025:BRD

[Cam25] Alexander Campolo. Book review: David R. Bellhouse, The Flawed Genius of William Playfair: The Story of the Father of Statistical Graphics, Toronto: University of Toronto Press, 2023. Pp. 356. ISBN 978-1-4875-4503-

1. CA$65.00 (hardcover). British Journal for the History of Science, 58(2):361–362, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/david-r-bellhouse-
the-flawed-genius-of-
william-playfair-the-story-
of-the-father-of-
statistical-graphics-
toronto-university-of-
toronto-press-2023-pp-356-
isbn-9781487545031-ca6500-hardcover/
07B0CAB536F2757ADC87A11DC25ED976.

Carson:2021:AMN

[Car21] Sarah Carson. Anticipating the monsoon: the necessity and impossibility of the seasonal weather forecast for South Asia, 1886– 1953. British Journal for the History of Science, 54(3):305– 325, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/
                                        anticipating-the-monsoon-
                                        the-necessity-and-
                                        impossibility-of-the-
                                        seasonal-weather-forecast-
                                        for-south-asia-18861953/
                                        1FC69001769FA9B58A1CFA7A69E167A2.

Carolino:2024:HDL

[Car24] Luís Miguel Carolino. How did a Lutheran astronomer get converted into a Catholic authority? The Jesuits and their reception of Tycho Brahe in Portugal. British Journal for the History of Science, 57(4): 581–602, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/how-did-a-
                                        lutheran-astronomer-get-
                                        converted-into-a-catholic-
                                        authority-the-jesuits-and-
                                        their-reception-of-tycho-
                                        brahe-in-portugal/
                                        26312C7292E546F1BE04E43190519DF2.

Chowdhury:2025:BRN

[Cho25] Sharmin Jahan Chowdhury. Book review: Nandini Bhattacharya, Disparate Remedies: Making Medicines in Modern India, Montreal: McGill-Queen’s University Press, 2023. Pp. 272. ISBN 978-0-2280-1753-0. CA$47.95 (paperback). British Journal for the History of Science, 58(1): 181–183, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/nandini-
bhattacharya-disparate-
remedies-making-medicines-
in-modern-india-montreal-
mcgillqueens-university-
press-2023-pp-272-isbn-
9780228017530-ca4795-
paperback/B57F605222E699838C6E9056E6B60A4B.

Chu:2023:BRB

[Chu23] Leo Chu. Book review: Ben Nobbs-Thiessen, Landscape of Migration: Mobility and Environmental Change on Bolivia’s Tropical Frontier, 1952 to the Present. Chapel Hill, NC: University of North Carolina Press, 2020. Pp. 342. ISBN 978-1-4696-5609-0. $99.00 (hardback). British Journal for the History of Science, 56(1):124–126, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/ben-nobbsthiessen-
landscape-of-migration-
mobility-and-environmental-
change-on-bolivias-
tropical-frontier-1952-to-
the-present-chapel-hill-nc-
university-of-north-carolina-
press-2020-pp-342-isbn-9781469656090-
9900-hardback/E7278E0324BA70E0631B3FB2D8D48ECC.

Cooley:2024:BRR

[Coo24] Mackenzie Cooley. Book review:

Rebecca Whiteley, Birth Figures:

Early Modern Prints and the Pregnant Body, Chicago: The University of Chicago Press, 2023. Pp. 312. ISBN 978-0-226-82312-

6. $49.00 (cloth). British Journal for the History of Science, 57(3): 506–508, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/rebecca-
whiteley-birth-figures-
early-modern-prints-and-
the-pregnant-body-chicago-
the-university-of-chicago-
press-2023-pp-312-isbn-
9780226823126-4900-cloth/
2D83A7078F6C3BB4D964BB961BA92DD9.

Coueffe:2025:BRJ

[Cou25a] Louise Couëffé. Book review: Jean–François Bert, Le corps qui pense: Une anthropologie historique des pratiques savantes, Basel: Schwabe Verlag, 2023. Pp. 176. ISBN 978-3-7965-4873-4. CHF 28.00 (paperback). British Journal for the History of Science, 58(2):385–386, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/jeanfrancois-bert-
le-corps-qui-pense-une-
anthropologie-historique-
des-pratiques-savantes-
basel-schwabe-verlag-2023-
pp-176-isbn-9783796548734-

chf-2800-paperback/
C223429A89A250A1941C131A2274D65A.

Cournil:2025:SSP

[Cou25b] Mélanie Cournil. Science ‘subservient to profit’? William Jackson Hooker and the first Glasgow Botanic Gardens (1817– 1841). British Journal for the History of Science, 58(1):39–59, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/science-
subservient-to-profit-
william-jackson-hooker-and-
the-first-glasgow-botanic-
gardens-18171841/
4198CE2A02FB515D08B50A97D21D496C.

Cowie:2024:PTW

[Cow24] Helen Cowie. ‘Down pythons’ throats we thrust live goats’: snakes, zoos and animal welfare in nineteenthcentury Britain. British Journal for the History of Science, 57(3): 369–388, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/down-
pythons-throats-we-thrust-
live-goats-snakes-zoos-and-
animal-welfare-in-
nineteenthcentury-britain/
117E6F1146C1EE514DC0378733E79A2B.

Cozzoli:2020:BRM

[Coz20] Daniele Cozzoli. Book review: Marcos Cueto, Theodore M. Brown and Elizabeth Fee, The World Health Organization: A History. Cambridge: Cambridge University Press, 2019. Pp. ix + 373. ISBN 978-1-108-72884-3. £26.99 (paperback). British Journal for the History of Science, 53 (1):126–128, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/marcos-
cueto-theodore-m-brown-and-
elizabeth-fee-the-world-
health-organization-a-
history-cambridge-
cambridge-university-press-
2019-pp-ix-373-isbn-
9781108728843-2699-paperback/
BE3BBB107CCAA90F39AE8C7A84CDB8D8.

Crook:2024:BRR

[Cro24a] Tom Crook. Book review: Roland Jackson, Scientific Advice to the Nineteenth-Century British State, Pittsburgh: University of Pittsburgh Press, 2023. Pp. 464. ISBN 978-0-8229-4790-5. $65.00 (hardcover). British Journal for the History of Science, 57(3):505– 506, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/roland-

jackson-scientific-advice-
to-the-nineteenthcentury-
british-state-pittsburgh-
university-of-pittsburgh-
press-2023-pp-464-isbn-
9780822947905-6500-
hardcover/
CDC992B072D332D8AECA39178F87C54F.

Crowe:2024:BRK

[Cro24b] Nathan Crowe. Book review:

Karl S. Matlin, Crossing the Boundaries of Life: Günter Blobel and the Origins of Molecular Cell Biology Chicago: University of Chicago Press, 2022. Pp. 368. ISBN 978-0-226-81934-

1. $105.00 (cloth). British Journal for the History of Science, 57(2):299–301, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/karl-s-
matlin-crossing-the-
boundaries-of-life-gunter-
blobel-and-the-origins-of-
molecular-cell-biology-
chicago-university-of-
chicago-press-2022-pp-368-
isbn-9780226819341-10500-
cloth/11C56D4271F516BE4EBFC634F3C6B9DB.

Crowson:2025:BRJ

[Cro25] Jaspreet Crowson. Book review:

Jonathan R. Cole, Smoother Pebbles: Essays in the Sociology of Science, New York: Columbia University Press, 2024. Pp. 680. ISBN 978-0-231-21261-8. £35.00 (paperback). British Journal for the History of Science, 58 (2):387–388, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/jonathan-r-
                                        cole-smoother-pebbles-
                                        essays-in-the-sociology-of-
                                        science-new-york-columbia-
                                        university-press-2024-pp-
                                        680-isbn-9780231212618-
                                        3500-paperback/
                                        54278AF82D627C95C858504B50CA2E13.

Currie:2024:BRC

[Cur24] Adrian Currie. Book review: Chris Manias, The Age of Mammals: Nature, Development, and Paleontology in the Long Nineteenth Century, Pittsburgh: University of Pittsburgh Press, 2023. Pp. 488. ISBN 978-0-8229-4780-6. $65.00 (hardcover). British Journal for the History of Science, 57(3):495–497, September 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/chris-manias-the-
                                        age-of-mammals-nature-
                                        development-and-
                                        paleontology-in-the-long-
                                        nineteenth-century-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2023-pp-
                                        488-isbn-9780822947806-6500-
                                        hardcover/A7E33C035B1121B2025A8D11B02EF354.

deAsua:2020:SCP

[dA20] Miguel de Asúa. Science, Catholicism and politics in Argentina (1910–1935). British Journal for the History of Science, 53 (2):139–158, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/science-
catholicism-and-politics-
in-argentina-19101935/
31F7C79840A424B3E2D65A9BA0DD7713.

deAvila:2024:UEP

[dÁ24] Nydia Pineda de Ávila. Universal enough: the politics of nomenclature in seventeenth-century selenography. British Journal for the History of Science, 57(4): 623–643, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/universal-
enough-the-politics-of-
nomenclature-in-
seventeenthcentury-
selenography/
29BF8AC056D80A1ECDDB052D2A2DF100.

Davies:2021:PMR

[Dav21] Stephen M. Davies. Priorities in medical research: elite dynamics in a pivotal episode for British health research. British Journal for the History of Science, 54(2):195–211, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

                                        journal-for-the-history-of-
                                        science/article/priorities-
                                        in-medical-research-elite-
                                        dynamics-in-a-pivotal-
                                        episode-for-british-health-
                                        research/
                                        305F78561AB890CB7BD5CFCAF5F2E751.

deBont:2020:EGP

[dB20] Raf de Bont. Eating game: proteins, international conservation and the rebranding of African wildlife, 1955–1965. British Journal for the History of Science, 53(2):183–205, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/eating-
                                        game-proteins-
                                        international-conservation-
                                        and-the-rebranding-of-
                                        african-wildlife-19551965/
                                        50786E80B0037B1BBEDF9250988DB00C.

Deen:2020:BRL

[Dee20] Rosa Deen. Book review: Lorei Gruen (ed.), Critical Terms for Animal Studies. Chicago: The University of Chicago Press, 2018. Pp. 472. ISBN 978-0-226-35542-9. £32.50/£24.50 (paperback). British Journal for the History of Science, 53 (1):132–134, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/lorei-
                                        gruen-ed-critical-terms-

for-animal-studies-chicago-
the-university-of-chicago-
press-2018-pp-472-isbn-
9780226355429-32502450-
paperback/
68713C71201FDEE538C4B2C00293F23F.

Dessain:2020:BRI

[Des20] H.-F. Dessain. Book review:

Ian Duncan, Human Forms: The Novel in the Age of Evolution. Princeton, NJ: Princeton University Press, 2019. Pp. 304. ISBN 978-0-691-17507-2. £30.00 (hardcover). British Journal for the History of Science, 53 (2):285–287, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ian-duncan-
human-forms-the-novel-in-
the-age-of-evolution-
princeton-nj-princeton-
university-press-2019-pp-
304-isbn-9780691175072-
3000-hardcover/
9C9BBB377412CE54508746DE4BDA1AE3.

Dessain:2021:BRT

[Des21] H.-F. Dessain. Book review:

Tom Williamson, Humphry Repton: Landscape Design in an Age of Revolution London: Reaktion Books, 2020. Pp. 312. ISBN 978-1-78914-299-0. £35.00 (hardcover). British Journal for the History of Science, 54(1): 115–116, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/tom-
                                        williamson-humphry-repton-
                                        landscape-design-in-an-age-
                                        of-revolution-london-
                                        reaktion-books-2020-pp-312-
                                        isbn-9781789142990-3500-
                                        hardcover/
                                        1851102F7BB8358AD72022A2A49A7B7A.

DiMeo:2024:HCR

[DGJQ24] Michelle DiMeo, Andrew Gregory, Frank A. J. L. James, and Viviane Quirke. Horizontal chemistry: [Review]: Peter J.T. Morris and Alan Rocke (gen. eds.), A Cultural History of Chemistry, 6 vols. London: Bloomsbury Academic, 2022. Pp. 1,728. ISBN 978-1-4742-9492-8, £440.00 (hardback). British Journal for the History of Science, 57(3): 467–477, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/horizontal-
                                        chemistry/
                                        0F622B349EE1D32DB32B6A5D24A7742C.

Dharan:2024:BRA

[Dha24] Nikhil Joseph Dharan. Book review: Amanda Lanzillo, Pious Labour: Islam, Artisanship, and Technology in Colonial India, Berkeley: University of California Press, 2024. Pp. 246. ISBN 978-0-520-39857-3. £30.00 (paperback). British Journal for the History of Science, 57 (2):301–303, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/amanda-
lanzillo-pious-labour-
islam-artisanship-and-
technology-in-colonial-
india-berkeley-university-
of-california-press-2024-
pp-246-isbn-9780520398573-
3000-paperback/
36F3C9AD125DD6D0A8700F1060C16C86.

DiCastri:2023:BRP

[Di 23] Theo Di Castri. Book reviews:

Philippe Fontaine and Jefferson D. Pooley, Society on the Edge: Social Science and Public Policy in the Postwar United States. Cambridge: Cambridge University Press, 2020. Pp. 280. ISBN 978-1-1084-8713-9. £74.99 (hardback). ISBN 978-1-1087-3219-2. £26.99 (paperback). British Journal for the History of Science, 56(1):132–134, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/philippe-fontaine-
and-jefferson-d-pooley-
society-on-the-edge-social-
science-and-public-policy-
in-the-postwar-united-
states-cambridge-cambridge-
university-press-2020-pp-
280-isbn-9781108487139-7499-
hardback-isbn-9781108732192-
2699-paperback/40F69300EB1735D15A27A21AFF2D3378.

Dishington:2024:VTM

[Dis24]  Rachel  Dish-

ington. ‘The very term mensuration sounds engineer-like’: measurement and engineering authority in nineteenth-century river management. British Journal for the History of Science, 57 (1):21–41, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/very-term-
                                        mensuration-sounds-
                                        engineerlike-measurement-
                                        and-engineering-authority-
                                        in-nineteenthcentury-river-
                                        management/
                                        3E96E060014B10F4D73B61234406F961.

Doley:2021:BRT

[Dol21] Caitlin Doley. Book review: Tom Stammers, The Purchase of the Past: Collecting Culture in Postrevolutionary Paris c. 1790–1890. Cambridge: Cambridge University Press, 2020. Pp. 370. ISBN: 978-1-108-78126-8. £90.00 (hardback). British Journal for the History of Science, 54(3):398– 399, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/tom-
                                        stammers-the-purchase-of-
                                        the-past-collecting-
                                        culture-in-
                                        postrevolutionary-paris-

c17901890-cambridge-
cambridge-university-press-
2020-pp-370-isbn-
9781108781268-9000-hardback/
E6C3C9352EF1C5D65EE8F1867BF38813.

Doria:2025:BRM

[Dor25] Corinne Doria. Book review:

Maria Cristina Galmarini, Ambassadors of Social Progress: A History of International Blind Activism in the Cold Warm Ithaca, NY: Northern Illinois University Press, 2024. Pp. 302. ISBN 978-1-5017-7377-8. $56.95 (hardcover). British Journal for the History of Science, 58 (2):355–356, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/maria-
cristina-galmarini-
ambassadors-of-social-
progress-a-history-of-
international-blind-
activism-in-the-cold-war-
ithaca-ny-northern-
illinois-university-press-2024-
pp-302-isbn-9781501773778-5695-
hardcover/D2155F1084CFFE04A01E1EF9BB840341.

Duancic:2021:BRA

[Dua21] Vedran Duancić. Book reviews:

Audra Wolfe, Freedom’s Laboratory: The Cold War Struggle for the Soul of Science. Baltimore: Johns Hopkins University Press, 2018. Pp. x + 302. ISBN 978-1-4214-2673-0. $29.95 (hardback). ISBN 978-1-4214-3908-2. $19.95 (paperback). British Journal for the History of Science, 54 (1):120–122, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/audra-
wolfe-freedoms-laboratory-
the-cold-war-struggle-for-
the-soul-of-science-
baltimore-johns-hopkins-
university-press-2018-pp-x-
302-isbn-9781421426730-
2995-hardback-isbn-
9781421439082-1995-paperback/
F3E75245BC3CB663E7B5314B3AC1FFAF.

Dumpelmann:2025:BRP

[Düm25] Sonja Dümpelmann. Book review: Paul Warde, Libby Robin and Sverker Sörlin, The Environment: A History of the Idea, Baltimore: Johns Hopkins University Press, 2021. Pp. 256. ISBN 978-1-4214-4002-6. $27.00 (paperback). British Journal for the History of Science, 58 (1):153–154, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/paul-warde-
libby-robin-and-sverker-
sorlin-the-environment-a-
history-of-the-idea-
baltimore-johns-hopkins-
university-press-2021-pp-
256-isbn-9781421440026-
2700-paperback/
4DFC74F98C8FA446C717AEE5F0D9D3A7.

Dunn:2022:BRJ

[Dun22] Richard Dunn. Book review: Jennifer M. Rampling, The Experimental Fire: Inventing English Alchemy, 1300–1700, Chicago and London: The University of Chicago Press, 2020. Pp. 416. ISBN 978-0-226-71070-9. £28.00/ $35.00 (hardback). British Journal for the History of Science, 55 (1):119–120, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jennifer-m-
rampling-the-experimental-
fire-inventing-english-
alchemy-13001700-chicago-
and-london-the-university-
of-chicago-press-2020-pp-
416-isbn-9780226710709-
28003500-hardback/
4E90AE5EF256E7C9949EAF260A8CAA82.

Dunning:2023:BRA

[Dun23] David E. Dunning. Book review: Andrew Fiss, A History of Communication and Anxiety in the American Mathematics Classroom, New Brunswick, NJ: Rutgers University Press, 2020. Pp. 216. ISBN 978-1-9788-2021-0. $150.00 (cloth). British Journal for the History of Science, 56(3): 411–412, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/andrew-

                                        fiss-a-history-of-
                                        communication-and-anxiety-
                                        in-the-american-
                                        mathematics-classroom-new-
                                        brunswick-nj-rutgers-
                                        university-press-2020-pp-
                                        216-isbn-9781978820210-
                                        15000-cloth/C5321F2F19886C777235023C0FA6A49B.

Duncan:2024:BRC

[Dun24] Katy Duncan. Book review: Catherine Jackson, Molecular World: Making Modern Chemistry, Cambridge, MA: MIT Press, 2023. Pp. 460. ISBN 978-0-262-54554-

9. $75.00 (paperback). British Journal for the History of Science, 57(2):303–304, June 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/catherine-jackson-
                                        molecular-world-making-
                                        modern-chemistry-cambridge-
                                        ma-mit-press-2023-pp-460-
                                        isbn-9780262545549-7500-
                                        paperback/
                                        A3315AF70B54DA622B1336FDF1100BCC.

Dunham:2025:BRE

[Dun25] Jeremy Dunham. Book review: Emma K. Sutton, William James, MD: Philosopher, Psychologist, Physician, Chicago: University of Chicago Press, 2023. Pp. 240. ISBN 978-0-226-82898-

5. $30.00 (paper). British Journal for the History of Science, 58 (1):149–150, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/emma-k-
sutton-william-james-md-
philosopher-psychologist-
physician-chicago-
university-of-chicago-
press-2023-pp-240-isbn-
9780226828985-3000-paper/
199B83DB8681AFF9DB63236412D21811.

Durlo:2023:BRK

[Dur23] Andrea Durlo. Book review:

Kevin McCain and Kostas Kampourakis: What Is Scientific Knowledge? An Introduction to Contemporary Epistemology of Science. London: Routledge, 2019. Pp. 328. ISBN 978-1-1385-7015-

3. £36.99 (paperback). British Journal for the History of Science, 56(1):131–132, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/kevin-mccain-and-
kostas-kampourakis-what-is-
scientific-knowledge-an-
introduction-to-
contemporary-epistemology-
of-science-london-
routledge-2019-pp-328-isbn-9781138570153-
3699-paperback/267DEB1282AB77D7B98B0B344CB16DAA.

Diogo:2023:CDV

[DUS23] Maria Paula Diogo, Paula Urze, and Ana Simões. Cartoon diplomacy: visual strategies, imperial rivalries and the 1890 British Ultimatum to Portugal. British Journal for the History of Science, 56:147–166, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/cartoon-
                                        diplomacy-visual-
                                        strategies-imperial-
                                        rivalries-and-the-1890-
                                        british-ultimatum-to-
                                        portugal/
                                        40313FFA1FE96BC0EEC321285D61D02A.

Eichberg:2021:BRV

[Eic21] Stephanie Eichberg. Book review: Vaughn Scribner, Merpeople: A Human History. London: Reaktion Books, 2020. Pp. 320. ISBN: 978-1-78914-314-0. £20.00 (hardback). British Journal for the History of Science, 54(1): 122–124, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/vaughn-
                                        scribner-merpeople-a-human-
                                        history-london-reaktion-
                                        books-2020-pp-320-isbn-
                                        9781789143140-2000-
                                        hardback/
                                        4A9AFBA084576FC81CED7EBB49EE03F7.

Endersby:2025:PE

[End25] Jim Endersby. Picturing evolution. British Journal for the History of Science, 58(2): 335–337, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/picturing-
evolution/
E68E522D97363B202308B07A8E81B4EB.

Esposito:2025:BRK

[Esp25] Maurizio Esposito. Book review: Kate MacCord, How Does Germline Regenerate?, Chicago: University of Chicago Press, 2024. Pp. 168. ISBN 978-0-226-83051-

3. $26.00 (paper). British Journal for the History of Science, 58(2):373–374, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/kate-
maccord-how-does-germline-
regenerate-chicago-
university-of-chicago-
press-2024-pp-168-isbn-
9780226830513-2600-paper/
53F2622F6ECB60AB4B9A702ECC8F5737.

Eriksen:2023:CFB

[EW23] Christoffer Basse Eriksen and Xinyi Wen. Colouring flowers: books, art, and experiment in the household of Margery and Henry Power. British Journal for the History of Science, 56(1):21–43, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/

                                        article/colouring-flowers-
                                        books-art-and-experiment-
                                        in-the-household-of-
                                        margery-and-henry-power/
                                        6BA23F60CC9962D8045AF81F8B75799D.

Exley:2023:BRC

[Exl23] Grace Exley. Book review: Claire G. Jones, Alison E. Martin and Alexis Wolf (eds), The Palgrave Handbook of Women and Science since 1660. London: Palgrave Macmillan, 2022. Pp. 658. ISBN 978-3-030-78972-5. £149.99 (hardcover). British Journal for the History of Science, 56(4): 586–587, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/claire-g-
                                        jones-alison-e-martin-and-
                                        alexis-wolf-eds-the-
                                        palgrave-handbook-of-women-
                                        and-science-since-1660-
                                        london-palgrave-macmillan-
                                        2022-pp-658-isbn-
                                        9783030789725-14999-
                                        hardcover/19EFB0C2DD6B80D23E3049786F8D349F.

Fara:2020:BRB

[Far20] Patricia Fara. Book review: Ben Marsh, Unravelled Dreams: Silk and the Atlantic World, 1500– 1840. Cambridge: Cambridge University Press, 2020. Pp. xiv + 487. ISBN 978-1-108-41828-7. £29.99 (hardback). British Journal for the History of Science, 53 (4):594–596, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ben-marsh-
unravelled-dreams-silk-and-
the-atlantic-world-
15001840-cambridge-
cambridge-university-press-
2020-pp-xiv-487-isbn-
9781108418287-2999-
hardback/
DBA61DF3CF2542AA430033C8AF332325.

Fara:2021:BRZ

[Far21]  Patricia Fara.  Book

review: Zachary Dorner, Merchants of Medicines: The Commerce and Coercion of Health in Britain’s Long Eighteenth Century. Chicago: The University of Chicago Press, 2020. Pp. 280. ISBN: 978-0-226-70680-1. $50.00 (hardback). British Journal for the History of Science, 54(4): 531–532, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/zachary-
dorner-merchants-of-
medicines-the-commerce-and-
coercion-of-health-in-
britains-long-eighteenth-
century-chicago-the-
university-of-chicago-
press-2020-pp-280-isbn-9780226706801-
5000-hardback/2F5D8423E34CC446A0941566CA1457BD.

Fara:2023:BRJ

[Far23] Patricia Fara. Book review:

Jacqueline Mitton and Simon Mitton, Vera Rubin: A Life.

Cambridge, MA: Belknap Press, 2021. Pp. x + 309. ISBN 978-0-674-91919-8. £23.95 (hardback). British Journal for the History of Science, 56:278–279, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/jacqueline-mitton-
and-simon-mitton-vera-
rubin-a-life-cambridge-ma-
belknap-press-2021-pp-x-
309-isbn-9780674919198-
2395-hardback/
C001FF83EB9177B6468D91F6062CDDB2.

Farley:2025:BRE

[Far25] Simon Farley. Book review: Emma Kowal, Haunting Biology: Science and Indigeneity in Australia, Durham, NC: Duke University Press, 2023. Pp. 248. ISBN 978-1-4780-2537-5. $27.95 (paperback). British Journal for the History of Science, 58 (1):141–142, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/emma-kowal-
haunting-biology-science-
and-indigeneity-in-
australia-durham-nc-duke-
university-press-2023-pp-
248-isbn-9781478025375-
2795-paperback/
154737A3833770C50525EB04D13F6352.

Figueroa:2021:BRA

[Fig21] Javier Poveda Figueroa. Book reviews: Alexander Wragge-Morley, Aesthetic Science: Representing Nature in the Royal Society of London, 1650–1720. Chicago: The University of Chicago Press, 2020. Pp. 272. ISBN: 978-0-226-68086-6. $120.00 (hardback). ISBN: 978-0-226-68072-9. $40.00 (paperback). British Journal for the History of Science, 54(3): 394–396, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/alexander-
wraggemorley-aesthetic-
science-representing-
nature-in-the-royal-
society-of-london-16501720-
chicago-the-university-of-
chicago-press-2020-pp-272-
isbn-9780226680866-12000-
hardback-isbn-9780226680729-
4000-paperback/3CA2611693D97FAF422E9F0BEC55AF6B.

Fischer:2024:BRJ

[Fis24] Suzanne Fischer. Book review:

Jennifer Lisa Koslow, Exhibiting Health: Public Health Displays in the Progressive Era, New Brunswick, NJ: Rutgers University Press, 2020. Pp. 160. ISBN 978-1-9788-0326-8. $33.95 (paperback). British Journal for the History of Science, 57 (1):148–150, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/jennifer-
                                        lisa-koslow-exhibiting-
                                        health-public-health-
                                        displays-in-the-
                                        progressive-era-new-
                                        brunswick-nj-rutgers-
                                        university-press-2020-pp-
                                        160-isbn-9781978803268-
                                        3395-paperback/2F9844DB29A10A1A397192D9560CDF77.

Fisher:2025:BRA

[Fis25] Amy A. Fisher. Book review: Anita Guerrini, Experimenting with Humans and Animals: From Aristotle to CRISPR, 2nd edn., Baltimore: Johns Hopkins University Press, 2022. Pp. 216. ISBN 978-1-4214-4405-5. $28.95 (paperback). British Journal for the History of Science, 58 (2):343–344, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/anita-
                                        guerrini-experimenting-
                                        with-humans-and-animals-
                                        from-aristotle-to-crispr-
                                        2nd-edn-baltimore-johns-
                                        hopkins-university-press-
                                        2022-pp-216-isbn-
                                        9781421444055-2895-
                                        paperback/1E60E9CEEA16B50E9600D81872574D87.

Forster:2023:SSC

[For23] Laura C. Forster. Sex, science and curated community at the World League for Sexual Reform 1929 conference. British Journal for the History of Science, 56(4):469–484, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/sex-science-and-
curated-community-at-the-
world-league-for-sexual-
reform-1929-conference/
B08C19B5A69BBDCAC1A42D0ECC414B3F.

Frampton:2020:BRJ

[Fra20]  Sally Frampton.  Book re-

view: Jacqueline H. Wolf, Cesarean Section: An American History of Risk, Technology, and Consequence. Baltimore: Johns Hopkins University Press, 2018. Pp. 320. ISBN 978-1-4214-2552-8. $49.95 (hardback). British Journal for the History of Science, 53 (1):136–137, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jacqueline-
h-wolf-cesarean-section-an-
american-history-of-risk-
technology-and-consequence-
baltimore-johns-hopkins-
university-press-2018-pp-
320-isbn-9781421425528-
4995-hardback/
B390AD3BFDC91541E7201878B7D1E65B.

Frank:2021:FAO

[Fra21] Mark E. Frank. Frontier atmosphere: observation and regret at Chinese weather stations in Tibet, 1939–1949. British Journal for the History of Science, 54(3): 361–379, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/frontier-
                                        atmosphere-observation-and-
                                        regret-at-chinese-weather-
                                        stations-in-tibet-19391949/
                                        355E8143AB3B26F40222D703084568EA.

French:2024:CHA

[Fre24] Steven French. Charting the hybrid architectural style of quantum theory: [review]: Anthony Duncan and Michel Janssen, Constructing Quantum Mechanics, vol. 1: The Scaffold 1900– 1923, Oxford: Oxford University Press, 2019. Pp. xvi + 487. ISBN 978-0-19-884547-8. £84.00 (hardback). British Journal for the History of Science, 57(3): 485–488, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/charting-
                                        the-hybrid-architectural-
                                        style-of-quantum-theory/
                                        935AD66CCA9AE5E8467EEBC45A0D3E13.

Freeborn:2025:BRH

[Fre25] Alfred Freeborn. Book review: Heidi Morefield, Developing to Scale: Technology and the Making of Global Health, Chicago: University of Chicago Press, 2023. Pp. 232. ISBN 978-0-226-82863-3. $30.00 (paper). British Journal for the History of Science, 58 (1):169–170, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/heidi-
morefield-developing-to-
scale-technology-and-the-
making-of-global-health-
chicago-university-of-
chicago-press-2023-pp-232-
isbn-9780226828633-3000-
paper/
6FBB985F346F2E584CFD8D2E9A9E670A.

Garau:2024:PCP

[Gar24] Rodolfo Garau. Patronage, cultural politics and the marginalization of astrology in seventeenthcentury France: the case of J.- B. Morin and of his polemics with Pierre Gassendi and his circle. British Journal for the History of Science, 57(4):603– 622, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/patronage-
cultural-politics-and-the-
marginalization-of-
astrology-in-
seventeenthcentury-france-
the-case-of-jb-morin-and-
of-his-polemics-with-
pierre-gassendi-and-his-
circle/D368D8F9C4DF03AD51755CE15FB97960.

Grevsmuhl:2023:SIT

[GB23] Sebastian V. Grevsmühl and Régis Briday. Satellite images as tools of visual diplomacy: NASA’s ozone hole visualizations and the Montreal Protocol negotiations. British Journal for the History of Science, 56:247–267, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/satellite-
                                        images-as-tools-of-visual-
                                        diplomacy-nasas-ozone-hole-
                                        visualizations-and-the-
                                        montreal-protocol-
                                        negotiations/
                                        EA86D11886E97CBABCDFA42126184129.

Geppert:2025:RSS

[Gep25] Alexander C. T. Geppert. Rocket stars, space personas and the global Space Age. British Journal for the History of Science, 58(2):195–213, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/rocket-
                                        stars-space-personas-and-
                                        the-global-space-age/
                                        FEEB9911374DE6CEB287B69228D365E5.

Gillin:2020:TTR

[Gil20] Edward J. Gillin. Tremoring transits: railways, the Royal Observatory and the capitalist challenge to Victorian astronomical science. British Journal for the History of Science, 53 (1):1–24, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/tremoring-
transits-railways-the-
royal-observatory-and-the-
capitalist-challenge-to-
victorian-astronomical-
science/
3EA125839341F1C175D83538F3D9BA78.

Gitelman:2025:BRA

[Git25] Lisa Gitelman. Book review: Adrian Johns, ‘The Science of Reading: Information, Media, and Mind in Modern America, Chicago: University of Chicago Press, 2023. Pp. 504. ISBN 978-0-226-82148-

1. $32.50 (cloth). British Journal for the History of Science, 58(2):377–378, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/adrian-
johns-the-science-of-
reading-information-media-
and-mind-in-modern-america-
chicago-university-of-
chicago-press-2023-pp-504-
isbn-9780226821481-3250-
cloth/9BBBA76EE975741DE0EA271E636A7834.

Geppert:2025:CQX

[GL25] Alexander C. T. Geppert and Lu Liu. The celebrification of Qian Xuesen. British Journal for the History of Science, 58 (2):237–260, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/
                                        celebrification-of-qian-
                                        xuesen/
                                        F209877E627E702FA52E8EF79418A379.

Garau:2024:AAS

[GO24] Rodolfo Garau and Pietro Daniel Omodeo. Arguing about the stars on the southern side of the confessional divide. British Journal for the History of Science, 57(4):521–528, December 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/arguing-about-the-
                                        stars-on-the-southern-side-
                                        of-the-confessional-divide/
                                        5C17D1527708150AA7F4143C624FA71F.

Goodrum:2020:BRE

[Goo20] Matthew R. Goodrum. Book review: Erika Lorraine Milam, Creatures of Cain: The Hunt for Human Nature in Cold War America. Princeton, NJ: Princeton University Press, 2019. Pp. 408. ISBN 978-0-691-18188-2. $29.95 (hardcover). British Journal for the History of Science, 53 (4):591–593, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/erika-
lorraine-milam-creatures-
of-cain-the-hunt-for-human-
nature-in-cold-war-america-
princeton-nj-princeton-
university-press-2019-pp-
408-isbn-9780691181882-
2995-hardcover/736A38931501C3A93F0798782C96F6EE.

Gooday:2025:BRD

[Goo25]  Graeme Gooday.  Book re-

view: Dagmar Schäfer, Annapurna Mamidipudi and Marius Buning (eds.), Ownership of Knowledge: Beyond Intellectual Property, Cambridge, MA: MIT Press, 2023. Pp. 394. ISBN 978-0-262-54559-4. $65.00 (paperback); 978-0-262-37464-4. $0.00 (openaccess PDF). British Journal for the History of Science, 58 (2):389–390, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/dagmar-
schafer-annapurna-
mamidipudi-and-marius-
buning-eds-ownership-of-
knowledge-beyond-
intellectual-property-
cambridge-ma-mit-press-
2023-pp-394-isbn-9780262545594-
6500-paperback-9780262374644-
000-openaccess-pdf/159ECCA45592B72B51EC5B6B2F969F68.

Gordin:2022:ELM

[Gor22] Michael D. Gordin. Einsteinian language: Max Talmey, Benjamin Lee Whorf and linguistic relativity. British Journal for the History of Science, 55 (2):145–165, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/
                                        einsteinian-language-max-
                                        talmey-benjamin-lee-whorf-
                                        and-linguistic-relativity/
                                        ECE744AA2067BA1EC96296C2E89DAC18.

Green:2025:BRM

[Gre25]  Leore Joanne Green.  Book

review: Matthew Wale, Making Entomologists: How Periodicals Shaped Scientific Communities in Nineteenth-Century Britain, Pittsburgh: University of Pittsburgh Press, 2022. Pp. 252. ISBN 978-0-8229-4751-6. $55.00 (hardcover). British Journal for the History of Science, 58 (2):379–380, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/matthew-
                                        wale-making-entomologists-
                                        how-periodicals-shaped-
                                        scientific-communities-in-
                                        nineteenthcentury-britain-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2022-pp-
                                        252-isbn-9780822947516-
                                        5500-hardcover/99830F8F6E0C7B04D050F370F48C5240.

Grodwohl:2023:BRM

[Gro23] Jean-Baptiste Grodwohl. Book review: Michel Morange, The Black Box of Biology: A History of the Molecular Revolution. Cambridge, MA: Harvard University Press, 2020. Pp. 528. ISBN 978-0-674-28136-3. £40.95 (hardcover). British Journal for the History of Science, 56(4): 588–589, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michel-
morange-the-black-box-of-
biology-a-history-of-the-
molecular-revolution-
cambridge-ma-harvard-
university-press-2020-pp-
528-isbn-9780674281363-
4095-hardcover/
B2A4C3723C18910BF9BBE1DB6E245F2C.

Gross:2024:BRC

[Gro24] Benjamin Gross. Book review:

Cyrus C. M. Mody, The Squares: US Physical and Engineering Scientists in the Long 1970s, Cambridge, MA: MIT Press, 2023. Pp. 422. ISBN 978-0-262-54361-

3. $65.00 (paperback). British Journal for the History of Science, 57(1):121–123, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/cyrus-cm-mody-the-
squares-us-physical-and-
engineering-scientists-in-
the-long-1970s-cambridge-

                                        ma-mit-press-2023-pp-422-
                                        isbn-9780262543613-6500-
                                        paperback/
                                        0CF21C826A2FB4D2C4830B783A306DF0.

Giovannetti-Singh:2021:BRP

[GS21]  Gianamar  Giovannetti-Singh.

Book review: Pratik Chakrabarti, Inscriptions of Nature: Geology and the Naturalization of Antiquity. Baltimore: Johns Hopkins University Press, 2020. Pp. 280. ISBN: 987-1-4214-3874-0. $54.95 (hardback). British Journal for the History of Science, 54(1): 118–120, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/pratik-
                                        chakrabarti-inscriptions-
                                        of-nature-geology-and-the-
                                        naturalization-of-
                                        antiquity-baltimore-johns-
                                        hopkins-university-press-
                                        2020-pp-280-isbn-
                                        9817421438740-5495-hardback/
                                        DA7541550E65A554C49E11DA3EEFD8A9.

The title and URL have an incorrect ISBN-13 prefix: 987 and 9817 should both be 978. The correct ISBN-13 is 978-1-4214-3874-0 (hardcover).

Giovannetti-Singh:2022:BRW

[GS22] Gianamar Giovannetti-Singh. Book review: William Beinart and Saul Dubow, The Scientific Imagination in South Africa: 1700 to the Present, Cambridge: Cambridge University Press, 2022, pp. 406. ISBN 987-1-1088-3708-8. £64.99 (hardback). British Journal for the History of Science, 55 (1):121–122, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/william-
beinart-and-saul-dubow-the-
scientific-imagination-in-
south-africa-1700-to-the-
present-cambridge-
cambridge-university-press-
2022-pp-406-isbn-
9871108837088-6499-hardback/
19A29A988363D007DAFB5B894027D25F.

The title and URL have an invalid ISBN-13 prefix: 987 should be 978. The correct ISBN-13 values are 978-1-108-83708-8 (hardcover) and 978-1-108-94008-5 (paperback).

Guerrini:2024:BRG

[Gue24] Anita Guerrini. Book review: Gunnar Broberg, The Man Who Organized Nature: The Life of Linnaeus, Princeton, NJ: Princeton University Press, 2023. Pp. 512. ISBN 978-0-691-21342-2. £35.00 (hardcover). British Journal for the History of Science, 57(3): 497–499, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/gunnar-
broberg-the-man-who-

                                        organized-nature-the-life-
                                        of-linnaeus-princeton-nj-
                                        princeton-university-press-
                                        2023-pp-512-isbn-
                                        9780691213422-3500-
                                        hardcover/40C93FC26E400424570CDC788C373A78.

Guffroy:2021:BRM

[Guf21]  Yohann Guffroy.  Book re-

views: Malcolm Dick and Caroline Archer–Parré (eds.), James Watt, 1736–1819: Culture, Innovation, Enlightenment. Liverpool: Liverpool University Press, 2020. Pp. 280. ISBN: 978-1-78962-081-8. £80.00 (hardback). ISBN: 978-1-78962-082-5. £24.99 (paperback). British Journal for the History of Science, 54(3): 400–401, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/malcolm-
                                        dick-and-caroline-
                                        archerparre-eds-james-watt-
                                        17361819-culture-
                                        innovation-enlightenment-
                                        liverpool-liverpool-
                                        university-press-2020-pp-
                                        280-isbn-9781789620818-8000-
                                        hardback-isbn-9781789620825-
                                        2499-paperback/8A972503C086D2856849735E58A5E44D.

Honeybun-Arnolda:2020:BRM

[HA20] Elliot Honeybun-Arnolda. Book review: Michael Oppenheimer, Naomi Oreskes, Dale Jamieson, Keynyn Brysse, Jessica O’reilly, Matthew Shindell and Milena Wazeck, Discerning Experts: The Practices of Scientific Assessment for Environmental Policy. Chicago and London: The University of Chicago Press, 2019. Pp. ix + 281. ISBN 978-0-226-60201-1. $35.00 (paperback). British Journal for the History of Science, 53 (1):128–129, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michael-
oppenheimer-naomi-oreskes-
dale-jamieson-keynyn-
brysse-jessica-oreilly-
matthew-shindell-and-
milena-wazeck-discerning-
experts-the-practices-of-
scientific-assessment-for-environmental-
policy-chicago-and-london-the-
university-of-chicago-press-
2019-pp-ix-281-isbn-9780226602011-
3500-paperback/C0FFD7F2846745895EE20848B8F81957.

Honeybun-Arnolda:2021:BRE

[HA21] Elliott Honeybun-Arnolda. Book reviews: Etienne S. Benson, Surroundings: A History of Environments and Environmentalisms. Chicago: The University of Chicago Press, 2020. Pp. 296. ISBN: 978-0-226-70615-3. $82.50 (hardback). ISBN: 978-0-226-70629-0. $27.50 (paperback). British Journal for the History of Science, 54(3): 392–394, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/etienne-s-
                                        benson-surroundings-a-
                                        history-of-environments-
                                        and-environmentalisms-
                                        chicago-the-university-of-
                                        chicago-press-2020-pp-296-
                                        isbn-9780226706153-8250-
                                        hardback-isbn-9780226706290-
                                        2750-paperback/C33FA7C6C8E158939C3CC68E7F711562.

Honeybun-Arnolda:2024:SES

[HA24] Elliot Honeybun-Arnolda. Scientizing the ‘environment’: Solly Zuckerman and the idea of the School of Environmental Sciences. British Journal for the History of Science, 57(1):99–112, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/scientizing-the-
                                        environment-solly-
                                        zuckerman-and-the-idea-of-
                                        the-school-of-
                                        environmental-sciences/
                                        9A0307D4D7E158EC7332F6482B11443C.

Haberer:2024:BRV

[Hab24] Maximilian Haberer. Book review: Victoria Tkaczyk, Thinking with Sound: A New Program in the Sciences and Humanities around 1900, Chicago: The University of Chicago Press, 2023. Pp. 304. ISBN 978-0-226-82328-

7. $55.00 (cloth). British Journal for the History of Science, 57 (1):129–131, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/victoria-
tkaczyk-thinking-with-
sound-a-new-program-in-the-
sciences-and-humanities-
around-1900-chicago-the-
university-of-chicago-
press-2023-pp-304-isbn-
9780226823287-5500-cloth/
380C6DC3D9988C6DFE2E15E0D053B4B5.

Hall:2024:BRD

[Hal24] Karl Hall. Book review: David Zimmerman, Ensnared between Hitler and Stalin: Refugee Scientists in the USSR, Toronto: University of Toronto Press, 2023. Pp. 376. ISBN 978-1-4875-4365-5. $85.00 (hardcover). British Journal for the History of Science, 57(2):305–307, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/david-
zimmerman-ensnared-between-
hitler-and-stalin-refugee-
scientists-in-the-ussr-
toronto-university-of-
toronto-press-2023-pp-376-
isbn-9781487543655-8500-
hardcover/
2861E0F67F00012B8F3AEC47980B02B8.

Hanley:2020:BRA

[Han20] Anne Hanley. Book review: Amelia Bonea, Melissa Dickson, Sally Shuttleworth and Jennifer Wallis, Anxious Times: Medicine and Modernity in Nineteenth-Century Britain. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 320. ISBN 978-0-8229-4551-2, $50.00 (hardcover). British Journal for the History of Science, 53 (1):119–121, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/amelia-
bonea-melissa-dickson-
sally-shuttleworth-and-
jennifer-wallis-anxious-
times-medicine-and-
modernity-in-
nineteenthcentury-britain-
pittsburgh-university-of-pittsburgh-
press-2019-pp-320-isbn-9780822945512-
5000-hardcover/DA18E80463DFEAE7F3271139BED89B19.

Hardy:2021:BRA

[Har21] Penelope K. Hardy. Book review: Antony Adler, Neptune’s Laboratory: Fantasy, Fear, and Science at Sea Cambridge, MA and London: Harvard University Press, 2019. Pp. 256. ISBN: 978-0-674-97201-8. £31.95 (hardback). British Journal for the History of Science, 54(1): 111–113, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/antony-
adler-neptunes-laboratory-
fantasy-fear-and-science-
at-sea-cambridge-ma-and-

london-harvard-university-
press-2019-pp-256-isbn-
9780674972018-3195-
hardback/9717B6A0406F13136698DA170329DEC0.

Harris:2025:BRM

[Har25] Jim Harris. Book review: Michael Bresalier, Modern Flu: British Medical Science and the Viralisation of Influenza, 1890–1950, London: Palgrave Macmillan, 2023. Pp. xxvii + 458. ISBN 978-1-137-33953-9. £95.00 (hardcover). British Journal for the History of Science, 58(1): 151–152, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michael-
bresalier-modern-flu-
british-medical-science-
and-the-viralisation-of-
influenza-18901950-london-
palgrave-macmillan-2023-pp-
xxvii-458-isbn-
9781137339539-9500-hardcover/
FB4BBB19484278E4A6D30B4828998814.

Haskell:2024:BRM

[Has24] Yasmin Haskell. Book review: Martin Korenjak, Latin Scientific Literature, 1450–1850, Oxford: Oxford University Press, 2023. Pp. 544. ISBN 978-0-19-886605-3. £120.00 (hardback). British Journal for the History of Science, 57(4): 671–673, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.

                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/martin-
                                        korenjak-latin-scientific-
                                        literature-14501850-oxford-
                                        oxford-university-press-
                                        2023-pp-544-isbn-
                                        9780198866053-12000-
                                        hardback/
                                        D085FC546387421000FAA184B9B755F1.

Hau:2024:BRJ

[Hau24]  Michael Hau.  Book review:

James A. Stark, The Cult of Youth: Anti-ageing in Modern Britain, Cambridge: Cambridge University Press, 2020. Pp. 262. ISBN 978-1-108-48415-2. $108.00 (hardcover). British Journal for the History of Science, 57(4): 659–660, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/james-a-
                                        stark-the-cult-of-youth-
                                        antiageing-in-modern-
                                        britain-cambridge-
                                        cambridge-university-press-
                                        2020-pp-262-isbn-
                                        9781108484152-10800-
                                        hardcover/
                                        1831372059AB73F273F2790BEE7A128B.

Hayes:2021:AHC

[Hay21a] Emily Hayes. The Anthropocene & the humanities: From climate change to a new age of sustainability. British Journal for the History of Science, 54(3): 381–385, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
anthropocene-the-
humanities-from-climate-
change-to-a-new-age-of-
sustainability/
DC0FF14BFC0FE24FAD82B2EA5F26D7EF.

Hayes:2021:BRA

[Hay21b] Emily Hayes. Book review: Andrei Pop, A Forest of Symbols: Art, Science, and Truth in the Long Nineteenth Century. New York: Zone Books, 2019. Pp. 320. ISBN 978-1-935408-36-9. £25.00 (hardback). British Journal for the History of Science, 54(4): 536–539, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/andrei-pop-
a-forest-of-symbols-art-
science-and-truth-in-the-
long-nineteenth-century-
new-york-zone-books-2019-
pp-320-isbn-9781935408369-
2500-hardback/
46871408D8490351C54A49F68FCA6973.

Hayes:2023:BRM

[Hay23]  Emily Hayes.  Book review:

Matthew H. Edney, Cartography: The Ideal and Its History. Chicago: The University of Chicago Press, 2019. Pp. 296. ISBN: 978-0-226-60568-5. $32.00 (paperback). British Journal for the History of Science, 56:

269–271, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/matthew-h-
                                        edney-cartography-the-
                                        ideal-and-its-history-
                                        chicago-the-university-of-
                                        chicago-press-2019-pp-296-
                                        isbn-9780226605685-3200-
                                        paperback/
                                        0DBD2E8881ED082587CD0A3ABEBF065D.

Heering:2023:TMR

[Hee23] Peter Heering. Transformations: the material representation of historical experiments in science teaching. British Journal for the History of Science, 56(3): 351–368, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/
                                        transformations-the-
                                        material-representation-of-
                                        historical-experiments-in-
                                        science-teaching/
                                        30408D9327C495A54C1DBBC1E390BF89.

Heffernan:2024:TDR

[Hef24] Anne Heffernan. Technopolitics, development and the residues of the South African state. British Journal for the History of Science, 57 (4):645–649, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.

org/core/journals/british-
journal-for-the-history-of-
science/article/
technopolitics-development-
and-the-residues-of-the-
south-african-state/
FB25FFEB66889B3D79CF03F1A59B5D35.

Helsing:2020:BRM

[Hel20] Daniel Helsing. Book review:

Michael Bravo, North Pole: Nature and Culture. London: Reaktion Books, 2019. Pp. 254. ISBN 978-1-78914-008-8. $24.95/£14.96 (paperback). British Journal for the History of Science, 53(1): 131–132, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/michael-
bravo-north-pole-nature-
and-culture-london-
reaktion-books-2019-pp-254-
isbn-9781789140088-
24951496-paperback/
37B3884E330C086F5FE7F651218FD375.

Hemmens:2022:BRM

[Hem22] Sue Hemmens. Book review:

Michel Anctil, Luminous Creatures: The History and Science of Light Production in Living Organisms, Montreal and Kingston: McGill-Queen’s University Press, 2018. Pp. xvii + 467. ISBN: 978-0-7735-5312-5. $49.95 (hardback). British Journal for the History of Science, 55(4):527– 529, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/michel-
                                        anctil-luminous-creatures-
                                        the-history-and-science-of-
                                        light-production-in-living-
                                        organisms-montreal-and-
                                        kingston-mcgillqueens-
                                        university-press-2018-pp-
                                        xvii-467-isbn-
                                        9780773553125-4995-hardback/
                                        2E3FE162F6076EE73D8AFE43E299D42D.

Henry:2024:LNC

[Hen24] John Henry. ‘Like nets or cobwebs’: Kenelm Digby, Isaac Newton and the problem of rarefaction. British Journal for the History of Science, 57(3): 447–466, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/like-nets-
                                        or-cobwebs-kenelm-digby-
                                        isaac-newton-and-the-
                                        problem-of-rarefaction/
                                        56CA95582BD6487F011224D0348E963A.

Herbert:2024:BRR

[Her24] Amanda E. Herbert. Book review: Rachel E. Walker, Beauty and the Brain: The Science of Human Nature in Early America, Chicago: University of Chicago Press, 2022. Pp. 288. ISBN 978-0-226-82256-3. $45.00 (cloth). British Journal for the History of Science, 57(1): 158–159, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/rachel-e-
walker-beauty-and-the-
brain-the-science-of-human-
nature-in-early-america-
chicago-university-of-
chicago-press-2022-pp-288-
isbn-9780226822563-4500-
cloth/394000DBB574A189252F66A5FF1D91AF.

Heslop:2023:BRH

[Hes23] Beck Chamberlain Heslop. Book review: Hans Radder, From Commodification to the Common Good: Reconstructing Science, Technology and Society. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 312. ISBN 978-0-8229-4579-6. $50.00 (hardcover). British Journal for the History of Science, 56:271–273, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/hans-radder-from-
commodification-to-the-
common-good-reconstructing-
science-technology-and-
society-pittsburgh-
university-of-pittsburgh-
press-2019-pp-312-isbn-
9780822945796-5000-hardcover/
EF17907F39AD6E900F73634610187881.

Hesketh:2025:DSG

[Hes25a] Ian Hesketh. Darwin’s scientific gardener: John Scott, the ‘physiological test’ and the importance of character in Victorian science. British Journal for the History of Science, 58(1):19–37, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/darwins-scientific-
gardener-john-scott-the-
physiological-test-and-the-
importance-of-character-in-
victorian-science/
F6DABBD3CEA5EFB2C7DA00B939569C9C.

Heslop:2025:BRG

[Hes25b] Beck Chamberlain Heslop. Book review: Gemma Almond-Brown, Spectacles and the Victorians: Measuring, Defining and Shaping Visual Capacity, Manchester: Manchester University Press, 2023. Pp. 296. ISBN 978-1-5261-6135-2. £85.00 (hardcover). British Journal for the History of Science, 58 (1):189–190, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/gemma-
almondbrown-spectacles-and-
the-victorians-measuring-
defining-and-shaping-
visual-capacity-manchester-
manchester-university-
press-2023-pp-296-isbn-
9781526161352-8500-hardcover/
EFCDE6177352CFFFD946CC04003006AE.

Higgitt:2024:BRS

[Hig24]  Rebekah Higgitt.  Book re-

view: Scott Alan Johnston, The Clocks Are Telling Lies: Science, Society, and the Construction of Time, Montreal: McGill-Queen’s University Press, 2022. Pp. 264. ISBN 978-0-2280-0843-9. CA$49.95 (cloth). British Journal for the History of Science, 57(2):307–309, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/scott-alan-
johnston-the-clocks-are-
telling-lies-science-
society-and-the-
construction-of-time-
montreal-mcgillqueens-
university-press-2022-pp-
264-isbn-9780228008439-
ca4995-cloth/D6ACB9EEA76A1B1092D34474F4956082.

Hinckley:2024:BRP

[Hin24] Marlis Hinckley. Book review:

Paola Bertucci, In the Land of Marvels: Science, Fabricated Realities, and Industrial Espionage in the Age of the Grand Tour, Baltimore: Johns Hopkins University Press, 2023. Pp. 184. ISBN 978-1-4214-4710-0. $54.95 (hardcover). British Journal for the History of Science, 57(4): 673–675, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-

                                        science/article/paola-
                                        bertucci-in-the-land-of-
                                        marvels-science-fabricated-
                                        realities-and-industrial-
                                        espionage-in-the-age-of-
                                        the-grand-tour-baltimore-
                                        johns-hopkins-university-
                                        press-2023-pp-184-isbn-
                                        9781421447100-5495-hardcover/
                                        7DCE9AE3BEAC0F4FBD72687A30F8D6B1.

Houghton:2020:RSU

[HLDS20] Jemma Houghton, Alexander Longworth-Dunbar, and Nicola Sugden. ‘Research sharing’ using social media: online conferencing and the experience of #BSHSGlobalHist. British Journal for the History of Science, 53 (4):555–573, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/research-
                                        sharing-using-social-media-
                                        online-conferencing-and-
                                        the-experience-of-
                                        bshsglobalhist/
                                        AB6B1C7AB14DA2EE2C827EC8EFD0D0BC.

Heumann:2022:IID

[HMB22] Ina Heumann, Anne Greenwood MacKinney, and Rainer Buschmann. Introduction: the issue of duplicates. British Journal for the History of Science, 55(3):257–278, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-

for-the-history-of-science/
article/introduction-the-
issue-of-duplicates/
B8811D4F80ED5CB40E357A41D4BBE5C2.

Hof:2024:BRC

[Hof24] Barbara Hof. Book review: Carola Sachse, Wissenschaft und Diplomatie: Die Max-Planck-Gesellschaft im Feld der internationale Politik (1945– 2000), Göttingen: Vandenhoeck & Ruprecht, 2023. Pp. 594. ISBN 978-3-525-30206-4. EUR80.00 (hardback). British Journal for the History of Science, 57 (2):309–310, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/carola-
sachse-wissenschaft-und-
diplomatie-die-
maxplanckgesellschaft-im-
feld-der-internationale-
politik-19452000-gottingen-
vandenhoeck-ruprecht-2023-pp-
594-isbn-9783525302064-8000-
hardback/E6478BA2068A9D0346BE7A9FC4DBCC36.

Holmes:2020:BRN

[Hol20]  Tarquin Holmes.  Book re-

view: Nicole C. Nelson, Model Behavior: Animal Experiments, Complexity, and the Genetics of Psychiatric Disorders. Chicago and London: The University of Chicago Press, 2018. Pp. 255. ISBN 978-0-226-54608-7. $30.00 (paperback). British Journal for the History of Science, 53 (2):279–280, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/nicole-c-
                                        nelson-model-behavior-
                                        animal-experiments-
                                        complexity-and-the-
                                        genetics-of-psychiatric-
                                        disorders-chicago-and-
                                        london-the-university-of-
                                        chicago-press-2018-pp-255-isbn-
                                        9780226546087-3000-paperback/
                                        64925BD46110A47C46DBD4689BF5BE4A.

Holmes:2021:BRV

[Hol21]  Tarquin Holmes.  Book re-

view: Vanessa Heggie, Higher and Colder: A History of Extreme Physiology and Exploration Chicago and London: The University of Chicago Press, 2019. Pp. 253. ISBN 978-0-226-65088-3. $40.00 (hardback). British Journal for the History of Science, 54(2):239–240, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/vanessa-
                                        heggie-higher-and-colder-a-
                                        history-of-extreme-
                                        physiology-and-exploration-
                                        chicago-and-london-the-
                                        university-of-chicago-
                                        press-2019-pp-253-isbn-
                                        9780226650883-4000-hardback/
                                        DBD5292B2F66BA87E8413C4E3FADC455.

Holmes:2022:BRC

[Hol22] Charlotte Holmes. Book review: Claire Hickman, The Doctor’s Garden: Medicine, Science, and Horticulture in Britain. New Haven: Yale University Press, 2021. Pp. xiv + 238. ISBN 978-0-300-23610-1. £30.00 (hardback). British Journal for the History of Science, 55(2):253–255, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/claire-hickman-the-
doctors-garden-medicine-
science-and-horticulture-
in-britain-new-haven-yale-
university-press-2021-pp-
xiv-238-isbn-9780300236101-
3000-hardback/
F20C5EC56ECBE1F443E4003CB791D726.

Holland:2023:BRM

[Hol23a] Sarah Holland. Book review:

Mikkel Borch-Jacobsen, Freud’s Patients: A Book of Lives. London: Reaktion Books, 2021. Pp. 304. ISBN 978-1-7891-4455-0. £20.00 (hardback). British Journal for the History of Science, 56(3): 413–414, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mikkel-
borchjacobsen-freuds-
patients-a-book-of-lives-
london-reaktion-books-2021-
pp-304-isbn-9781789144550-

                                        2000-hardback/
                                        011EE83E2A71B945B091347FD90B61AE.

Holmes:2023:BRJ

[Hol23b] Matthew Holmes. Book review: Johan Alfredo Linthorst, Research between Science, Society and Politics: The History and Scientific Development of Green Chemistry. Utrecht: Eburon, 2023. Pp. 268. ISBN 978-94-6301-434-2. EUR36.00 (paperback). British Journal for the History of Science, 56(4):589– 591, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/johan-
                                        alfredo-linthorst-research-
                                        between-science-society-
                                        and-politics-the-history-
                                        and-scientific-development-
                                        of-green-chemistry-utrecht-
                                        eburon-2023-pp-268-isbn-
                                        9789463014342-3600-paperback/
                                        1FA60F7CAD4769251F0FC50153C273A5.

Holmes:2023:PWT

[Hol23c] Matthew Holmes. A plague of weasels and ticks: animal introduction, ecological disaster, and the balance of nature in Jamaica, 1870–1900. British Journal for the History of Science, 56(3): 391–407, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/plague-of-

weasels-and-ticks-animal-
introduction-ecological-
disaster-and-the-balance-
of-nature-in-jamaica-
18701900/
111337F9B312A6994CBE64FCEF46E963.

Hopwood:2022:BMD

[Hop22]  Nick Hopwood.  ‘Not birth,

marriage or death, but gastrulation’: the life of a quotation in biology. British Journal for the History of Science, 55(1):1–26, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/not-birth-
marriage-or-death-but-
gastrulation-the-life-of-a-
quotation-in-biology/
277AA3AE743772CB4335A98836564E48.

Horder:2023:BRM

[Hor23]  Tim Horder.  Book review:

Marco Tamborini, The Architecture of Evolution: The Science of Form in Twentieth-Century Evolutionary Biology. Pittsburgh: University of Pittsburgh Press, 2022. Pp. 283. ISBN 978-0-8229-4735-6. $455.00 (hardcover). British Journal for the History of Science, 56(4):591–593, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/marco-tamborini-
the-architecture-of-

                                        evolution-the-science-of-
                                        form-in-twentiethcentury-
                                        evolutionary-biology-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2022-pp-
                                        283-isbn-9780822947356-45500-
                                        hardcover/C188762767439D9D845410586685230F.

Hall:2021:DTF

[HS21] Kersten Hall and Neeraja Sankaran. DNA translated: Friedrich Miescher’s discovery of nuclein in its original context. British Journal for the History of Science, 54(1):99–107, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/dna-translated-
                                        friedrich-mieschers-
                                        discovery-of-nuclein-in-
                                        its-original-context/
                                        60A9706BE7610FFD42F805AE636670FA.

Huang:2021:HKC

[Hua21] Chi Chi Huang. ‘Hong Kong can afford a typhoon or two’: British discussions of revolving storms. British Journal for the History of Science, 54(3): 327–339, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/hong-kong-
                                        can-afford-a-typhoon-or-
                                        two-british-discussions-of-
                                        revolving-storms/
                                        75D7F1587210624FAC6E84C211D9CFA4.

Hunt:2023:BRR

[Hun23] Bruce J. Hunt. Book review:

Richard Noakes, Physics and Psychics: The Occult and the Sciences in Modern Britain. Cambridge: Cambridge University Press, 2019. Pp. 419. ISBN 978-1-316-88243-6. £90.00 (hardcover). British Journal for the History of Science, 56(3):415–416, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/richard-noakes-
physics-and-psychics-the-
occult-and-the-sciences-in-
modern-britain-cambridge-
cambridge-university-press-
2019-pp-419-isbn-
9781316882436-9000-
hardcover/6C9F3AE766ED6473C2C4EDA493255E8B.

Hyun:2024:NCC

[Hyu24] Jaehwan Hyun. Negotiating conservation and competition: national parks and ‘victory-over-communism’ diplomacy in South Korea. British Journal for the History of Science, 57(2): 239–255, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
negotiating-conservation-
and-competition-national-
parks-and-
victoryovercommunism-

                                        diplomacy-in-south-korea/
                                        CD1C11D581897D1329324F3060D9E745.

Iqbal:2024:JAA

[Iqb24] Aashique Ahmed Iqbal. Jodhpur and the aeroplane: aviation and diplomacy in an Indian state 1924–1952. British Journal for the History of Science, 57(2):175–189, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/jodhpur-
                                        and-the-aeroplane-aviation-
                                        and-diplomacy-in-an-indian-
                                        state-19241952/
                                        CC0EEC0EEA4F282D7344095485FE5B40.

Ishibashi:2020:CAG

[Ish20] Yuto Ishibashi. Constructing the ‘automatic’ Greenwich time system: George Biddell Airy and the telegraphic distribution of time, c. 1852–1880. British Journal for the History of Science, 53 (1):25–46, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/
                                        constructing-the-automatic-
                                        greenwich-time-system-
                                        george-biddell-airy-and-
                                        the-telegraphic-
                                        distribution-of-time-
                                        c18521880/
                                        BC3BFBBA2C010C3D73F7468EEA45AF90.

Ito:2024:TSA

[Ito24] Kenji Ito. Transnational scientific advising: occupied Japan, the United States National Academy of Sciences and the establishment of the Science Council of Japan. British Journal for the History of Science, 57(2): 257–271, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
transnational-scientific-
advising-occupied-japan-
the-united-states-national-
academy-of-sciences-and-
the-establishment-of-the-
science-council-of-japan/
8845DA5025BDD2AF9A7E61049E345FA7.

Iwata:2025:BRJ

[Iwa25]  Naoya Iwata.  Book review:

Justin Humphreys, The Invention of Imagination: Aristotle, Geometry, and the Theory of the Psyche, Pittsburgh: University of Pittsburgh Press, 2023. Pp. 176. ISBN 978-0-8229-4740-0. $50.00 (hardcover). British Journal for the History of Science, 58 (1):155–156, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/justin-
humphreys-the-invention-of-
imagination-aristotle-
geometry-and-the-theory-of-

                                        the-psyche-pittsburgh-
                                        university-of-pittsburgh-
                                        press-2023-pp-176-isbn-
                                        9780822947400-5000-
                                        hardcover/8279BBABB594B8A8BFF648F401F4A78A.

Jackson:2021:BRB

[Jac21] Roland Jackson. Book review: Bernard Lightman and Bennet Zon (eds.), Victorian Culture and the Origin of Disciplines, New York: Routledge, 2020. Pp. 336. ISBN 978-0-367-22842-2. £120.00 (hardback). British Journal for the History of Science, 54 (2):248–250, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/bernard-
                                        lightman-and-bennet-zon-
                                        eds-victorian-culture-and-
                                        the-origin-of-disciplines-
                                        new-york-routledge-2020-pp-
                                        336-isbn-9780367228422-
                                        12000-hardback/
                                        8E25BBDC5E6B0ADFB6D6B3E2147195FE.

Jagessar:2025:BRK

[Jag25]  Philip Jagessar.  Book re-

view: Kamlesh Mohan, Science and Technology in Colonial India, London: Routledge, 2022. Pp. 176. ISBN 978-1-032-36479-

7. £96.00 (hardback). British Journal for the History of Science, 58(1):185–186, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-

for-the-history-of-science/
article/kamlesh-mohan-
science-and-technology-in-
colonial-india-london-
routledge-2022-pp-176-isbn-
9781032364797-9600-
hardback/
1F1B3D6D3BF6E226391A911B49841824.

Jalloh:2025:BRS

[Jal25] Mahmoud Jalloh. Book review:

Steven French, A Phenomenological Approach to Quantum Mechanics: Cutting the Chain of Correlations, Oxford: Oxford University Press, 2023. Pp. 288. ISBN 978-0-19-889795-8. £80.00 (hardcover). British Journal for the History of Science, 58 (1):165–166, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/steven-
french-a-phenomenological-
approach-to-quantum-
mechanics-cutting-the-
chain-of-correlations-
oxford-oxford-university-
press-2023-pp-288-isbn-
9780198897958-8000-
hardcover/A80BC7171C2ACDCEE15D215213C13C71.

James:2021:BRM

[Jam21] Frank A. J. L. James. Book review: Michael Wheeler, The Athenæum: More than Just Another London Club New Haven: Yale University Press, 2020. Pp. 440. ISBN 978-0-300-24677-3. £35.00 (hardback). British Journal for the History of Science, 54(3): 390–391, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/michael-
                                        wheeler-the-athenaeum-more-
                                        than-just-another-london-
                                        club-new-haven-yale-
                                        university-press-2020-pp-
                                        440-isbn-9780300246773-
                                        3500-hardback/
                                        B3FF9F34C6B0CC6F91A4E9C197C160C0.

Johnston:2021:MOD

[Joh21] Scott Alan Johnston. Managing the observatory: discipline, order and disorder at Greenwich, 1835–1933. British Journal for the History of Science, 54(2):155–175, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/managing-
                                        the-observatory-discipline-
                                        order-and-disorder-at-
                                        greenwich-18351933/
                                        0AAEEF7AC342DC4D5AC3AAE7D2DFC2E6.

Jones:2021:BRI

[Jon21] Andrew Jones. Book review: Ido Hartogsohn, The American Trip: Set, Setting, and the Psychedelic Experience in the Twentieth Century Cambridge, MA: MIT Press, 2020. Pp. 432. ISBN: 978-0-262-53914-2. $35.00 (paperback). British Journal for the History of Science, 54 (1):109–111, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ido-
hartogsohn-the-american-
trip-set-setting-and-the-
psychedelic-experience-in-
the-twentieth-century-
cambridge-ma-mit-press-
2020-pp-432-isbn-
9780262539142-3500-paperback/
FEE939414B35F753A8F93AE80AECA12D.

Jones:2023:SSS

[Jon23] Elin Jones. Stratifying seamanship: sailors’ knowledge and the mechanical arts in eighteenthcentury Britain. British Journal for the History of Science, 56(1):45–63, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
stratifying-seamanship-
sailors-knowledge-and-the-
mechanical-arts-in-
eighteenthcentury-britain/
C5A91723E125B35035F20F27412E1909.

Jones:2024:BRS

[Jon24] Kevin Matthew Jones. Book review: Susanne Schmidt, Midlife Crisis: The Feminist Origins of a Chauvinist Cliché, Chicago: University of Chicago Press, 2020. Pp. 280. ISBN 978-0-226-63714-

3. $89.00 (cloth). British Journal for the History of Science, 57(3):

494–495, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/susanne-
                                        schmidt-midlife-crisis-the-
                                        feminist-origins-of-a-
                                        chauvanist-cliche-chicago-
                                        university-of-chicago-
                                        press-2020-pp-280-isbn-
                                        9780226637143-8900-cloth/
                                        20AFB50CB698FC46F2E3A84E24AD52F6.

See corrigendum [Jon25].

Jones:2025:SSM

[Jon25] Kevin Matthew Jones. Susanne Schmidt, Midlife Crisis: The Feminist Origins of a Chauvinist Cliché — Corrigendum. British Journal for the History of Science, 58(1):193, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/susanne-
                                        schmidt-midlife-crisis-the-
                                        feminist-origins-of-a-
                                        chauvinist-cliche-
                                        corrigendum/
                                        9262D7148264C8A6CB6B5999DEB87447.

See [Jon24].

Jordanova:2023:BRA

[Jor23] Ludmilla Jordanova. Book review: Andrew Cunningham, ‘I Follow Aristotle’: How William Harvey Discovered the Circulation of the Blood. London: Routledge, 2022. Pp. xii + 180.

ISBN 978-1-032-16223-2. £130.00 (hardback). British Journal for the History of Science, 56: 273–274, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/andrew-
cunningham-i-follow-
aristotle-how-william-
harvey-discovered-the-
circulation-of-the-blood-
london-routledge-2022-pp-
xii-180-isbn-9871032162232-
13000-hardback/00E95A84A19D6F3F5AFEFF6DBA34D699.

Kadia:2024:BRR

[Kad24] Miriam Kingsberg Kadia. Book review: Ran Zwigenberg, Nuclear Minds: Cold War Psychological Science and the Bombings of Hiroshima and Nagasaki, Chicago: The University of Chicago Press, 2023. Pp. 304. ISBN 978-0-226-82676-9. $35.00 (paperback). British Journal for the History of Science, 57 (1):159–161, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/ran-
zwigenberg-nuclear-minds-
cold-war-psychological-
science-and-the-bombings-
of-hiroshima-and-nagasaki-
chicago-the-university-of-
chicago-press-2023-pp-304-
isbn-9780226826769-3500-paperback/
5D49F31E9A52108201107D08D244E2BC.

Kaiser:2022:DNB

[Kai22] Katja Kaiser. Duplicate networks: the Berlin botanical institutions as a ‘clearing house’ for colonial plant material, 1891– 1920. British Journal for the History of Science, 55(3):279– 296, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/duplicate-
                                        networks-the-berlin-
                                        botanical-institutions-as-
                                        a-clearing-house-for-
                                        colonial-plant-material-
                                        18911920/
                                        FC18720ACE9E3B4DA1748D55D8971406.

Kallstrand:2022:SNC

[Käl22] Gustav Källstrand. Science by Nobel committee: decision making and norms of scientific practice in the early physics and chemistry prizes. British Journal for the History of Science, 55(2):187–205, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/science-by-
                                        nobel-committee-decision-
                                        making-and-norms-of-
                                        scientific-practice-in-the-
                                        early-physics-and-
                                        chemistry-prizes/
                                        48B80DE55848512740AE7E05A5C77717.

Kana:2020:BRS

[Kan20] Sandip Kana. Book review:

Sabine Clarke, Science at the End of Empire: Experts and the Development of the British Caribbean, 1940–62. Manchester: Manchester University Press, 2018. Pp. 224. ISBN 978-1-5261-3138-6. £80.00 (hardback). British Journal for the History of Science, 53 (4):593–594, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/sabine-
clarke-science-at-the-end-
of-empire-experts-and-the-
development-of-the-british-
caribbean-194062-
manchester-manchester-
university-press-2018-pp-
224-isbn-9781526131386-8000-
hardback/90D253E51E1B9D3B351B903774A45CB6.

Kanazawa:2024:PME

[Kan24] Reiko Kanazawa. The politics of medical expertise and substance control: WHO consultants for addiction rehabilitation and pharmacy education in Thailand and India during the Cold War. British Journal for the History of Science, 57(2):221–238, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/politics-
of-medical-expertise-and-

                                        substance-control-who-
                                        consultants-for-addiction-
                                        rehabilitation-and-
                                        pharmacy-education-in-
                                        thailand-and-india-during-
                                        the-cold-war/
                                        333942551DF2C12F7629CBA777EB5E2C.

Katz:2020:BRR

[Kat20] Osnat Katz. Book review: Roger D. Launius, Reaching for the Moon: A Short History of the Space Race. New Haven, CT: Yale University Press, 2019. Pp. 256. ISBN 978-0-300-23046-8. $30.00 (hardback). British Journal for the History of Science, 53(4): 603–604, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/roger-d-
                                        launius-reaching-for-the-
                                        moon-a-short-history-of-
                                        the-space-race-new-haven-
                                        ct-yale-university-press-
                                        2019-pp-256-isbn-
                                        9780300230468-3000-
                                        hardback/
                                        F10517977303F5486CAE4278316021B7.

Kempton:2023:CTP

[Kem23] Miles Kempton. Commercial television and primate ethology: facial expressions between Granada and London Zoo. British Journal for the History of Science, 56(1):83–102, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-

journal-for-the-history-of-
science/article/commercial-
television-and-primate-
ethology-facial-
expressions-between-
granada-and-london-zoo/
FEF5D5661892E52B718B001396B190C9.

Kent:2025:BRP

[Ken25] Deborah Kent. Book review:

Philip Beeley and Christopher D. Hollings (eds.), Beyond the Learned Academy: The Practice of Mathematics, 1600–1850, Oxford: Oxford University Press, 2024. Pp. 512. ISBN 978-0-19-886395-3. £35.00 (hardback). British Journal for the History of Science, 58(2):363–364, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/philip-
beeley-and-christopher-d-
hollings-eds-beyond-the-
learned-academy-the-
practice-of-mathematics-
16001850-oxford-oxford-
university-press-2024-pp-
512-isbn-9780198863953-3500-
hardback/1C6F44E1A911FC7579C4854513D9BDB1.

Kern:2020:AEA

[Ker20] Emily M. Kern. Archaeology enters the ‘atomic age’: a short history of radiocarbon, 1946– 1960. British Journal for the History of Science, 53(2):207–227, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/archaeology-enters-
                                        the-atomic-age-a-short-
                                        history-of-radiocarbon-
                                        19461960/
                                        855D6E2E66D09F7467C6051C77C2D6B0.

Khramov:2024:FDS

[Khr24] Alexander V. Khramov. A forerunner of Darwin in the service of nihilists: the translation and reception of Vestiges in Russia. British Journal for the History of Science, 57 (1):65–79, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/forerunner-
                                        of-darwin-in-the-service-
                                        of-nihilists-the-
                                        translation-and-reception-
                                        of-vestiges-in-russia/
                                        891ACFF6C6BB9C45FB635D9B15C6DDCF.

Kirwan:2020:SRM

[Kir20] Adrian Kirwan. Sociability, radium and the maintenance of scientific culture and authority in twentieth-century Ireland: a case study of the Royal Dublin Society. British Journal for the History of Science, 53(1):47–66, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/sociability-radium-

and-the-maintenance-of-
scientific-culture-and-
authority-in-
twentiethcentury-ireland-a-
case-study-of-the-royal-
dublin-society/
F72934AB17A85C3BD5E533EDD1D6DF55.

Khayat:2022:MAL

[KK22] Nicole Khayat and Liat Kozma.

Medicine and Arabic literary production in the Ottoman Empire during the nineteenth century. British Journal for the History of Science, 55(4):515– 524, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/medicine-
and-arabic-literary-
production-in-the-ottoman-
empire-during-the-
nineteenth-century/
9EDBADC6DFE673E6B47C37672ABD4E0F.

Klerk:2020:NHP

[Kle20]  Saskia  Klerk.

Natural history in the physician’s study: Jan Swammerdam (1637– 1680), Steven Blankaart (1650– 1705) and the ‘paperwork’ of observing insects. British Journal for the History of Science, 53(4): 497–525, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/natural-
history-in-the-physicians-

                                        study-jan-swammerdam-
                                        16371680-steven-blankaart-
                                        16501705-and-the-paperwork-
                                        of-observing-insects/
                                        8D7EA3137DC5F069D32F98F38BA08896.

Kline:2024:BRS

[Kli24] Cameron L. Kline. Book review: Sarah E. Naramore, Benjamin Rush, Civic Health, and Human Illness in the Early American Republic, Rochester: University of Rochester Press, 2023. Pp. 308. ISBN 978-1-64825-069-9. £97.00 (hardcover). British Journal for the History of Science, 57(4):670–671, December 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/sarah-e-naramore-
                                        benjamin-rush-civic-health-
                                        and-human-illness-in-the-
                                        early-american-republic-
                                        rochester-university-of-
                                        rochester-press-2023-pp-
                                        308-isbn-9781648250699-
                                        9700-hardcover/6B0A807ECA95CF031AA0AB7525B836CE.

Kohlstedt:2023:BRH

[Koh23] Sally Gregory Kohlstedt. Book review: Hannah Wills, Sadie Harrison, Erika Jones, Rebecca Martin and Farrah Lawrence-Mackey (eds.), Women in the History of Science: A Sourcebook. London: UCL Press, 2023. Pp. xxviii + 446. ISBN 978-1-80008-415-

5. £50.00 (hardback); £30.00 (paperback); £0.00 (open-access PDF). British Journal for the History of Science, 56(4):593– 595, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/hannah-
wills-sadie-harrison-erika-
jones-rebecca-martin-and-
farrah-lawrencemackey-eds-
women-in-the-history-of-
science-a-sourcebook-
london-ucl-press-2023-pp-
xxviii-446-isbn-
9781800084155-5000-hardback-
3000-paperback-000-openaccess-
pdf/7BBA7A0D78540A181E0FCDF044CE7805.

Komel:2024:PIS

[Kom24] Svit Komel. Petty’s instruments: the Down Survey, territorial natural history and the birth of statistics. British Journal for the History of Science, 57(1):43–64, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/pettys-
instruments-the-down-
survey-territorial-natural-
history-and-the-birth-of-
statistics/
0C35B4A6C753DE4DF81B523B4AD6B62B.

Korolija:2021:BRK

[Kor21] Maja Korolija. Book review:

Keith Ewing, Joan Mahoney and Andrew Moretta, MI5, the Cold War, and the Rule of Law. Oxford: Oxford University Press, 2020. Pp. 528. ISBN: 978-0-19-881862-5. £84.00 (hardback). British Journal for the History of Science, 54(4):534–536, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/keith-ewing-joan-
                                        mahoney-and-andrew-moretta-
                                        mi5-the-cold-war-and-the-
                                        rule-of-law-oxford-oxford-
                                        university-press-2020-pp-
                                        528-isbn-9780198818625-
                                        8400-hardback/
                                        6EFA68E8DABFC6F6722B1C5484DE1993.

Kotsou:2023:FIC

[Kot23] Georgiana Kotsou. Functional informality: crafting social interaction toward scientific productivity at the Gordon Research Conferences, 1950–1980. British Journal for the History of Science, 56 (4):519–534, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/functional-
                                        informality-crafting-
                                        social-interaction-toward-
                                        scientific-productivity-at-
                                        the-gordon-research-
                                        conferences-19501980/
                                        533749102A39D7A7EA41DDA0F5596E7D.

Krige:2023:BRG

[Kri23] John Krige. Book review: Gordon Barrett, China’s Cold War Science Diplomacy. Cambridge:

Cambridge University Press, 2022. Pp. 300. ISBN 978-1-1088-4457-4. £75.00 (hardback). British Journal for the History of Science, 56 (1):119–121, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/gordon-
barrett-chinas-cold-war-
science-diplomacy-
cambridge-cambridge-
university-press-2022-pp-
300-isbn-9781108844574-
7500-hardback/
8B2858DB0642844D0C474798942D3DED.

Kuchinskaya:2025:BRJ

[Kuc25] Olga Kuchinskaya. Book review:

Jacob Darwin Hamblin and Linda Marie Richards (eds.), Making the Unseen Visible: Science and the Contested Histories of Radiation Exposure, Corvallis: Oregon State University Press, 2023. Pp. 400. ISBN 978-0-87071-253-1. $39.95 (paperback). British Journal for the History of Science, 58(2):391–392, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jacob-
darwin-hamblin-and-linda-
marie-richards-eds-making-
the-unseen-visible-science-
and-the-contested-
histories-of-radiation-
exposure-corvallis-oregon-

                                        state-university-press-
                                        2023-pp-400-isbn-9780870712531-
                                        3995-paperback/D5F1757A257752221B991DFD4B2C0944.

Kaalund:2021:ALM

[KW21] Nanna Katrine Lüders Kaalund and John Woitkowitz. ‘Ancient lore with modern appliances’: networks, expertise, and the making of the Open Polar Sea, 1851–1853. British Journal for the History of Science, 54(3):277–299, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/ancient-lore-with-
                                        modern-appliances-networks-
                                        expertise-and-the-making-
                                        of-the-open-polar-sea-
                                        18511853/
                                        9FB26207EDAE5DBD31EE96D5EC829B0B.

Langstaff:2022:BRE

[Lan22] Alex Langstaff. Book review: Elena Aronova, Scientific History: Experiments in History and Politics from the Bolshevik Revolution to the End of the Cold War, Chicago: The University of Chicago Press, 2021. Pp. 256. ISBN 978-0-226-76138-1. $45.00 (hardback). British Journal for the History of Science, 55 (1):124–126, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/elena-

aronova-scientific-history-
experiments-in-history-and-
politics-from-the-
bolshevik-revolution-to-
the-end-of-the-cold-war-
chicago-the-university-of-
chicago-press-2021-pp-256-
isbn-9780226761381-4500-hardback/
B1B940818A69D19F72FCD39BB374A060.

Lanzoni:2024:BRR

[Lan24] Susan Lanzoni. Book review:

Roger Smith, Kinasthesia in the Psychology, Philosophy and Culture of Human Experience, London: Routledge, 2023. Pp. 156. ISBN 978-1-032-43590-9. £48.99 (hardback). British Journal for the History of Science, 57(4): 666–668, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/roger-
smith-kinasthesia-in-the-
psychology-philosophy-and-
culture-of-human-
experience-london-
routledge-2023-pp-156-isbn-
9781032435909-4899-
hardback/
B03380E554DB77644E105ECF964C4B72.

Law:2021:BRD

[Law21a] Harry Law. Book review: Dylan Mulvin, Proxies: The Cultural Work of Standing In London: MIT Press, 2021. Pp. 228. ISBN 978-0-262-04514-8. £40.00 (paperback). British Journal for the History of Science, 54(4): 532–534, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/dylan-
                                        mulvin-proxies-the-
                                        cultural-work-of-standing-
                                        in-london-mit-press-2021-
                                        pp-228-isbn-9780262045148-
                                        4000-paperback/
                                        E01DA7CD7A6D033918C93B89506716C9.

Lawrence:2021:RMY

[Law21b] Christopher Lawrence. Robert M. Young’s mind, brain and adaptation revisited. British Journal for the History of Science, 54(1):61–77, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/robert-m-
                                        youngs-mind-brain-and-
                                        adaptation-revisited/
                                        6896719470FBB886C23398325DA81035.

Law:2023:BRD

[Law23a] Harry Law. Book review: Dylan Mulvin, Proxies: The Cultural Work of Standing In London: MIT Press, 2021. Pp. 228. ISBN 978-0-262-04514-8. £40.00 (paperback). — Corrigendum. British Journal for the History of Science, 56:281, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/dylan-

mulvin-proxies-the-
cultural-work-of-standing-
in-london-mit-press-2021-
pp-228-isbn-9780262045148-
4000-paperback-corrigendum/
7C4031DC2A1A212EB4EA8111953FCBD0.

Laws:2023:BRD

[Law23b] Harry Laws. Book review: Dylan Mulvin, Proxies: The Cultural Work of Standing In London: MIT Press, 2021. Pp. 228. ISBN 978-0-2620-4514-8. £40.00 (paperback). British Journal for the History of Science, 56 (1):115–117, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/dylan-
mulvin-proxies-the-
cultural-work-of-standing-
in-london-mit-press-2021-
pp-228-isbn-9780262045148-
4000-paperback/
21FD56578E1074C5F6BEA6BD0E82987D.

Law:2025:BRB

[Law25]  Harry Law.  Book review:

Bernardo Gonçalves, The Turing Test Argument, New York: Routledge, 2023. Pp. 238. ISBN 978-1-032-29157-4. £130.00 (hardback). British Journal for the History of Science, 58(2):365–366, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bernardo-goncalves-

                                        the-turing-test-argument-
                                        new-york-routledge-2023-pp-
                                        238-isbn-9781032291574-
                                        13000-hardback/
                                        C5E15DE777D7496F6AF54E1AA8D69B09.

Lidwell-Durnin:2020:CFD

[LD20] John Lidwell-Durnin. Cultivating famine: data, experimentation and food security, 1795– 1848. British Journal for the History of Science, 53(2):159–181, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/cultivating-famine-
                                        data-experimentation-and-
                                        food-security-17951848/
                                        A7FD0427A16C56306914F54969948331.

Lidwell-Durnin:2021:BRR

[LD21] John Lidwell-Durnin. Book review: Rebecca Earle, Feeding the People: The Politics of the Potato, Cambridge: Cambridge University Press, 2020. Pp. 308. ISBN 978-1-108-68845-1. £17.99 (hardback). British Journal for the History of Science, 54 (2):242–244, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/rebecca-
                                        earle-feeding-the-people-
                                        the-politics-of-the-potato-
                                        cambridge-cambridge-
                                        university-press-2020-pp-
                                        308-isbn-9781108688451-

1799-hardback/
EEB700F5AAA9275294E5217CC92B0BD7.

Ledgerwood:2021:BRC

[Led21] Emmeline Ledgerwood. Book review: C. Bruce Tarter, The American Lab: An Insider’s History of the Lawrence Livermore National Laboratory, Baltimore: Johns Hopkins University Press, 2018. Pp. 453. ISBN 978-1-4214-2531-3. $79.95 (hardback). British Journal for the History of Science, 54 (2):246–248, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/c-bruce-
tarter-the-american-lab-an-
insiders-history-of-the-
lawrence-livermore-
national-laboratory-
baltimore-johns-hopkins-
university-press-2018-pp-
453-isbn-9781421425313-7995-
hardback/CAB848C2FE28EE1C2B1982C0E2D2ECAC.

Ledgerwood:2022:ANB

[Led22] Emmeline Ledgerwood. ‘Armed with the necessary background of knowledge’: embedding science scrutiny mechanisms in the UK Parliament. British Journal for the History of Science, 55(2):167–185, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/armed-with-

                                        the-necessary-background-
                                        of-knowledge-embedding-
                                        science-scrutiny-
                                        mechanisms-in-the-uk-
                                        parliament/
                                        4A0A4BB7B04D5AC31EBA8E69AEDC3B5A.

Lehnen:2023:BRM

[Leh23] Odile Manon Lehnen. Book review: Michelle DiMeo, Lady Ranelagh: The Incomparable Life of Robert Boyle’s Sister. Chicago: University of Chicago Press, 2021. Pp. 296. ISBN 978-0-226-73160-

5. $45.00 (cloth). British Journal for the History of Science, 56 (4):595–596, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/michelle-
                                        dimeo-lady-ranelagh-the-
                                        incomparable-life-of-
                                        robert-boyles-sister-
                                        chicago-university-of-
                                        chicago-press-2021-pp-296-
                                        isbn-9780226731605-4500-
                                        cloth/
                                        2495BC5A2BDB810CD566A9AFFDED2C1B.

Liang:2024:TDE

[Lia24] Yue Liang. Technology diplomacy in early Communist China: the visit to the Jingjiang Flood Diversion Project in 1952. British Journal for the History of Science, 57(2):191–203, June 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-

science/article/technology-
diplomacy-in-early-
communist-china-the-visit-
to-the-jingjiang-flood-
diversion-project-in-1952/
4BB4DFDE8BFACB6C51B60D3F74C3A657.

Lockley:2024:BRK

[Loc24]  Tim Lockley.  Book review:

Kalle Kananoja, Healing Knowledge in Atlantic Africa: Medical Encounters, 1500–1800, Cambridge: Cambridge University Press, 2021. Pp. 272. ISBN 978-1-108-49125-9. $29.99 (paperback). British Journal for the History of Science, 57(1): 150–152, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/kalle-
kananoja-healing-knowledge-
in-atlantic-africa-medical-
encounters-15001800-
cambridge-cambridge-
university-press-2021-pp-
272-isbn-9781108491259-
2999-paperback/C59A7A2919913A27D8640D520E81A2F6.

Long:2020:CBN

[Lon20] Max Long. The ciné-biologists:

natural history film and the coproduction of knowledge in interwar Britain. British Journal for the History of Science, 53(4): 527–551, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-

                                        science/article/
                                        cinebiologists-natural-
                                        history-film-and-the-
                                        coproduction-of-knowledge-
                                        in-interwar-britain/
                                        ED1CAE4195B4C3030705CD8A1D80B5AC.

Lorek:2024:BRM

[Lor24] Timothy Lorek. Book review: Marci R. Baranski, The Globalization of Wheat, Pittsburgh: University of Pittsburgh Press, 2022. Pp. 256. ISBN 978-0-8229-4734-9. $55.00 (hardcover). British Journal for the History of Science, 57(1):140–142, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/marci-r-baranski-
                                        the-globalization-of-wheat-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2022-pp-
                                        256-isbn-9780822947349-
                                        5500-hardcover/
                                        66346F2CF31E50C08200E002552F1CE9.

Loske:2025:BRT

[Los25] Alexandra Loske. Book review: Tanja C. Kleinwächter, Sarah Lowengard and Friedrich Steinle (eds.), Ordering Colours in 18th and Early 19th Century Europe, Cham: Springer Nature, 2023. Pp. 201. ISBN 978-3-031-34955-3. £99.99 (hardcover). British Journal for the History of Science, 58(2):349–350, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.

org/core/journals/british-
journal-for-the-history-of-
science/article/tanja-c-
kleinwachter-sarah-
lowengard-and-friedrich-
steinle-eds-ordering-
colours-in-18th-and-early-
19th-century-europe-cham-
springer-nature-2023-pp-
201-isbn-9783031349553-
9999-hardcover/9D7CB9501D68A03A2444D4D62E77F2F8.

Lowe:2021:APH

[Low21] James W. E. Lowe. Adjusting to precarity: how and why the Roslin Institute forged a leading role for itself in international networks of pig genomics research. British Journal for the History of Science, 54(4):507–530, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/adjusting-to-
precarity-how-and-why-the-
roslin-institute-forged-a-
leading-role-for-itself-in-
international-networks-of-
pig-genomics-research/
A4B4243748211FCEAAA55362393906EA.

Li:2024:IAP

[LW24] Tianyu Li and Chadwick Wang.

Innovation amidst post-socialist reform: Jonas Salk and the birth of the Sabin strains-derived inactivated polio vaccine in China. British Journal for the History of Science, 57(3):407–423, September 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/innovation-amidst-
                                        postsocialist-reform-jonas-
                                        salk-and-the-birth-of-the-
                                        sabin-strainsderived-
                                        inactivated-polio-vaccine-
                                        in-china/
                                        E91233BFB8E3F633DD3CEFFFAFCCC809.

Luciano:2023:ASA

[LZ23]  Eugenio  Luciano  and

Elena Zanoni. Antonio Stoppani’s ‘Anthropozoic’ in the context of the Anthropocene. British Journal for the History of Science, 56 (1):103–114, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/antonio-
                                        stoppanis-anthropozoic-in-
                                        the-context-of-the-
                                        anthropocene/
                                        1E6A458F79F7028D603D68BCC8F1086B.

Luk:2024:BRJ

[LZ24] Christine Y. L. Luk and Longkai Zang. Book review: Jeffrey Womack, Radiation Evangelists: Technology, Therapy, and Uncertainty at the Turn of the Century, Pittsburgh: University of Pittsburgh Press, 2020. Pp. 288. ISBN 978-0-8229-4609-0. $35.00 (hardcover). British Journal for the History of Science, 57 (1):123–125, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jeffrey-
womack-radiation-
evangelists-technology-
therapy-and-uncertainty-at-
the-turn-of-the-century-
pittsburgh-university-of-
pittsburgh-press-2020-pp-
288-isbn-9780822946090-
3500-hardcover/35E3643D68F72CAE27C19E57C43276DE.

Ma:2025:LIVa

[Ma25a] Haitian Ma. Leapfrogging India: Vikram Sarabhai and the developmental promise of geocentric spaceflight. British Journal for the History of Science, 58(2):261–280, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
leapfrogging-india-vikram-
sarabhai-and-the-
developmental-promise-of-
geocentric-spaceflight/
E342E7B439CD0A2F7661247164C135C2.

See erratum [Ma25b].

Ma:2025:LIVb

[Ma25b] Haitian Ma. Leapfrogging India:

Vikram Sarabhai and the developmental promise of geocentric spaceflight — erratum. British Journal for the History of Science, 58(2):401, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

                                        journal-for-the-history-of-
                                        science/article/
                                        leapfrogging-india-vikram-
                                        sarabhai-and-the-
                                        developmental-promise-of-
                                        geocentric-spaceflight-
                                        erratum/
                                        55B8C853AFA29A16DC820A43EF1F6539.

See [Ma25a].

MacDonald:2020:BRP

[Mac20] Lee T. MacDonald. Book review: Peter R. Broughton, Northern Star: J. S. Plaskett. Toronto and London: University of Toronto Press, 2018. Pp. xx + 539. ISBN 978-1-4426-3017-8. $90.00 (hardback). British Journal for the History of Science, 53(1): 125–126, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/peter-r-
                                        broughton-northern-star-js-
                                        plaskett-toronto-and-
                                        london-university-of-
                                        toronto-press-2018-pp-xx-
                                        539-isbn-9781442630178-
                                        9000-hardback/
                                        D3E7D01CDD7A34B05F0267A2B7CB3AE2.

MacKinney:2022:DUH

[Mac22] Anne Greenwood MacKinney. Duplicates under the hammer: natural-history auctions in Berlin’s early nineteenth-century collection landscape. British Journal for the History of Science, 55(3):319–339, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/duplicates-under-
the-hammer-naturalhistory-
auctions-in-berlins-early-
nineteenthcentury-
collection-landscape/
1BED5F7051264DCD6CBA0AC7CA750C25.

Manias:2024:BRI

[Man24] Chris Manias. Book reviews: Irina Podgorny, Florentino Ameghino y hermanos, Buenos Aires: Edhasa, 2021. Pp. 346. ISBN 978-987-628-598-8. AR$9,150.00 (softcover) — Irina Podgorny, Los Argentinos vienen de los peces: Ensayo de filogenia nacional, Rosario: Beatriz Viterbo, 2021. Pp. 184. ISBN 978-9500845-399-0. AR$3,978.00 (softcover). British Journal for the History of Science, 57(1): 142–144, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/irina-
podgorny-florentino-
ameghino-y-hermanos-buenos-
aires-edhasa-2021-pp-346-
isbn-9789876285988-
ar915000-softcover-irina-
podgorny-los-argentinos-
vienen-de-los-peces-ensayo-de-
filogenia-nacional-rosario-beatriz-
viterbo-2021-pp-184-isbn-97895008453990-
ar397800-softcover/99D7B6D9B2B6DBE0087F944F1B271BCE.

Marschall:2020:BRC

[Mar20] Wythe Marschall. Book review: Chris Thorogood and Simon Hiscock, The Botany of Gin. Oxford: Bodleian Library, University of Oxford, 2020. Pp. 102. ISBN 978-1-85124-553-6. £15.00 (hardback). British Journal for the History of Science, 53(4): 600–602, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/chris-
                                        thorogood-and-simon-
                                        hiscock-the-botany-of-gin-
                                        oxford-bodleian-library-
                                        university-of-oxford-2020-
                                        pp-102-isbn-9781851245536-
                                        1500-hardback/
                                        190D9E8C30DBC5AE63F61BB5F3D9EDB1.

Margocsy:2022:MMG

[Mar22] Dániel Margócsy. Malinowski and malacology: global value systems and the issue of duplicates. British Journal for the History of Science, 55(3): 389–409, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/malinowski-
                                        and-malacology-global-
                                        value-systems-and-the-
                                        issue-of-duplicates/
                                        9727A997935430F98BF396C86C98718A.

Martin:2024:MHP

[Mar24a] Craig Martin. Medicine and the heavens in Padua’s Faculty of Arts, 1570–1630. British Journal for the History of Science, 57 (4):545–559, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/medicine-
and-the-heavens-in-paduas-
faculty-of-arts-15701630/
933FECA7301E23A58D934570E3C0B50E.

Martin:2024:BRC

[Mar24b] Rebecca Martin. Book review: Christopher D.E. Willoughby, Masters of Health: Racial Science and Slavery in U.S. Medical Schools, Chapel Hill: University of North Carolina Press, 2022. Pp. 282. ISBN 978-1-469-67184-0. $99.00 (hardcover). British Journal for the History of Science, 57(3): 492–493, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
christopher-de-willoughby-
masters-of-health-racial-
science-and-slavery-in-us-
medical-schools-chapel-
hill-university-of-north-
carolina-press-2022-pp-282-
isbn-9781469671840-9900-hardcover/
3CC860C3610DC3C5307A47263B5DA070.

Marino:2025:BRM

[Mar25a] Julia E. Marino. Book review: Mario Daniels and John Krige, Knowledge Regulation and National Security in Postwar America, Chicago: University of Chicago Press, 2022. Pp. 432. ISBN 978-0-226-81753-8. $40.00 (paper). British Journal for the History of Science, 58(1): 171–172, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mario-
daniels-and-john-krige-
knowledge-regulation-and-
national-security-in-
postwar-america-chicago-
university-of-chicago-
press-2022-pp-432-isbn-
9780226817538-4000-paper/07BD7BC174FD62A06331750A35F90CE6.

Martinez:2025:BRC

[Mar25b] Jean-Phillipe Martinez. Book review: Climério Paulo da Silva Neto, Materializing the Foundations of Quantum Mechanics: Instruments and the First Bell Tests, Cham: Springer Nature, 2023. Pp. 73. ISBN 978-3-031-29796-0. £39.99 (softcover). British Journal for the History of Science, 58(2):393–394, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/climerio-paulo-da-

silva-neto-materializing-
the-foundations-of-quantum-
mechanics-instruments-and-
the-first-bell-tests-cham-
springer-nature-2023-pp-73-
isbn-9783031297960-3999-softcover/
D3F6979710CE41B90C1BD877AF7A21E6.

Mattes:2022:MIK

[Mat22] Johannes Mattes. Mapping the invisible: knowledge, credibility and visions of earth in early modern cave maps. British Journal for the History of Science, 55(1):53–80, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mapping-
the-invisible-knowledge-
credibility-and-visions-of-
earth-in-early-modern-cave-
maps/
9A626594DED16EB9785EDA54E150EF2B.

Mattes:2025:BRB

[Mat25]  Johannes Mattes.  Book re-

view: Benoît de Maillet (Édition établie, présentée et annotée par Claudine Cohen), Telliamed: Entretiens d’un philosophe indien avec un missionnaire français sur la diminution de la merm Grenoble: Jérôme Millon, 2023. Pp. 408. ISBN 978-2-84137-408-

3. EUR33.00 (paperback). British Journal for the History of Science, 58(2):351–352, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/

journals/british-journal-
for-the-history-of-science/
article/benoit-de-maillet-
edition-etablie-presentee-
et-annotee-par-claudine-
cohen-telliamed-entretiens-
dun-philosophe-indien-avec-
un-missionnaire-francais-
sur-la-diminution-de-la-
mer-grenoble-jerome-millon-2023-
pp-408-isbn-9782841374083-3300-
paperback/C9958E555AC06FFBBC3BD494B2F17777.

Mauuarin:2022:VDS

[Mau22] Anaïs Mauuarin. Visual duplication: specimens, works of art and photographs at the Musée d’ethnographie du Trocadéro (1928–1935). British Journal for the History of Science, 55(3):365–388, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/visual-duplication-
specimens-works-of-art-and-
photographs-at-the-musee-
dethnographie-du-trocadero-
19281935/
0618FC4683D60D7B595E7424D9B36C34.

Murphy:2022:REG

[MB22] Jane H. Murphy and Sahar Bazzaz. Re-examining globalization and the history of science: Ottoman and Middle Eastern experiences. British Journal for the History of Science, 55(4): 411–422, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
reexamining-globalization-
and-the-history-of-science-
ottoman-and-middle-eastern-
experiences/
7EE29C456A52DD9DAF5417C94AD82615.

McAleer:2022:TCW

[McA22] John McAleer. ‘The troubles of collecting’: William Henry Harvey and the practicalities of natural-history collecting in Britain’s nineteenthcentury world. British Journal for the History of Science, 55 (1):81–100, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/troubles-
of-collecting-william-
henry-harvey-and-the-
practicalities-of-
naturalhistory-collecting-
in-britains-
nineteenthcentury-world/
80BD873A0B1A3B2B49C595F12322C7E1.

McCahey:2023:BRN

[McC23] Daniella McCahey. Book review:

Nanna Katrine Lüders Kaalund, Explorations in the Icy North: How Travel Narratives Shaped Arctic Science in the Nineteenth Century. Pittsburgh: University of Pittsburgh Press, 2021. Pp. 230. ISBN 978-0-8229-4659-5. $40.00 (hardback). British Journal for the History of Science, 56 (1):123–124, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/nanna-
katrine-luders-kaalund-
explorations-in-the-icy-
north-how-travel-
narratives-shaped-arctic-
science-in-the-nineteenth-
century-pittsburgh-
university-of-pittsburgh-
press-2021-pp-230-isbn-9780822946595-
4000-hardback/836F6A8B87B59DC1C0A7EF6A5AF90141.

McCray:2025:BRP

[McC25] W. Patrick McCray. Book review: Philip Glahn and Cary Levine, The Future Is Present: Art, Technology, and the Work of Mobile Image, Cambridge, MA: MIT Press, 2024. Pp. 336. ISBN 978-0-262-54807-6. $45.00 (paperback). British Journal for the History of Science, 58 (1):191–192, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/philip-
glahn-and-cary-levine-the-
future-is-present-art-
technology-and-the-work-of-
mobile-image-cambridge-ma-
mit-press-2024-pp-336-isbn-
9780262548076-4500-
paperback/
8394A08F2B1017A75F4CD2D371F6095E.

McGovern:2021:GGD

[McG21] Michael F. McGovern. Genes go digital: Mendelian Inheritance in Man and the genealogy of electronic publishing in biomedicine. British Journal for the History of Science, 54 (2):213–231, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/genes-go-
digital-mendelian-
inheritance-in-man-and-the-
genealogy-of-electronic-
publishing-in-biomedicine/
CFB5307BA6645D9FA9A9634B03DB0F4B.

McMullan:2020:BRS

[McM20] Joshua McMullan. Book review: Sergei Belyakov, Liquidator: The Chernobyl Story. Singapore: World Scientific Publishing, 2019. Pp. 188. ISBN 978-981-322-868-9. £25.00 (paperback). British Journal for the History of Science, 53(4):599–600, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/sergei-belyakov-
liquidator-the-chernobyl-
story-singapore-world-
scientific-publishing-2019-
pp-188-isbn-9789813228689-
2500-paperback/
7378AD3866DDADC45DF7C378170BE252.

Medori:2023:VDC

[Med23] Beatriz Medori. The visual diplomacy of cancer treatments: the mediatic legacy of the Curies in the early transnational fight against cancer. British Journal for the History of Science, 56:167–183, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/visual-
                                        diplomacy-of-cancer-
                                        treatments-the-mediatic-
                                        legacy-of-the-curies-in-
                                        the-early-transnational-
                                        fight-against-cancer/
                                        19E901018D2329B8BFBF6A34B150171C.

Meyer:2020:BRH

[Mey20] Tirza Meyer. Book review: Helen M. Rozwadowski, Vast Expanses: A History of the Oceans. London: Reaktion Books, 2018. Pp. 268. ISBN 978-1-78023-997-2. £16.00 (hardback). British Journal for the History of Science, 53 (1):123–125, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/helen-m-
                                        rozwadowski-vast-expanses-
                                        a-history-of-the-oceans-
                                        london-reaktion-books-2018-
                                        pp-268-isbn-9781780239972-
                                        1600-hardback/
                                        C09B1BC3A65C3AE2607486BB62CE3C93.

Meyer:2022:ILS

[Mey22] Victoria N. Meyer. Innovations from the Levant: smallpox inoculation and perceptions of scientific medicine. British Journal for the History of Science, 55(4):423–444, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/innovations-from-
the-levant-smallpox-
inoculation-and-
perceptions-of-scientific-
medicine/
E3C38F05248CF4393EA2EFD26D743536.

Middleton:2021:BRE

[Mid21] Lindsay Middleton. Book review: Emma Griffin, Bread Winner: An Intimate History of the Victorian Economy. New Haven, CT: Yale University Press, 2020. Pp. 320. ISBN 978-0-300-23006-

2. £20.00 (hardback). British Journal for the History of Science, 54(1):117–118, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/emma-griffin-bread-
winner-an-intimate-history-
of-the-victorian-economy-
new-haven-ct-yale-
university-press-2020-pp-
320-isbn-9780300230062-

2000-hardback/
68EBFE468C62FF294000445E3679235B.

Miller:2021:BRA

[Mil21] Ian Miller. Book review: Andrew B. Liu, Tea War: A History of Capitalism in China and India. New Haven, CT, and London: Yale University Press, 2020. Pp. 360. ISBN 978-0-300-24373-4. £35.00 (hardcover). British Journal for the History of Science, 54(3):391– 392, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/andrew-b-
liu-tea-war-a-history-of-
capitalism-in-china-and-
india-new-haven-ct-and-
london-yale-university-
press-2020-pp-360-isbn-
9780300243734-3500-
hardcover/286A1F6141DC7781B74BD48ED874D5D2.

Morris:2021:BRT

[Mor21a] Andrew M. A. Morris. Book review: Toby Musgrave, The Multifarious Mr Banks: From Botany Bay to Kew, the Natural Historian Who Shaped the World, New Haven, CT: Yale University Press, 2020. Pp. 386. ISBN 978-0-300-22383-5. $35.00 (hardcover). British Journal for the History of Science, 54 (2):245–246, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

journal-for-the-history-of-
science/article/toby-
musgrave-the-multifarious-
mr-banks-from-botany-bay-
to-kew-the-natural-
historian-who-shaped-the-
world-new-haven-ct-yale-
university-press-2020-pp-
386-isbn-9780300223835-
3500-hardcover/2CE89541304076C7522A1B94E23DCB7B.

Morus:2021:FWH

[Mor21b] Iwan Rhys Morus. Out on the fringe: Wales and the history of science. British Journal for the History of Science, 54(1):87–97, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/out-on-the-
fringe-wales-and-the-
history-of-science/
69B7B8BEB81ACCA9796117210D8EB5C0.

Mougey:2023:NNI

[Mou23] Thomas Mougey. Negotiating the norms of an international science: standardization work at the International Geological Congress, 1878–1891. British Journal for the History of Science, 56(4): 435–451, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
negotiating-the-norms-of-
an-international-science-
standardization-work-at-

                                        the-international-
                                        geological-congress-
                                        18781891/
                                        73ADE4A8ADCA0671C8106A816D19E100.

Moxham:2021:BRM

[Mox21] Noah Moxham. Book review: Maria M. Portuondo, The Spanish Disquiet: The Biblical Natural Philosophy of Benito Arias Montano. Chicago and London: The University of Chicago Press, 2019. Pp. xvi + 446. ISBN 978-0-226-59226-8. $65.00 (hardback). British Journal for the History of Science, 54(1):126–128, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/maria-m-portuondo-
                                        the-spanish-disquiet-the-
                                        biblical-natural-
                                        philosophy-of-benito-arias-
                                        montano-chicago-and-london-
                                        the-university-of-chicago-
                                        press-2019-pp-xvi-446-isbn-
                                        9780226592268-6500-hardback/
                                        E80CFCA92FA68434AA51992BFA6C36EB.

Michel:2021:MAD

[MS21] Nicolas Michel and Ivahn Smadja. Mathematics in the archives: deconstructive historiography and the shaping of modern geometry (1837–1852). British Journal for the History of Science, 54(4):423–441, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/

journals/british-journal-
for-the-history-of-science/
article/mathematics-in-the-
archives-deconstructive-
historiography-and-the-
shaping-of-modern-geometry-
18371852/
8C08FCC4DEE5B9C2C846D072808C727F.

Muller:2024:BRO

[Mül24] Janna K. Müller. Book review: Omar W. Nasim, The Astronomer’s Chair: A Visual and Cultural History, Cambridge, MA: MIT Press, 2021. Pp. 312. ISBN 978-0-262-04553-7. $60.00 (paperback). British Journal for the History of Science, 57 (1):131–132, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/omar-w-
nasim-the-astronomers-
chair-a-visual-and-
cultural-history-cambridge-
ma-mit-press-2021-pp-312-
isbn-9780262045537-6000-
paperback/
95A8C2A09B8CA466A70DE2DB953026E2.

Morgan:2025:BRL

[MWH25] Gregory J.

Morgan, Réka Weinacht, and Madison Heller. Book review: Laura Mamo, Sexualizing Cancer: HPV and the Politics of Cancer Prevention, Chicago: University of Chicago Press, 2024. Pp. 360. ISBN 978-0-226-82929-

6. $35.00 (paper). British Journal for the History of Science, 58(2):357–358, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/laura-mamo-
                                        sexualizing-cancer-hpv-and-
                                        the-politics-of-cancer-
                                        prevention-chicago-
                                        university-of-chicago-
                                        press-2024-pp-360-isbn-
                                        9780226829296-3500-paper/
                                        439C7411B0E53DFFA99965D5A5ACDB82.

Niala:2024:BRA

[Nia24]  J. C. Niala.  Book review:

Anna Marie Roos and Vera Keller (eds), Collective Wisdom: Collecting in the Early Modern Academy, Turnhout: Brepols, 2022. Pp. 325. ISBN 978-2-503-58806-3. EUR85.00 (hardback). British Journal for the History of Science, 57(1): 136–138, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/anna-marie-
                                        roos-and-vera-keller-eds-
                                        collective-wisdom-
                                        collecting-in-the-early-
                                        modern-academy-turnhout-
                                        brepols-2022-pp-325-isbn-
                                        9782503588063-8500-
                                        hardback/4D8443581C2EB4C6BE5520A1536718F7.

Nichols:2022:CDO

[Nic22] Catherine A. Nichols. Curating duplicates: operationalizing similarity in the Smithsonian Institution with Haida rattles, 1880– 1926. British Journal for the History of Science, 55(3):341– 363, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/curating-
duplicates-
operationalizing-
similiarity-in-the-
smithsonian-institution-
with-haida-rattles-
18801926/
845161923F4D901FB9388C4E723A2435.

Nixon:2024:WBO

[Nix24] Sean Nixon. Watching birds: observation, photography and the ‘ethological eye’. British Journal for the History of Science, 57(1):1–19, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/watching-
birds-observation-
photography-and-the-
ethological-eye/
BCFB24B7609DFB5A113C34F5D46E87ED.

Nutton:2024:BRC

[Nut24] Vivian Nutton. Book review:

Colin Webster, Tools and Organisms: Technology and the Body in Ancient Greek and Roman Medicine, Chicago: University of Chicago Press, 2023. Pp. 320. ISBN 978-0-226-82877-

0. $45.00 (cloth). British Journal for the History of Science, 57 (4):665–666, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/colin-
                                        webster-tools-and-
                                        organisms-technology-and-
                                        the-body-in-ancient-greek-
                                        and-roman-medicine-chicago-
                                        university-of-chicago-
                                        press-2023-pp-320-isbn-
                                        9780226828770-4500-cloth/
                                        946273780200CD9E61740DCF7A1AF4A5.

Offerhaus:2025:BRM

[Off25]  Aleida Offerhaus.  Book re-

view: Maria Eulàlia Gassó Miracle, Coenraad Jacob Temminck and the Emergence of Systematics (1800–1850) Leiden: Brill, 2021. Pp. 328. ISBN 978-90-04-41917-

9. EUR129.00 (hardback). British Journal for the History of Science, 58(1):157–158, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/maria-eulalia-
                                        gasso-miracle-coenraad-
                                        jacob-temminck-and-the-
                                        emergence-of-systematics-
                                        18001850-leiden-brill-2021-
                                        pp-328-isbn-9789004419179-
                                        12900-hardback/
                                        7F690C36A1F397F22A6938E47AFA8125.

Olshin:2023:BRA

[Ols23] Benjamin B. Olshin. Book review: Alida C. Metcalf, Mapping an Atlantic World, circa 1500. Baltimore: Johns Hopkins University Press, 2020. Pp. 224. ISBN 978-1-4214-3852-8. $57.00 (hardback). British Journal for the History of Science, 56(4): 597–598, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/alida-c-
metcalf-mapping-an-
atlantic-world-circa-1500-
baltimore-johns-hopkins-
university-press-2020-pp-
224-isbn-9781421438528-
5700-hardback/
4A5C1785AF591FEB4F061F64FB3704AB.

OMalley:2025:BRM

[O’M25] Maureen A. O’Malley. Book review: Mark A. Ragan, Kingdoms, Empires, and Domains: The History of High-Level Biological Classification, Oxford: Oxford University Press, 2023. Pp. 816. ISBN 978-0-19-764303-7. £107.50 (hardback). British Journal for the History of Science, 58(1): 159–160, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mark-a-
ragan-kingdoms-empires-and-
domains-the-history-of-
highlevel-biological-
classification-oxford-
oxford-university-press-

                                        2023-pp-816-isbn-
                                        9780197643037-10750-
                                        hardback/0261005FCE343494ACC2A05DFD7DD939.

Omodeo:2024:CCN

[Omo24] Pietro Daniel Omodeo. Cesare Cremonini’s non-theological cosmology: a contribution to Padua’s secular culture in times of wars of religion. British Journal for the History of Science, 57(4):561–579, December 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/cesare-cremoninis-
                                        nontheological-cosmology-a-
                                        contribution-to-paduas-
                                        secular-culture-in-times-
                                        of-wars-of-religion/
                                        7AAF6AC871BBC5E4D4F3A570973E8C4A.

Palmieri:2021:BRA

[Pal21] Kristine Palmieri. Book reviews: Alex Csiszar, The Scientific Journal: Authorship and the Politics of Knowledge in the Nineteenth Century Cambridge, MA: Harvard University Press, 2018. Pp. 368. ISBN 978-0-226-55323-8. $45.00 (hardback). ISBN 978-0-226-75250-1. $35.00 (paperback). British Journal for the History of Science, 54 (2):241–242, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/alex-

csiszar-the-scientific-
journal-authorship-and-the-
politics-of-knowledge-in-
the-nineteenth-century-
cambridge-ma-harvard-
university-press-2018-pp-
368-isbn-9780226553238-4500-
hardback-isbn-9780226752501-
3500-paperback/EED2D4E3ED6727E73F7114024795DC1D.

Penn:2024:BRC

[Pen24] Jonnie Penn. Essay review: Computing as sugar? The sweet and the bitter of social histories of computing. Janet Abbate and Stephanie Dick (eds.), Abstractions and Embodiments: New Histories of Computing and Society, Baltimore: Johns Hopkins University Press, 2022. Pp. 472. ISBN 978-1-4214-4437-6. $39.95 (paperback). British Journal for the History of Science, 57(3): 479–483, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/computing-
as-sugar-the-sweet-and-the-
bitter-of-social-histories-
of-computing/
8C2B60E952E3FA0CEFABEF3A3057EB03.

Pfeffer:2021:SAC

[Pfe21] Michelle Pfeffer. The Society of Astrologers (c. 1647–1684): sermons, feasts and the resuscitation of astrology in seventeenthcentury London. British Journal for the History of Science, 54(2):133–153, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/society-of-
                                        astrologers-c16471684-
                                        sermons-feasts-and-the-
                                        resuscitation-of-astrology-
                                        in-seventeenthcentury-
                                        london/
                                        B292AB39DACD6474CBAB1000BAE4E655.

Piel:2020:SBS

[Pie20] Helen Piel. Scientific broadcasting as a social responsibility? John Maynard Smith on radio and television in the 1960s and 1970s. British Journal for the History of Science, 53 (1):89–108, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/scientific-
                                        broadcasting-as-a-social-
                                        responsibility-john-
                                        maynard-smith-on-radio-and-
                                        television-in-the-1960s-
                                        and-1970s/
                                        6F06499A7EDA8C0282E56DB79BA95F9A.

Raj:2024:WSM

[Raj24]  Kapil Raj.  When Scot-

tish medicine hospitalized Indian magic: Dr James Esdaile’s mesmeric surgery in mid-nineteenth-century Bengal. British Journal for the History of Science, 57(3):325–347, September 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.

cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/when-scottish-
medicine-hospitalized-
indian-magic-dr-james-
esdailes-mesmeric-surgery-
in-midnineteenthcentury-
bengal/
DCAE917EB83F11C5503708A41D6FC68E.

Ramberg:2024:BRG

[Ram24] Peter J. Ramberg. Book reviews: Gisela Boeck and Alan J. Rocke, Lothar Meyer: Modern Theories and Pathways to Periodicity, Cham: Birkhäuser, 2022. Pp. xi + 193. ISBN 978-0-303-78341-9. £79.99 (hardcover). — Gisela Boeck and Alan J. Rocke, Lothar Meyer: Moderne Theorien und Wege zum Periodensystem, Berlin: Springer Spektrum, 2022. Pp. ix + 217. ISBN 978-3-662-63932-0. £79.99 (softcover). British Journal for the History of Science, 57(3):499– 501, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/gisela-
boeck-and-alan-j-rocke-
lothar-meyer-modern-
theories-and-pathways-to-
periodicity-cham-
birkhauser-2022-pp-xi-193-
isbn-9780303783419-7999-
hardcover-gisela-boeck-and-
alan-j-rocke-lothar-meyer-moderne-
theorien-und-wege-zum-periodensystem-
berlin-springer-spektrum-2022-

                                        pp-ix-217-isbn-9783662639320-
                                        7999-softcover/3298221CFF5B02299F88FCB08D3BAD9B.

Ranford:2023:BRA

[Ran23] Paul Ranford. Book review: Aileen Fyfe, Noah Moxham, Julie McDougal-Waters and Camilla Mørk Røstvik, A History of Scientific Journals: Publishing at the Royal Society, 1665–2015. London: UCL Press, 2022. Pp. 643. ISBN 978-1-8000-8234-2. £60.00 (hardcover), £0.00 (open-access pdf). Doi:10.14324/111.9781800082328. British Journal for the History of Science, 56(3):416–418, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/aileen-fyfe-noah-
                                        moxham-julie-
                                        mcdougalwaters-and-camilla-
                                        mork-rostvik-a-history-of-
                                        scientific-journals-
                                        publishing-at-the-royal-
                                        society-16652015-london-ucl-
                                        press-2022-pp-643-isbn-9781800082342-
                                        6000-hardcover-000-openaccess-
                                        pdf-doi10143241119781800082328/
                                        0C60483D34C81972D3188340AB6D81AD.

Robinson:2020:ICR

[RBB^{+}20] Sam

Robinson, Megan Baumhammer, Lea Beiermann, Daniel Belteki, Amy C. Chambers, Kelcey Gibbons, Edward Guimont, Kathryn Heffner, Emma-Louise Hill, Jemma Houghton, Daniella Mccahey, Sarah Qidwai, Charlotte Sleigh, Nicola Sugden, and James Sumner. Innovation in a crisis: rethinking conferences and scholarship in a pandemic and climate emergency. British Journal for the History of Science, 53(4): 575–590, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/innovation-
in-a-crisis-rethinking-
conferences-and-
scholarship-in-a-pandemic-
and-climate-emergency/
601948CB0B684F833AC3E96643E46F2C.

Reeves:2021:BRT

[Ree21] Carole Reeves. Book review: Tom Sharpe, The Fossil Woman: A Life of Mary Anning Wimborne, Minster: The Dovecote Press, 2020. Pp. 240. ISBN 978-0-9955462-9-5. £20.00 (hardback). British Journal for the History of Science, 54(2):250–251, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/tom-sharpe-
the-fossil-woman-a-life-of-
mary-anning-wimborne-
minster-the-dovecote-press-
2020-pp-240-isbn-
9780995546295-2000-
hardback/
3095C57F79274F868CC99B7AD6BFC2BF.

Reeves:2025:BRN

[Ree25] Eileen Reeves. Book review: Nicole Howard, Loath to Print: The Reluctant Scientific Author 1500–1750, Baltimore: Johns Hopkins University Press, 2022. Pp. 232. ISBN 978-1-4214-4368-3. $55.00 (hardcover). British Journal for the History of Science, 58(2):381–382, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/nicole-
                                        howard-loath-to-print-the-
                                        reluctant-scientific-
                                        author-15001750-baltimore-
                                        johns-hopkins-university-
                                        press-2022-pp-232-isbn-
                                        9781421443683-5500-
                                        hardcover/
                                        1D8095AAB55BA574FC56144BBC5BB775.

Reinisch:2023:TCT

[Rei23] Jessica Reinisch. Technical conferences as a technique of internationalism. British Journal for the History of Science, 56(4): 485–502, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/technical-
                                        conferences-as-a-technique-
                                        of-internationalism/
                                        72E819978160AF024A4D37CFA7C3A83F.

Revuelta:2025:BSA

[Rev25] Gloria Maritza Gómez Revuelta.

Black in space: Arnaldo Tamayo and the Cuban cosmic revolution. British Journal for the History of Science, 58(2):303–324, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/black-in-space-
arnaldo-tamayo-and-the-
cuban-cosmic-revolution/
4B1D250F90C2B85C82D8FE8C1107768E.

Reese:2025:VKR

[RFV25] Henry Reese, Vanessa Finney, and Simon Ville. Value, knowledge and reputation: zoological exchange by Australian museums, 1870–1900. British Journal for the History of Science, 58(1):81–100, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/value-
knowledge-and-reputation-
zoological-exchange-by-
australian-museums-
18701900/
69663791EAA785CDEAE08D6F18A812CE.

Rispoli:2020:BRN

[Ris20] Giulia Rispoli. Book review: Nicholas B. Breyfogle (ed.), Eurasian Environments: Nature and Ecology in Imperial Russian and Soviet History. Pittsburgh: University of Pittsburgh Press, 2018. Pp. 424. ISBN: 978-0-8229-6563-3. $34.95 (paperback). British Journal for the History of Science, 53 (1):134–136, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/nicholas-b-
breyfogle-ed-eurasian-
environments-nature-and-
ecology-in-imperial-
russian-and-soviet-history-
pittsburgh-university-of-
pittsburgh-press-2018-pp-
424-isbn-9780822965633-3495-
paperback/3560B54516BF0D92C6EC2C7844BF9390.

Roberti:2025:BRK

[Rob25] Valentina Roberti. Book review: Kenneth L. Caneva, Helmholtz and the Conservation of Energy: Contexts of Creation and Reception, Cambridge, MA: MIT Press, 2021. Pp. 760. ISBN 978-0-262-04573-5. $125.00 (hardcover). British Journal for the History of Science, 58(1): 161–162, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/kenneth-l-
caneva-helmholtz-and-the-
conservation-of-energy-
contexts-of-creation-and-
reception-cambridge-ma-mit-
press-2021-pp-760-isbn-
9780262045735-12500-
hardcover/82CE5A489ABCF9691DC8FF3DF08EA5BA.

Rocha:2024:BRE

[Roc24] Gustavo Rodrigues Rocha. Book review: Efran Sera-Shriar, Psychic Investigators: Anthropology, Modern Spiritualism, and Credible Witnessing in the Late Victorian Age, Pittsburgh: University of Pittsburgh Press, 2022. Pp. 236. ISBN 978-0-8229-4707-3. $50.00 (hardcover). British Journal for the History of Science, 57 (1):161–163, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/efran-
serashriar-psychic-
investigators-anthropology-
modern-spiritualism-and-
credible-witnessing-in-the-
late-victorian-age-
pittsburgh-university-of-
pittsburgh-press-2022-pp-
236-isbn-9780822947073-5000-
hardcover/67D02E2AD040F5E2C8373837CC63D59B.

Rossini:2020:BRR

[Ros20] Paolo Rossini. Book review:

Richard J. Oosterhoff, Making Mathematical Culture: University and Print in the Circle of Lefèvre d’Etaples, Oxford: Oxford University Press, 2018. Pp. 304. ISBN 978-0-19-882352-0. £65.00 (hardback). British Journal for the History of Science, 53 (2):282–283, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

                                        journal-for-the-history-of-
                                        science/article/richard-j-
                                        oosterhoff-making-
                                        mathematical-culture-
                                        university-and-print-in-
                                        the-circle-of-lefevre-
                                        detaples-oxford-oxford-
                                        university-press-2018-pp-
                                        304-isbn-9780198823520-
                                        6500-hardback/03675126FF50009A9FA3E88FDFEDF381.

Rose:2022:BRT

[Ros22] Sebastian James Rose. Book review: Thomas Simpson, The Frontier in British India: Space, Science and Power in the Nineteenth Century. Cambridge: Cambridge University Press, 2021. Pp. xvi + 298. ISBN 978-1-108-87915-

6. £75.00 (hardback). British Journal for the History of Science, 55(2):250–252, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/thomas-simpson-the-
                                        frontier-in-british-india-
                                        space-science-and-power-in-
                                        the-nineteenth-century-
                                        cambridge-cambridge-
                                        university-press-2021-pp-
                                        xvi-298-isbn-9781108879156-7500-
                                        hardback/0B9213454159BD3937E12A3324C28217.

Rossetter:2023:AVN

[Ros23] Thomas Rossetter. Anti-voluntarism, natural providence and miracles in Thomas Burnet’s theory of the Earth. British Journal for the History of Science, 56(1):1–20, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/antivoluntarism-
natural-providence-and-
miracles-in-thomas-burnets-
theory-of-the-earth/
029A224C155D5BCBD774177896007B40.

Rossetter:2025:BRV

[Ros25] Thomas Rossetter. Book review: Vera Keller, The Interlopers: Early Stuart Projects and the Undisciplining of Knowledge, Baltimore: Johns Hopkins University Press, 2023. Pp. 368. ISBN 978-1-4214-4592-2. $60.00 (hardcover). British Journal for the History of Science, 58 (1):145–146, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/vera-
keller-the-interlopers-
early-stuart-projects-and-
the-undisciplining-of-
knowledge-baltimore-johns-
hopkins-university-press-
2023-pp-368-isbn-
9781421445922-6000-
hardcover/561180EFF8205D66D5F8DDD60356E0AB.

Sachdev:2021:BRD

[Sac21] Vorathep Sachdev. Book review: David Herzberg, White Market Drugs: Big Pharma and the Hidden History of Addiction in America, Chicago: The University of Chicago Press, 2020. ISBN 978-0-226-73188-9. $27.50 (hardback). British Journal for the History of Science, 54 (2):252–253, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/david-
                                        herzberg-white-market-
                                        drugs-big-pharma-and-the-
                                        hidden-history-of-
                                        addiction-in-america-
                                        chicago-the-university-of-
                                        chicago-press-2020-isbn-
                                        9780226731889-2750-
                                        hardback/172E9835A008956F1E5C2FDE05377EA8.

Skogerboe:2025:PBA

[SB25]  David  Skogerboe

and David Baneke. The prophet business: Arthur C. Clarke, Sri Lanka and the making of a global space persona. British Journal for the History of Science, 58 (2):215–236, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/prophet-
                                        business-arthur-c-clarke-
                                        sri-lanka-and-the-making-
                                        of-a-global-space-persona/
                                        5D67DD2B87CFE3E0692F219019D08BD0.

Schleper:2022:CCT

[Sch22a] Simone Schleper. Caribou crossings: the Trans-Alaska Pipeline System, conservation, and stakeholdership in the Anthropocene. British Journal for the History of Science, 55(2):127–143, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/caribou-crossings-
the-transalaska-pipeline-
system-conservation-and-
stakeholdership-in-the-
anthropocene/
B1CB3827444C3D4CA70654544D59F217.

Schmechel:2022:DFD

[Sch22b] Carmen Schmechel. Descartes on fermentation in digestion: iatromechanism, analogy and teleology. British Journal for the History of Science, 55(1): 101–116, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/descartes-
on-fermentation-in-
digestion-iatromechanism-
analogy-and-teleology/
7E372D66D0B035A876A7FB92C0A0A525.

Schirrmacher:2025:BRM

[Sch25a] Arne Schirrmacher. Book reviews: Mitchell G. Ash, Die Max-Planck-Gesellschaft im Prozess der deutschen Vereinigung 1989– 2002, Göttingen: Vandenhoeck & Ruprecht, 2023. Pp. 394. ISBN 978-3-525-30209-5. EUR70.00 (hardcover). ISBN 978-3-666-30209-1. EUR0.00 (open-access PDF). British Journal for the History of Science, 58(1):173–174, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/mitchell-g-ash-die-
maxplanckgesellschaft-im-
prozess-der-deutschen-
vereinigung-19892002-
gottingen-vandenhoeck-
ruprecht-2023-pp-394-isbn-9783525302095-
7000-hardcover-isbn-9783666302091-
000-openaccess-pdf/4BC08CF1B0DAA285B6093F05B0350B78.

Schley:2025:BRA

[Sch25b] David Schley. Book review: Aaron W. Marrs, The American Transportation Revolution: A Social and Cultural History, Baltimore: Johns Hopkins University Press, 2024. Pp. 262. ISBN 978-1-4214-4849-7. $64.95 (hardcover). British Journal for the History of Science, 58(2):395–396, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/aaron-w-marrs-the-
american-transportation-
revolution-a-social-and-
cultural-history-baltimore-
johns-hopkins-university-
press-2024-pp-262-isbn-
9781421248497-6495-
hardcover/8C9A95B4EDFF5CD7D46B39F222499149.

Schuetz:2025:BRP

[Sch25c] Thomas Schuetz. Book review:

Peter Thompson, The Gas Mask in Interwar Germany, Cambridge: Cambridge University Press, 2023. Pp. 320. ISBN 978-1-009-31486-2. $130.00 (hardback). British Journal for the History of Science, 58(2):397–398, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/peter-
thompson-the-gas-mask-in-
interwar-germany-cambridge-
cambridge-university-press-
2023-pp-320-isbn-
9781009314862-13000-
hardback/
F2AB3C4A92DD2A5D08AE2072541F6090.

Suarez-Diaz:2024:BRJ

[SD24] Edna Suárez-Díaz. Book review: J. Justin Castro and James A. Garza (eds), Technocratic Visions: Engineers, Technology, and Society in Mexico, Pittsburgh: University of Pittsburgh Press, 2022. Pp. 282. ISBN: 978-0-8229-4748-6. $55.00 (cloth). British Journal for the History of Science, 57(3): 503–504, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/j-justin-
castro-and-james-a-garza-

                                        eds-technocratic-visions-
                                        engineers-technology-and-
                                        society-in-mexico-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2022-pp-
                                        282-isbn-9780822947486-5500-
                                        cloth/F210366CCFF46BBB535C9119F146FFD2.

Secord:2021:RHDa

[Sec21a] James A. Secord. Revolutions in the head: Darwin, Malthus and Robert M. Young. British Journal for the History of Science, 54(1):41–59, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/revolutions-in-the-
                                        head-darwin-malthus-and-
                                        robert-m-young/
                                        9DBEACA1949208DD36BF77C8653DF21E.

See corrigendum [Sec21b].

Secord:2021:RHDb

[Sec21b] James A. Secord. Revolutions in the head: Darwin, Malthus and Robert M. Young — corrigendum. British Journal for the History of Science, 54(2):255, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/revolutions-in-the-
                                        head-darwin-malthus-and-
                                        robert-m-young-corrigendum/
                                        073D3E929F2BB8B5C21C8F62A8EFCB63.

See [Sec21a].

Secord:2022:BRD

[Sec22] James A. Secord. Book review:

Diarmid A. Finnegan, The Voice of Science: British Scientists on the Lecture Circuit in Gilded Age America, Pittsburgh: University of Pittsburgh Press, 2021. Pp. xiii + 286. ISBN 978-0-8229-4681-6. $60.00 (hardback). British Journal for the History of Science, 55 (1):117–119, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/diarmid-a-
finnegan-the-voice-of-
science-british-scientists-
on-the-lecture-circuit-in-
gilded-age-america-
pittsburgh-university-of-
pittsburgh-press-2021-pp-
xiii-286-isbn-
9780822946816-6000-hardback/
8A582F1D72E0A337B17DCB3818CDAB28.

Secord:2025:TA

[Sec25] James A. Secord. A tradition from the ancestors. British Journal for the History of Science, 58(1):1–17, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/tradition-
from-the-ancestors/
9159D7D4D8D6C1628251B5B898C976FC.

Shalvey:2022:BRA

[Sha22] Aisling Shalvey. Book review:

Amir Teicher, Social Mendelism: Genetics and the Politics of Race in Germany, 1900–1948. Cambridge: Cambridge University Press, 2021. Pp. xiv + 268. ISBN 978-1-108-49949-1. £26.99 (hardback). British Journal for the History of Science, 55 (2):252–253, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/amir-
teicher-social-mendelism-
genetics-and-the-politics-
of-race-in-germany-
19001948-cambridge-
cambridge-university-press-
2021-pp-xiv-268-isbn-
9781108499491-2699-
hardback/86E83B483AFB11EFB7174BBCC0B2BD06.

Shields:2024:CHB

[Shi24]  Brit Shields.  The ‘Courant

Hilton’: building the mathematical sciences at New York University. British Journal for the History of Science, 57(3): 425–446, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/courant-
hilton-building-the-
mathematical-sciences-at-
new-york-university/
A676B4AD7FBCAB05B767BB5FEC4DA9FD.

Short:2022:BRA

[Sho22] Devin Short. Book review: Alex Wellerstein, Restricted Data: The History of Nuclear Secrecy in the United States, Chicago: The University of Chicago Press, 2021. Pp. 528. ISBN 978-0-2260-2038-9. $35.00 (hardback). British Journal for the History of Science, 55(4):525–527, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/alex-wellerstein-
restricted-data-the-
history-of-nuclear-secrecy-
in-the-united-states-
chicago-the-university-of-
chicago-press-2021-pp-528-
isbn-9780226020389-3500-
hardback/F090D1B68D646FCABA0F6BBB42A8FA06.

Sialaros:2021:BRA

[Sia21] Michalis Sialaros. Book review:

Alexander Jones and Liba Taub (eds.), The Cambridge History of Science, vol. 1, Ancient Science Cambridge: Cambridge University Press, 2018. Pp. xix + 642. ISBN 978-0-511-98014-5. £120.00 (hardback). British Journal for the History of Science, 54(1): 124–125, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/alexander-
jones-and-liba-taub-eds-
the-cambridge-history-of-
science-vol-1-ancient-
science-cambridge-

                                        cambridge-university-press-
                                        2018-pp-xix-642-isbn-
                                        9780511980145-12000-
                                        hardback/FE05FCF78E3727110F679FC440C82CA5.

Siddiqi:2025:EMWa

[Sid25a] Asif Siddiqi. Epilogue: the many worlds of rock(et) stars. British Journal for the History of Science, 58(2):325–333, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/epilogue-the-many-
                                        worlds-of-rocket-stars/
                                        4E1804F5A0172DC69920785E7DB77D6C.

See corrigendum [Sid25b].

Siddiqi:2025:EMWb

[Sid25b] Asif Siddiqi. Epilogue: the many worlds of rock(et) stars

— corrigendum. British Journal for the History of Science, 58(2):403, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/epilogue-
                                        the-many-worlds-of-rocket-
                                        stars-corrigendum/
                                        4EDFE73AFFDDDA572111534E09F8768A.

See [Sid25a].

Siebeneichner:2025:SGS

[Sie25a] Tilmann Siebeneichner. Showcasing Germany in space: the lives and afterlives of Cold War rocket stars Sigmund Jähn and Ulf Merbold. British Journal for the History of Science, 58 (2):281–302, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/showcasing-
germany-in-space-the-lives-
and-afterlives-of-cold-war-
rocket-stars-sigmund-jahn-
and-ulf-merbold/
49359D84DEFBCB711A4753BE8839BC1E.

Siegel:2025:BRT

[Sie25b] Benjamin Siegel. Book review:

Thomas B. Robertson and Jenny Leigh Smith (eds.), Transplanting Modernity: The Environmental Legacy of International Development, Pittsburgh: University of Pittsburgh Press, 2023. Pp. 320. ISBN 978-0-8229-4639-7. $50.00 (hardcover). British Journal for the History of Science, 58 (1):175–176, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/thomas-b-
robertson-and-jenny-leigh-
smith-eds-transplanting-
modernity-the-
environmental-legacy-of-
international-development-
pittsburgh-university-of-
pittsburgh-press-2023-pp-
320-isbn-9780822946397-5000-
hardcover/71BB5C161361910BE4A6412DEB5B0124.

Sime:2022:BRD

[Sim22] Zane Sime. Book review: Dipesh Chakrabarty, The Climate of History in a Planetary Age. Chicago: The University of Chicago Press, 2021. Pp. 296. ISBN 978-0-226-73286-2. $25.00 (paperback). British Journal for the History of Science, 55(2):247–248, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/dipesh-
                                        chakrabarty-the-climate-of-
                                        history-in-a-planetary-age-
                                        chicago-the-university-of-
                                        chicago-press-2021-pp-296-
                                        isbn-9780226732862-2500-
                                        paperback/
                                        18775C1057A57BA5A145286D5A3F0CFA.

Sleigh:2020:BRJ

[Sle20]  Charlotte Sleigh.  Book re-

view: Joshua Nall, News from Mars: Mass Media and the Forging of a New Astronomy, 1860– 1910. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 287. ISBN 978-0-8229-4552-9. $50.00 (hardcover). British Journal for the History of Science, 53 (2):281–282, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/joshua-
                                        nall-news-from-mars-mass-
                                        media-and-the-forging-of-a-

new-astronomy-18601910-
pittsburgh-university-of-
pittsburgh-press-2019-pp-
287-isbn-9780822945529-
5000-hardcover/
F1CAFB3BB21EAF146C16937F14590BB2.

Sleigh:2024:BRK

[Sle24]  Charlotte Sleigh.  Book re-

view: Karel Capek (ed. Jitka Cejková), R.U.R. and the Vision of Artificial Life Cambridge, MA: MIT Press, 2023. Pp. 312. ISBN 978-0-262-54450-4. $29.95 (paperback). British Journal for the History of Science, 57(3): 501–502, September 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/karel-
capek-ed-jitka-cejkova-rur-
and-the-vision-of-
artificial-life-cambridge-
ma-mit-press-2023-pp-312-
isbn-9780262544504-2995-
paperback/
D47F6C0AF5D8621B2E4024A9BCB41290.

Smith:2020:BRK

[Smi20] Roger Smith. Book review: Kerri Andrews, Wanderers: A History of Women Walking. London: Reaktion Books, 2020. Pp. 303. ISBN 978-1-78914-342-3. £14.99 (hardback). British Journal for the History of Science, 53(4): 602–603, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-

                                        journal-for-the-history-of-
                                        science/article/kerri-
                                        andrews-wanderers-a-
                                        history-of-women-walking-
                                        london-reaktion-books-2020-
                                        pp-303-isbn-9781789143423-
                                        1499-hardback/
                                        5485980149031EE578A093D9DA06771B.

Smith:2021:MPC

[Smi21a] Crosbie Smith. Maurice Pierre Crosland (1931–2020): an appreciation. British Journal for the History of Science, 54(1):79–85, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/maurice-pierre-
                                        crosland-19312020-an-
                                        appreciation/
                                        58825CA57DD603492B21B855A4410F15.

Smith:2021:BRE

[Smi21b] David F. Smith. Book review: Elizabeth A. Williams, Appetite and Its Discontents: Science, Medicine and the Urge to Eat, 1750–1950. Chicago: The University of Chicago Press, 2020. Pp. 416. ISBN 978-0-226-69304-0. $35.00 (paperback). British Journal for the History of Science, 54 (1):128–129, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/elizabeth-
                                        a-williams-appetite-and-
                                        its-discontents-science-

medicine-and-the-urge-to-
eat-17501950-chicago-the-
university-of-chicago-
press-2020-pp-416-isbn-
9780226693040-3500-
paperback/751ED7088372EA7728CF388E24CAAAA2.

Smith:2023:BRR

[Smi23]  Roger Smith.  Book review:

Robert J. Sternberg and Wade E. Pickren (eds.), The Cambridge Handbook of the Intellectual History of Psychology, Cambridge: Cambridge University Press, 2019. Pp. 533. ISBN 978-1-108-41869-0. $210.00 (hardback). British Journal for the History of Science, 56(3):418–419, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/robert-j-sternberg-
and-wade-e-pickren-eds-the-
cambridge-handbook-of-the-
intellectual-history-of-
psychology-cambridge-
cambridge-university-press-
2019-pp-533-isbn-
9781108418690-21000-hardback/
63BC8389BE66B42918D52FC046D86545.

Smith:2024:BRM

[Smi24] Roger Smith. Book review: Margarete Vöhringer, Avant-Garde and Psychotechnics: Science, Art and Technology in the Early Soviet Union, London: Routledge, 2023. Pp. 254. ISBN 978-1-032-53264-6. £104.00 (hardback). British Journal for the History of Science, 57(1):

133–134, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/margarete-
                                        vohringer-avantgarde-and-
                                        psychotechnics-science-art-
                                        and-technology-in-the-
                                        early-soviet-union-london-
                                        routledge-2023-pp-254-isbn-
                                        9781032532646-10400-
                                        hardback/8022A3E305194225A8747BA104FA1E95.

Somsen:2023:GWS

[Som23] Geert Somsen. ‘The goddess that we serve’: projecting international community at the first serial chemistry conferences, 1893–1914. British Journal for the History of Science, 56(4): 453–467, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/goddess-
                                        that-we-serve-projecting-
                                        international-community-at-
                                        the-first-serial-chemistry-
                                        conferences-18931914/
                                        53349DC11128CC0F6CA961B59D24641D.

Spary:2025:BRL

[Spa25] E. C. Spary. Book review: Leonie Hannan, A Culture of Curiosity: Science in the Eighteenth-Century Home, Manchester: Manchester University Press, 2023. Pp. 272. ISBN 978-1-5261-5303-6. £85.00 (hardcover). British Journal for the History of Science, 58 (1):147–148, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/leonie-
hannan-a-culture-of-
curiosity-science-in-the-
eighteenthcentury-home-
manchester-manchester-
university-press-2023-pp-
272-isbn-9781526153036-
8500-hardcover/142543AB3F9004AC9E77E8A27FBD9F88.

Spinardi:2020:GAN

[Spi20] Graham Spinardi. The green airliner that never was: aerodynamic theory, fuel-efficiency and the role of the British state in aviation technology in the midtwentieth century. British Journal for the History of Science, 53(2):229–254, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/green-
airliner-that-never-was-
aerodynamic-theory-
fuelefficiency-and-the-
role-of-the-british-state-
in-aviation-technology-in-
the-midtwentieth-century/
62F9ABAB856994206E46E389CA0FF08A.

Spielmann:2023:BRP

[Spi23] Sophia Spielmann. Book review: Paula De Vos, Compound Remedies: Galenic Pharmacy from the Ancient Mediterranean to New Spain. Pittsburgh: University of Pittsburgh Press, 2021. Pp. 404. ISBN 978-0-8229-4649-6 $50.00 (hardback). British Journal for the History of Science, 56:275–276, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/paula-de-
                                        vos-compound-remedies-
                                        galenic-pharmacy-from-the-
                                        ancient-mediterranean-to-
                                        new-spain-pittsburgh-
                                        university-of-pittsburgh-
                                        press-2021-pp-404-isbn-
                                        9780822946496-5000-
                                        hardback/B7B096A7041F9ECA92403AC92F4DBE59.

Sera-Shriar:2023:MDF

[SS23] Efram Sera-Shriar. From museumization to decolonization: fostering critical dialogues in the history of science with a Haida eagle mask. British Journal for the History of Science, 56(3): 309–328, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/from-
                                        museumization-to-
                                        decolonization-fostering-
                                        critical-dialogues-in-the-
                                        history-of-science-with-a-
                                        haida-eagle-mask/
                                        E6B3D4C885685F0A1843F7A099CCE380.

Stark:2022:RDD

[Sta22] James F. Stark. ‘A remedy for this dread disease’: Achille Sclavo, anthrax and serum therapy in early twentieth-century Britain. British Journal for the History of Science, 55(2): 207–226, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/remedy-for-
this-dread-disease-achille-
sclavo-anthrax-and-serum-
therapy-in-early-
twentiethcentury-britain/
DBFFB0865A244BB7B34C290BB08FD690.

Staley:2024:BRD

[Sta24a] Richard Staley. Book review:

Dipesh Chakrabarty, One Planet, Many Worlds: The Climate Parallax, Waltham, MA: Brandeis University Press, 2023. Pp. 144. ISBN: 978-1-68458-157-3. $24.95 (paper). British Journal for the History of Science, 57(4): 657–659, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/dipesh-
chakrabarty-one-planet-
many-worlds-the-climate-
parallax-waltham-ma-
brandeis-university-press-
2023-pp-144-isbn-
9781684581573-2495-paper/
A32BB67FB0C263DD275F0791AC5E92B8.

Stark:2024:BRL

[Sta24b] James F. Stark. Book review: Lina Zeldovich, The Other Dark Matter: The Science and Business of Turning Waste into Wealth and Health, Chicago: University of Chicago Press, 2021. Pp. 259. ISBN 978-0-226-61557-

8. $26.00 (cloth). British Journal for the History of Science, 57 (1):152–153, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/lina-
zeldovich-the-other-dark-
matter-the-science-and-
business-of-turning-waste-
into-wealth-and-health-
chicago-university-of-
chicago-press-2021-pp-259-
isbn-9780226615578-2600-
cloth/5D50F78AE55071025882A06F97AE5BBD.

Stenhouse:2022:BRA

[Ste22] Brigitte Stenhouse. Book reviews: Allan Esterson and David C. Cassidy, Einstein’s Wife: The Real Story of Mileva Einstein-Marić Cambridge, MA: MIT Press, 2019. Pp. 336. ISBN: 978-0-2625-3897-8. $19.95 (paperback). ISBN: 978-0-2620-3961-1. $29.95 (hardback). British Journal for the History of Science, 55 (4):531–533, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/allan-
esterson-and-david-c-
cassidy-einsteins-wife-the-

real-story-of-mileva-
einsteinmaric-cambridge-ma-
mit-press-2019-pp-336-isbn-
9780262538978-1995-
paperback-isbn-9780262039611-
2995-hardback/EC290FBBFB3E58DC60E32E6AE4B6CD49.

Stepney:2024:VCO

[Ste24] Eva Surawy Stepney. ‘Visible’ compulsions: OCD and the politics of science in British clinical psychology, 1948–1975. British Journal for the History of Science, 57(1):81–97, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/visible-
compulsions-ocd-and-the-
politics-of-science-in-
british-clinical-
psychology-19481975/
D431B7D6003860F9E6ABE50476BA46A4.

Steingart:2025:BRB

[Ste25a] Alma Steingart. Book review:

Benjamin Labatut, The MA-NIAC, New York: Penguin, 2023. Pp. 368. ISBN 978-0-593-65447-7. $28.00 (hardcover). British Journal for the History of Science, 58(2):367–368, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/benjamin-
labatut-the-maniac-new-
york-penguin-2023-pp-368-
isbn-9780593654477-2800-

                                        hardcover/
                                        5B660CC3974704C421247CA8C7AE628F.

Stephens:2025:BRS

[Ste25b] Carlene Stephens. Book review: Shaul Katzir, Sonar to Quartz Clock: Technology and Physics in War, Academy, and Industry, Oxford: Oxford University Press, 2023. Pp. 368. ISBN 978-0-19-887873-5. £45.00 (hardback). British Journal for the History of Science, 58(2):399–400, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/shaul-katzir-sonar-
                                        to-quartz-clock-technology-
                                        and-physics-in-war-academy-
                                        and-industry-oxford-oxford-
                                        university-press-2023-pp-
                                        368-isbn-9780198878735-
                                        4500-hardback/
                                        A45F3718AFC8C685C3B92D38C6CC837F.

Storni:2021:DPD

[Sto21] Marco Storni. Denis Papin’s digester and its eighteenth-century European circulation. British Journal for the History of Science, 54 (4):443–463, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/denis-
                                        papins-digester-and-its-
                                        eighteenthcentury-european-

circulation/
81444C8189DA86253389A007FED65612.

Stolz:2022:IPA

[Sto22]  Daniel A. Stolz.  ‘Impossi-

ble to provide an accurate estimate’: the interested calculation of the Ottoman public debt, 1875–1881. British Journal for the History of Science, 55(4): 477–493, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/impossible-
to-provide-an-accurate-
estimate-the-interested-
calculation-of-the-ottoman-
public-debt-18751881/
9FCA5C2D392B2CB590DFF48DECA8847F.

Stoeger:2023:BRF

[Sto23] Alexander Stoeger. Book review: Felix Lüttge, Auf den Spuren des Wals: Geographien des Lebens im 19. Jahrhundert Göttingen: Wallstein Verlag, 2020. Pp. 279. ISBN 978-3-8353-3680-3. EUR28.00 (hardback). British Journal for the History of Science, 56(1): 129–131, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/felix-
luttge-auf-den-spuren-des-
wals-geographien-des-
lebens-im-19-jahrhundert-
gottingen-wallstein-verlag-

                                        2020-pp-279-isbn-
                                        9783835336803-2800-
                                        hardback/F449FCAD458A941AA4A92DBB738259AD.

Stoeger:2024:BRJ

[Sto24] Alexander Stoeger. Book review: James J.A. Blair, Salvaging Empire: Sovereignty, Natural Resources, and Environmental Science in the South Atlantic, Ithaca, NY: Cornell University Press, 2023. Pp. 318. ISBN 978-1-5017-7154-5. $31.95 (paperback). British Journal for the History of Science, 57 (1):144–146, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/james-ja-
                                        blair-salvaging-empire-
                                        sovereignty-natural-
                                        resources-and-
                                        environmental-science-in-
                                        the-south-atlantic-ithaca-
                                        ny-cornell-university-
                                        press-2023-pp-318-isbn-
                                        9781501771545-3195-paperback/
                                        B6AA42D780B5CE7DC0C78DE44D1E0B12.

Steere-Williams:2025:ERF

[SW25] Jacob Steere-Williams. Essay review: the fictive history of Victorian science and empire. British Journal for the History of Science, 58(1):133–135, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/

article/essay-review-the-
fictive-history-of-
victorian-science-and-
empire/
B097A1B16E0505260EE07CB23A32BABE.

Turchetti:2023:IPI

[TA23] Simone Turchetti and Matthew Adamson. Introduction: Power to the image! Science, technology and visual diplomacy. British Journal for the History of Science, 56:135–146, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
introduction-power-to-the-
image-science-technology-
and-visual-diplomacy/
FDE4F52D3091011E66B1068CF31FEB43.

Talairach:2021:BRC

[Tal21] Laurence Talairach. Book reviews: Charles Smith, James T. Costa and David Collard (eds.), An Alfred Russel Wallace Companion, Chicago and London: The University of Chicago Press, 2019. Pp. 416. ISBN 978-0-226-62210-1. $60.00 (hardback). Michael A. Flannery, Nature’s Prophet: Alfred Russel Wallace and His Evolution from Natural Selection to Natural Theology, Tuscaloosa: University of Alabama Press, 2018. Pp. 280. ISBN 978-0-8173-1985-4. $44.95 (hardback). British Journal for the History of Science, 54 (2):233–236, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/charles-
smith-james-t-costa-and-
david-collard-eds-an-
alfred-russel-wallace-
companion-chicago-and-
london-the-university-of-
chicago-press-2019-pp-416-
isbn-9780226622101-6000-hardback-
michael-a-flannery-natures-prophet-
alfred-russel-wallace-and-his-
evolution-from-natural-selection-
to-natural-theology-tuscaloosa-
university-of-alabama-press-
2018-pp-280-isbn-9780817319854-
4495-hardback/50DEC7EDA05762B4CAB0638B3053CCAC.

Tarrant:2022:BRM

[Tar22] Neil Tarrant. Book review: Mark A. Waddell, Magic, Science and Religion in Early Modern Europe, Cambridge: Cambridge University Press, 2021. Pp. 220. ISBN 978-1-108-34823-2. £69.99/£19.99 (hardback/paperback). British Journal for the History of Science, 55 (1):123–124, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mark-a-
waddell-magic-science-and-
religion-in-early-modern-
europe-cambridge-cambridge-
university-press-2021-pp-
220-isbn-9781108348232-
69991999-hardbackpaperback/
D67F2CAB7EC534A85DC4CCBBA60755AE.

Tarrant:2024:BRM

[Tar24]  Neil Tarrant.  Book review:

Mackenzie Cooley, The Perfection of Nature: Animals, Breeding, and Race in the Renaissance, Chicago: University of Chicago Press, 2022. Pp. 334. ISBN 978-0-226-82228-0. $112.50 (cloth). British Journal for the History of Science, 57(1): 138–140, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mackenzie-
cooley-the-perfection-of-
nature-animals-breeding-
and-race-in-the-
renaissance-chicago-
university-of-chicago-
press-2022-pp-334-isbn-
9780226822280-11250-cloth/
C18C2FD82F97E419C865CFC7F313D1D3.

Tekin:2022:IPG

[Tek22] Kenan Tekin. Islamic philosophy and the globalization of science: Ahmed Cevdet’s translation of the sixth chapter of Ibn Khaldun’s Muqaddimah. British Journal for the History of Science, 55 (4):459–475, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/islamic-
philosophy-and-the-
globalization-of-science-
ahmed-cevdets-translation-

                                        of-the-sixth-chapter-of-
                                        ibn-khalduns-muqaddimah/
                                        C060C8FFB6B8A6E51B0DA2E6835316C9.

teHennepe:2023:VLF

[tH23]  Mieneke te Hennepe.  Van

Leeuwenhoek — the film: remaking memory in Dutch science cinema 1925–c. 1960. British Journal for the History of Science, 56(3): 329–349, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/van-
                                        leeuwenhoek-the-film-
                                        remaking-memory-in-dutch-
                                        science-cinema-1925c1960/
                                        A9BDEFD1D3BA838EAFF76231C74EC530.

Thampuran:2024:BRL

[Tha24] Arya Thampuran. Book review: Lynn M. Thomas, Beneath the Surface: A Transnational History of Skin Lighteners, Durham, NC: Duke University Press, 2020. Pp. 368. ISBN 978-1-4780-0642-

8. $107.95 (cloth). British Journal for the History of Science, 57 (4):661–662, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/lynn-m-
                                        thomas-beneath-the-surface-
                                        a-transnational-history-of-
                                        skin-lighteners-durham-nc-
                                        duke-university-press-2020-
                                        pp-368-isbn-9781478006428-

10795-cloth/
A7AD7F8C44536E173580303A5FFFC5E9.

Thomas:2024:WWF

[Tho24] William Thomas. Why we fight about science. British Journal for the History of Science, 57(4): 651–654, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/why-we-
fight-about-science/
39C1C89CAC05F50C35FAD19E7B0CDA65.

Timberlake:2020:BRM

[Tim20]  Todd K. Timberlake.  Book

review: Martin Beech, Going Underground: The Science and History of Falling through the Earth. New Jersey, London, Singapore, Beijing, Shanghai, Hong Kong, Taipei, Chennai and Tokyo: World Scientific, 2019. Pp. xi + 276. ISBN 978-981-327-903-2. £35.00/ $38.00 (paperback). British Journal for the History of Science, 53(2):287–288, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/martin-
beech-going-underground-
the-science-and-history-of-
falling-through-the-earth-
new-jersey-london-
singapore-beijing-shanghai-
hong-kong-taipei-chennai-
and-tokyo-world-scientific-2019-

                                        pp-xi-276-isbn-9789813279032-
                                        35003800-paperback/FFB6C23B4A6BBD7D65A62AD52C606FE7.

Tomar:2025:BRN

[Tom25] Koyna Tomar. Book review: Neal A. Knapp, Making Machines of Animals: The International Livestock Exposition, Baltimore: Johns Hopkins University Press, 2023. Pp. 216. ISBN 978-1-4214-4655-4. $60.00 (hardcover). British Journal for the History of Science, 58 (2):345–346, June 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/neal-a-
                                        knapp-making-machines-of-
                                        animals-the-international-
                                        livestock-exposition-
                                        baltimore-johns-hopkins-
                                        university-press-2023-pp-
                                        216-isbn-9781421446554-
                                        6000-hardcover/2EEE89CDAE5A9043B46EA39F2BD55E58.

Turney:2021:BRP

[Tur21] Jon Turney. Book reviews: Patrick Manning, A History of Humanity: The Evolution of the Human System. Cambridge: Cambridge University Press, 2020. Pp. xiii + 363. ISBN: 978-1-108-47819-9, £59.99 (hardback). ISBN: 978-1-108-74709-7. £18.99 (paperback).

— Amanda Rees and Charlotte Sleigh, Human London: Reaktion Books, 2020. Pp. 216. ISBN 978-1-78914-214-3. £12.95 (paperback). British Journal for the History of Science, 54(3):

387–389, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/patrick-
manning-a-history-of-
humanity-the-evolution-of-
the-human-system-cambridge-
cambridge-university-press-
2020-pp-xiii-363-isbn-
9781108478199-5999-
hardback-isbn-9781108747097-
1899-paperback-amanda-rees-and-
charlotte-sleigh-human-london-
reaktion-books-2020-pp-216-isbn-
9781789142143-1295-paperback/
DF2F4E40EFFBB4B64BDC00C630622640.

Turchetti:2023:RNS

[Tur23] Simone Turchetti. Representing noise: stacked plots and the contrasting diplomatic ambitions of radio astronomy and postpunk. British Journal for the History of Science, 56:225–245, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/representing-noise-
stacked-plots-and-the-
contrasting-diplomatic-
ambitions-of-radio-
astronomy-and-postpunk/
085E6BEDC4FD45EE330E23395E387F3F.

Valenti:2025:WCT

[Val25] Salvatore Valenti. ‘Who can tell me what potable water means?’ The assessment of water quality in debates over hydraulic infrastructure in nineteenth-century Italy. British Journal for the History of Science, 58(1):101–116, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/who-can-tell-me-
what-potable-water-means-
the-assessment-of-water-
quality-in-debates-over-
hydraulic-infrastructure-
in-nineteenthcentury-italy/
2C126FB87EB63F75556FCEE0A30CCA9A.

VanAusdal:2024:BRT

[Van24] Shawn Van Ausdal. Book review: Timothy W. Lorek, Making the Green Revolution: Agriculture and Conflict in Colombia, Chapel Hill: University of North Carolina Press, 2023. Pp. 342. ISBN 978-1-4696-7382-0. $34.95 (paperback). British Journal for the History of Science, 57(4): 655–657, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/timothy-w-
lorek-making-the-green-
revolution-agriculture-and-
conflict-in-colombia-
chapel-hill-university-of-
north-carolina-press-2023-
pp-342-isbn-9781469673820-
3495-paperback/
417F5E7E3EA44ECD602966E7AAF366B7.

vanDijk:2023:BRL

[vD23] Alexander van Dijk. Book reviews: Lydia Barnett, After the Flood: Imagining the Global Environment in Early Modern Europe. Baltimore: Johns Hopkins University Press, 2019. Pp. 264. ISBN 978-1-4214-2951-9. $52.00 (hardback). ISBN 978-1-4214-4527-4. $28.95 (paperback). British Journal for the History of Science, 56 (1):121–123, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/lydia-
barnett-after-the-flood-
imagining-the-global-
environment-in-early-
modern-europe-baltimore-
johns-hopkins-university-
press-2019-pp-264-isbn-
9781421429519-5200-
hardback-isbn-9781421445274-
2895-paperback/D12F3BF1FF54941FE34C4A5C935778D3.

Virmajoki:2025:BRL

[Vir25] Veli Virmajoki. Book review:

Lukas M. Verburgt (ed.), Debating Contemporary Approaches to the History of Science, London: Bloomsbury Academic, 2024. Pp. 376. ISBN 978-1-3503-2623-1. £75.00 (hardback). British Journal for the History of Science, 58 (1):167–168, March 2025. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-

                                        science/article/lukas-m-
                                        verburgt-ed-debating-
                                        contemporary-approaches-to-
                                        the-history-of-science-
                                        london-bloomsbury-academic-
                                        2024-pp-376-isbn-
                                        9781350326231-7500-
                                        hardback/18D061B205050381B1BEC7A1A122421C.

Walsdorf:2025:BRS

[Wal25a] Anaïs Walsdorf. Book review: Simona Boscani Leoni, Sarah Baumgartner and Meike Knittel (eds.), Connecting Territories: Exploring People and Nature, 1700–1850, Leiden: Brill, 2021. Pp. 268. ISBN 978-90-04-41246-

0. EUR119.00 (hardback). British Journal for the History of Science, 58(1):143–144, March 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/simona-boscani-
                                        leoni-sarah-baumgartner-
                                        and-meike-knittel-eds-
                                        connecting-territories-
                                        exploring-people-and-
                                        nature-17001850-leiden-
                                        brill-2021-pp-268-isbn-
                                        9789004412460-11900-hardback/
                                        EFDD6C5BB8F51D78503305AAE3FFE2BE.

Walsh:2025:BRL

[Wal25b] Fionnuala Walsh. Book review: Laura Kelly, Contraception and Modern Ireland: A Social History, c. 1922–92, Cambridge: Cambridge University Press, 2023. Pp. 350. ISBN 978-1-108-96977-2. $89.99 (hardcover).

British Journal for the History of Science, 58(2):359–360, June 2025. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/laura-kelly-
contraception-and-modern-
ireland-a-social-history-c-
192292-cambridge-cambridge-
university-press-2023-pp-
350-isbn-9781108969772-
8999-hardcover/
F541226433BB2A75B239DAA5298054F7.

Weber:2020:BRJ

[Web20] Thomas P. Weber. Book review:

JoAnne Yates and Craig N. Murphy, Engineering Rules: Global Standard Setting since 1880. Baltimore: Johns Hopkins University Press, 2019. Pp. 440. ISBN 978-1-4214-2889-5. $64.95 (hardcover). British Journal for the History of Science, 53 (2):283–285, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/joanne-
yates-and-craig-n-murphy-
engineering-rules-global-
standard-setting-since-
1880-baltimore-johns-
hopkins-university-press-
2019-pp-440-isbn-
9781421428895-6495-
hardcover/8071585BF0754D4D41D014FE278E0C79.

Weber:2023:BRT

[Web23] Thomas P. Weber. Book review: Trais Pearson, Sovereign Necropolis: The Politics of Death in Semi-colonial Siam/ Ithaca, NY: Cornell University Press, 2020. Pp. 233. ISBN 978-1-5017-4015-2. $49.95 (hardback). British Journal for the History of Science, 56 (1):126–128, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/trais-
                                        pearson-sovereign-
                                        necropolis-the-politics-of-
                                        death-in-semicolonial-siam-
                                        ithaca-ny-cornell-
                                        university-press-2020-pp-
                                        233-isbn-9781501740152-
                                        4995-hardback/
                                        C8013ABE1A4F11556573656D73DF1D0E.

Webster:2024:BRN

[Web24a] Emily Webster. Book review: Neeraja Sankaran, A Tale of Two Viruses: Parallels in the Research Trajectories of Tumor and Bacterial Viruses, Pittsburgh: University of Pittsburgh Press, 2021. Pp.

312. ISBN: 978-0-8229-4630-4. $55.00 (hardcover). British Journal for the History of Science, 57 (1):153–155, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/neeraja-
                                        sankaran-a-tale-of-two-

viruses-parallels-in-the-
research-trajectories-of-
tumor-and-bacterial-
viruses-pittsburgh-
university-of-pittsburgh-
press-2021-pp-312-isbn-
9780822946304-5500-hardcover/
7DCDECBFB27B6DB27FDCE72100C591D5.

Webster:2024:BRS

[Web24b] Emily Webster. Book review:

Stig S. Frøland, Duel without End: Mankind’s Battle with Microbes, (trans. John Irons) London: Reaktion Books, 2022. Pp. 632. ISBN: 978-1-78914-505-2. £30.00 GBP (paperback). British Journal for the History of Science, 57 (4):668–669, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/stig-s-
froland-duel-without-end-
mankinds-battle-with-
microbes-trans-john-irons-
london-reaktion-books-2022-
pp-632-isbn-9781789145052-
3000-gbp-paperback/
2BA157D888642BD737CF0E56EF266041.

Wesner:2021:RES

[Wes21] Samantha Wesner. Revolutionary electricity in 1790: shock, consensus, and the birth of a political metaphor. British Journal for the History of Science, 54(3):257–275, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.

                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/revolutionary-
                                        electricity-in-1790-shock-
                                        consensus-and-the-birth-of-
                                        a-political-metaphor/
                                        1D8ABB62212B1BB32AA805DD543D8EC0.

Whiteley:2024:BRK

[Whi24] Rebecca Whiteley. Book review: Kathleen M. Crowther, Policing Pregnant Bodies from Ancient Greece to Post-Roe America, Baltimore: Johns Hopkins University Press, 2023. Pp. 288. ISBN 978-1-4214-4763-6. $29.95 (hardcover). British Journal for the History of Science, 57(4): 663–664, December 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/kathleen-m-
                                        crowther-policing-pregnant-
                                        bodies-from-ancient-greece-
                                        to-postroe-america-
                                        baltimore-johns-hopkins-
                                        university-press-2023-pp-
                                        288-isbn-9781421447636-
                                        2995-hardcover/
                                        5C10909C9DA7FC078F96E8324FE9C4D4.

Widmalm:2023:SBW

[Wid23] Sven Widmalm. ‘Super Bowl of the world conference circuit’? A network approach to highlevel science and policy conferencing. British Journal for the History of Science, 56(4): 535–551, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/super-bowl-
of-the-world-conference-
circuit-a-network-approach-
to-highlevel-science-and-
policy-conferencing/
9AA1752861077BDED686E99732841139.

Wills:2020:BRB

[Wil20] Hannah Wills. Book review:

Benjamin Wardhaugh, Gunpowder and Geometry: The Life of Charles Hutton, Pit Boy, Mathematician, and Scientific Rebel. London: William Collins, 2019. pp. 312. ISBN 978-0-00-829995-8. £20.00/ $39.99 (hardback). British Journal for the History of Science, 53 (1):129–131, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/benjamin-
wardhaugh-gunpowder-and-
geometry-the-life-of-
charles-hutton-pit-boy-
mathematician-and-
scientific-rebel-london-
william-collins-2019-pp-
312-isbn-9780008299958-20003999-
hardback/DE0F963903D0DE148EBE4495D6B11E3E.

Williamson:2021:FAA

[Wil21a] Fiona Williamson. Framing Asian atmospheres: imperial weather science and the problem of the local c. 1880-1950. British Journal for the History of Science, 54(3):

301–304, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/framing-
                                        asian-atmospheres-imperial-
                                        weather-science-and-the-
                                        problem-of-the-local-c-
                                        18801950/
                                        7BED570465846FF2ABB5FE578F414578.

Williamson:2021:JDT

[Wil21b] Fiona Williamson. Just doing their job: the hidden meteorologists of colonial Hong Kong c. 1883–1914. British Journal for the History of Science, 54(3): 341–359, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/just-doing-
                                        their-job-the-hidden-
                                        meteorologists-of-colonial-
                                        hong-kong-c18831914/
                                        B3FEAB52A790A800459FC8957A737B08.

Winters:2023:BRJ

[Win23a] Lisa Winters. Book review: J. David Archibald, Critical Lives: Charles Darwin. London: Reaktion Books, 2021. Pp. 240. ISBN 978-1-78914-440-6. $19.00 (paperback). British Journal for the History of Science, 56(3):420– 421, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-

journal-for-the-history-of-
science/article/j-david-
archibald-critical-lives-
charles-darwin-london-
reaktion-books-2021-pp-240-
isbn-9781789144406-1900-
paperback/
E5B3E6744516E02E46CA5EFBE30FE462.

Winters:2023:BRP

[Win23b] Lisa Winters. Book review:

Peter H. Hoffenberg, A Science of Our Own: Exhibitions and the Rise of Australian Public Science. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 206. ISBN 978-0-8229-4576-

5. $45.00 (cloth). British Journal for the History of Science, 56:276–277, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/peter-h-
hoffenberg-a-science-of-
our-own-exhibitions-and-
the-rise-of-australian-
public-science-pittsburgh-
university-of-pittsburgh-
press-2019-pp-206-isbn-
9780822945765-4500-cloth/
1FF3BFC7649127CA8B9FE6A904D51F65.

Winckel:2024:HGP

[Win24] Floris Winckel. How the Glaishers pictured snowflakes. British Journal for the History of Science, 57(3):349–368, September 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/

                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/how-the-glaishers-
                                        pictured-snowflakes/
                                        0B10C969CBEAA99F31B666A5FB17C031.

Worboys:2022:IEB

[Wor22] Michael Worboys. Imperial entomology: Boris P. Uvarov and locusts, c. 1920–c. 1950. British Journal for the History of Science, 55(1):27–51, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/imperial-
                                        entomology-boris-p-uvarov-
                                        and-locusts-c1920c1950/
                                        0B29C73859E2E6490020ED0B26956827.

Wynter:2021:AFL

[Wyn21] Rebecca Wynter. Ambition, ‘failure’ and the laboratory: Birmingham as a centre of twentiethcentury British scientific psychiatry. British Journal for the History of Science, 54(1):19–40, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/ambition-failure-
                                        and-the-laboratory-
                                        birmingham-as-a-centre-of-
                                        twentiethcentury-british-
                                        scientific-psychiatry/
                                        4263E8AE6733A2CAEAB6B72D73A2713B.

Yalcinkaya:2022:GSR

[Yal22] M. Alper Yalçinkaya. Globalizing ‘science and religion’: examples from the late Ottoman Empire. British Journal for the History of Science, 55(4): 445–458, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/
globalizing-science-and-
religion-examples-from-the-
late-ottoman-empire/
5BDB1E43A0A740A982C772EA5109787F.

Yeon-Jung:2021:UGI

[YJ21]  Ji  Yeon-

Jung. Unfriendly guardians: India’s first nuclear leadership change in 1966. British Journal for the History of Science, 54(1):1–17, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/unfriendly-
guardians-indias-first-
nuclear-leadership-change-
in-1966/
AAB306F55A8383C4C70B50C19F754A97.

Zaidi:2023:PSC

[Zai23] Waqar H. Zaidi. The Pugwash scientists’ conferences, Cyrus Eaton and the clash of internationalisms, 1954–1961. British Journal for the History of Science, 56(4):503–517, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/pugwash-scientists-
                                        conferences-cyrus-eaton-
                                        and-the-clash-of-
                                        internationalisms-19541961/
                                        17262B4D46E0F113F4B97A940294C654.